### v2

# Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values

Abstract.
The potential impact of an academic paper is determined by various factors, including its popularity and contribution. Existing models usually estimate original citation counts based on static graphs and fail to differentiate values from nuanced perspectives. In this study, we propose a novel graph neural network to Disentangle the Potential impacts of Papers into Diffusion, Conformity, and Contribution values (called DPPDCC). Given a target paper, DPPDCC encodes temporal and structural features within the constructed dynamic heterogeneous graph. Particularly, to capture the knowledge flow, we emphasize the importance of comparative and co-cited/citing information between papers and aggregate snapshots evolutionarily. To unravel popularity, we contrast augmented graphs to extract the essence of diffusion and predict the accumulated citation binning to model conformity. We further apply orthogonal constraints to encourage distinct modeling of each perspective and preserve the inherent value of contribution. To evaluate models’ generalization for papers published at various times, we reformulate the problem by partitioning data based on specific time points to mirror real-world conditions. Extensive experimental results on three datasets demonstrate that DPPDCC significantly outperforms baselines for previously, freshly, and immediately published papers. Further analyses confirm its robust capabilities. We will make our datasets and codes publicly available.
Impact Prediction, Hierarchical and Heterogeneous Graph, Graph Neural Network, Disentangled Representation Learning, Contrastive Learning
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Theory of computation Dynamic graph algorithmsccs: Mathematics of computing Graph algorithms
1. Introduction
In paper retrieval and recommendation, ranking papers based on their scientific impact can aid researchers to delve into intricate research efforts. This is particularly significant considering the exponential annual growth in the number of published papers (Lo et al., 2020; Chu and Evans, 2021; Xue et al., 2023). Given the inherent difficulties in quantifying the precise numerical value of scientific impact, citation count is regularly employed as a rough approximation (Evans and Reimer, 2009; Sinatra et al., 2016; Jiang et al., 2021). Nevertheless, the current ranking depends on the cumulative number of citations received to date, which merely represents the impact within the prior research environment. Therefore, it becomes imperative to cultivate a sophisticated model that can assess the potential impact of a research paper, placing significant emphasis on the future (Wang et al., 2013).
At a given point in time (e.g., at present), estimating the anticipated increment in citations can effectively facilitate a fair comparison of the potential impact between long-published papers and newly published ones. Actually, the citations received by papers are not solely derived from the influence of their contributions, but are also affected by a multitude of other factors (Bornmann and Daniel, 2008; Case and Higgins, 2000). Hence, rather than directly forecasting original citation counts, it is more effective to disentangle citation increments received from extrinsic factors and the actual contributing factor, making impact prediction more practical and interpretable. Afterwards, we can identify exceptional works amidst the vast array of publications by discerning the potential contributing impact. In this work, we emphasize the popularity factor of a paper as a prominent extrinsic factor affecting the increase of citations. More precisely, when referring to the popularity factor of a paper, we are considering its citation diffusion patterns and the collective conformity degree within the citation network.
Regarding citation diffusion, we posit that influential papers assume an amplifying effect in shaping the process of information dissemination within the citation network (Page et al., 1998; Wu et al., 2019). This leads to increased visibility of papers connected to them. In other words, papers cited by highly-cited papers will receive significantly more citations than others. The statistics of the dataset provides empirical support. For instance, in the field of Computer Science, the former papers will receive an average of 1.5475 citations (after applying a logarithmic transformation), while the latter ones will only receive an average of 0.8580 citations. Towards collective conformity, existing studies have found that some researchers may tend to cite papers associated with well-known entities, irrespective of their actual relevance or genuine contribution (Thornley et al., 2015). This phenomenon is substantiated by the future disparity (2.1917 vs. 0.8009) in papers with different accumulated citations, which aligns with the concept of Matthew effect in the academic domain (Allison and Stewart, 1974).
Therefore, by separating out the influence of the popularity, i.e., citation diffusion and collective conformity, we can gain a closer approximation to the genuine contributing impacts of papers. In essence, we aim to disentangle the citation increment into diffusion, conformity, and contribution values (as shown in Figure 1 (a)).
Figure 1. (a) We disentangle the citation increment into diffusion, conformity, and contribution values for better interpretability. (b) To imitate the evaluation of future dynamics of the citation context, we split the dataset by the observation time point and categorize the samples into previous, fresh, and immediate papers in terms of their publication time. (c) We represent the heterogeneous subgraph schema with all metadata, where each type of relation is bidirectional.
Typically, conventional research (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022; Xu et al., 2022) on citation prediction imposes stringent constraints to select favorable samples by removing lowly-cited papers. Moreover, the division of these data into training, validation, and testing is carried out randomly, disregarding the chronological order of paper publications (Ruan et al., 2020; Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023). However, this problem setting deviates from real-world scenarios, possibly causing a substantial distribution shift. Hence, the model may face challenges in handling recently published papers. Random data partitioning disregards new citation network updates, hampering the model’s practicality and robustness in upcoming real-world test. This limitation impedes the exploration of the potential value of papers reactivated in the evolved context (Van Raan, 2004).
In contrast, our work emulates reality by performing a novel problem setting. We include all papers without strict filtering to maintain consistency with the data distribution of real-world citation network. As illustrated in Figure 1 (b), we partition the dataset based on particular observation time points. This can introduce fresh papers that are not part of the training set (previous papers) into the test set. We further consider immediate papers (published at the test time point) that closely align with cold-start scenarios to validate the robustness of our model. Thus, papers in the test set can be divided into previous, fresh, and immediate papers, facilitating a more practical evaluation. Notably, previous papers have different contexts and future increments at the train and test time points. By employing this approach, our model is incentivized to fit the dynamic citation context, thereby enhancing its generalization.
Existing graph disentanglement studies usually aim at decomposing abstractive hidden factors from either static or homogeneous graphs (see Section 2) (Ma et al., 2019a; Yang et al., 2020; Wang et al., 2020). Instead, our focus lies in disentangling the categorical characteristics of citation behaviors within dynamic and heterogeneous citation graphs. In this study, tailored to each target paper, we construct a Dynamic Heterogeneous Graph from the citation network. The graph consists of a sequence of heterogeneous snapshots (see Figure 1 (c)). This graph retains both the content of the paper and its dynamic citation context. Based on the proposed graph, we devise a model that could Disentangle the Potential impacts of Papers into Diffusion, Conformity, and Contribution Values (called DPPDCC). This model utilizes Citation-aware Relational Graph Convolutional Networks (R-GCNs) to capture the structural features of the input sequence of snapshots. The outputs are then passed through a Transformer Encoder, which allows for the incorporation of temporal information. Specially, we design a novel module in modified R-GCN to integrate the comparative and co-cited/citing information between paper nodes. Furthermore, we propose a Type-specific Attention Readout mechanism to aggregate information from each snapshot effectively. Upon acquiring the final representation of the target paper, we proceed to disentangle the predicted citation increment into diffusion, conformity, and contribution values individually. For the diffusion, we employ Triplet Citation-aware Graph Contrastive Learning to capture the popularity influence inherent in the information diffusion process within the citation network. For the conformity, we introduce an auxiliary task to classify bins of accumulated citations. Additionally, we impose an orthogonal constraint on the representations of disentangled perspectives in pair, guaranteeing that they encode distinct features without overlap. Extensive experiments are conducted on three fields of the real-world scientific dataset S2AG (Kinney et al., 2023). The results illustrate the superior performance of DPPDCC compared to baseline models across previous, fresh, and immediate papers. Ablation tests further highlight the significance of the proposed components. Additionally, visualizations offer insights into the model’s interpretability.
To sum up, the contributions of this research are as follows:
$\bullet$ We revise the problem formulation of impact prediction to better align with real-world scenarios. Without filtering the predicted samples by their citation counts, we partition the datasets based on specific time points to maintain the dynamics of the citation context. Moreover, we employ disentangled representation learning to extract the contributing impacts of papers.
$\bullet$ We propose a novel model called DPPDCC that first encodes comparative and co-cited/citing information evolutionarily within the Dynamic Heterogeneous Graph through Citation-aware Graph Neural Network Encoder for potential impact prediction. It further disentangles the citation increment into diffusion, conformity, and contribution values.
$\bullet$ Experimental results show significantly superior to existing baselines for previous, fresh, and immediate papers. Further analyses illustrate that our model can reasonably make predictions.
2. Related Work
We review three lines of related work: Citation/Cascade Prediction, Dynamic and Heterogeneous Graph Neural Network (GNN), and Disentangled Representation Learning.
2.1. Citation/Cascade prediction
Citation prediction is a vital sub-task of automated academic evaluation as it enables the estimation of potential impact. It is closely examined within cascade prediction tasks, which share analogous graph structures and objectives (Zhou et al., 2021). The primary goal of cascade prediction is to anticipate the popularity of a post based on the interactions between users. Similarly, in the context of citation networks, the paper can serve as the post, while the authors can serve as the users. Citation count prediction places a focus on the content and distinctive features of individual papers, heavily relying on comprehensive feature extraction techniques. In contrast, cascade prediction delves into the intricate relationships between various entities, including papers and authors, directly exploiting the information of graph structure.
Their approaches can be broadly classified into three categories: stochastic models, feature-based models, and deep learning models. Stochastic models can predict future citation counts by fitting the curve of citation counts (Glänzel and Schubert, 1995), following Zipf-Mandelbrot’s law (Silagadze, 1997). Recently, machine learning models have shown promising results by utilizing manually extracted features from various meta-data (Yan et al., 2011; Yu et al., 2014; Ruan et al., 2020; Guille et al., 2013; Tatar et al., 2014). More presently, deep neural networks have dominated this task by applying advanced Natural Language Processing (NLP) and Computer Vision (CV) methods to extract abstractive representations from paper content (Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023). For cascade prediction, graph embeddings, sequence models, and GNN models are employed to extract structural information from the underlying graph, and then temporal information is encoded with sequence models (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022; Xu et al., 2022).
However, existing models for citation and cascade prediction exhibit sub-optimal performance. They fail to fully exploit the valuable information presented in paper content and scientific context within the citation network simultaneously. Furthermore, their strict data selection and random data splitting strategies may hinder their practicality. Many valuable samples will be filtered and the temporal information within dynamic context will be neglected. Thus, models in this setting are more prone to face distribution shift issues. In contrast, our study proposes to specifically revise these settings and utilize both content and dynamic context for impact estimation. We also emphasize retaining genuine contributions to bolster the robustness of our models.
2.2. Dynamic/Heterogeneous Graph Neural Networks
GNNs (Sperduti and Starita, 1997; Gori et al., 2005; Scarselli et al., 2008; Wu et al., 2020) are widely applied to handle non-Euclidean data like graphs. Recently, research interest has surged in dynamic graphs, which incorporate temporal information, as well as in heterogeneous graphs, which involve multiple node types or edge types. These types of graphs are more prevalent in practical scenarios and encapsulate more intricate information, compared to static homogeneous graphs, driving the formulation and refinement of diverse methodologies within the academic realm.
The dynamic graph can be divided into multiple snapshots at different time points. Previous models like DGCN (Manessi et al., 2020), Dysat (Sankar et al., 2020), and ROLAND (You et al., 2022) typically encode the snapshots with static GCN as the structural encoder, followed by sequential models like Recurrent Neural Networks (RNN) or Transformer Encoder. Moreover, certain models incorporate GCN into RNN models by configuring the parameters within RNN cells as GCN parameters (Pareja et al., 2020). Within heterogeneous graphs, on one side, R-GCN serves as a prominent model with its separate message-passing architecture (Schlichtkrull et al., 2018). It conducts independent message passing within different relations and aggregates them to update node representations. On the other side, the relation-learning method transforms heterogeneous graphs into homogeneous ones, while retaining node and edge type indicators to preserve heterogeneous information (Hu et al., 2020; You et al., 2022). An example of this technique is HGT (Hu et al., 2020), which utilizes multi-head attention modules with specific parameters tailored to different node and edge types, effectively modeling the complexity of heterogeneous graphs.
Citation network can be regarded as dynamic and heterogeneous graphs, as it evolves over time and encompasses diverse entities such as papers, authors, and venues. Addressing the complexities within this network involves two primary challenges: (1) How to encode the temporal target-centric information within a single snapshot? (2) How to model distinct characteristics of information diffusion in citation networks? To overcome these challenges, our study proposes a Citation-aware GNN Encoder that effectively models both dynamic and heterogeneous graphs of target papers. These graphs are constructed from the citation network on an annual basis and incorporate multiple metadata nodes beyond papers themselves. (1) To encode the temporal target-centric information within a single snapshot, we employ a type-aware attention readout. It aggregates snapshot information considering paper type and publication time. (2) To model distinct characteristics of information diffusion in citation networks, we propose a novel GCN module CompGAT. This module focuses on capturing comparisons and co-cited/citing information between papers, drawing inspiration from bibliometrics theories (Wu et al., 2019).
2.3. Disentangled Representation Learning
Disentangled representation learning, aiming to discern and separate fundamental explanatory factors (Bengio et al., 2013), stands as a pivotal method within deep learning. It strives to produce resilient, manageable, and interpretable representations. Existing research predominantly focuses on the field of CV, often employing methods like Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) to disentangle relevant latent factors (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021). Researchers explore techniques to align distinct representations with specific factors through explicit and implicit guidance, employing approaches like direct labeling, loss penalties, and regularization. For instance, DR-GAN (Tran et al., 2017) utilizes an adversarial network to generate diverse facial poses, aiming to disentangle pose information from facial attributes. In NLP, disentangled representation learning finds applications in various tasks like generation. For example, John et al. (John et al., 2018) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style. This separation allows precise control over content and style in text generation.
Recently, there has been a growing interest in applying disentangled representation learning to graph learning tasks. Given that graph data encompasses both structural and attributed information, numerous new methodologies have emerged to address this complexity. In the feature space, inspired by capsule networks, approaches have emerged to partition node features into multiple hidden channels (Ma et al., 2019a). Some methods introduce additional constraints to enable effective separation (Liu et al., 2020). Moving into the structural space, subsequent models strive to factorize the input graphs into multiple subgraphs, to facilitate distinct message-passing strategies (Yang et al., 2020; Li et al., 2021). Recent models have expanded their scope to isolate causal and biased information presented in both the graph structure and node features, with a particular emphasis on causal effects (Sui et al., 2022; Fan et al., 2022). These models commonly employ techniques to generate masks within the adjacency matrix and node features. This process aims to identify the crucial elements of the graph within the structural and attributed spaces. In addition, certain studies extend disentangled GNNs to specialized scenarios such as recommendation systems, handling intricate graph structures in heterogeneous or dynamic graphs (Wang et al., 2020; Zhang et al., 2022; Wen et al., 2022). Besides the masking strategy considering solely the graph structures and node features, they integrate classical methods like VAE and adversarial attack learning from CV and NLP to handle task-specific information from a broader perspective.
In contrast to previous studies, our approach acknowledges that the potential impact within the citation network is influenced not only by the paper’s contribution but also by factors related to its popularity. These popularity factors encompass the amplifying effect owing to the highly-cited paper nodes in information diffusion and the manifestation of the Matthew effect through collective conformity. Based on the dynamic heterogeneous graph, we disentangle the potential impact into diffusion, conformity, and contribution values, seeking a more comprehensive estimation. Specifically, we devise dedicated auxiliary tasks aimed at extracting values associated with these respective factors.
3. Methodology
In this section, we present a detailed exposition of our proposed model, DPPDCC. Initially, we formulate the constructed dynamic heterogeneous and the potential impact prediction task. Following this, we provide an in-depth elucidation of DPPDCC’s core components: the Citation-aware GNN Encoder (CGE) and the Popularity-aware Disentanglement Module (PDM), as depicted in Figure 2. The CGE integrates multiple layers of structural and temporal encoders, aimed at capturing intricate high-order temporal and structural information within the network. Concurrently, the PDM comprises three distinct modules, each targeting a distinct perspective of citation increment, all trained in parallel. Lastly, we delineate the training and optimization strategies employed for DPPDCC.
Figure 2. The overall architecture of DPPDCC. Initially, DPPDCC processes the dynamic heterogeneous graph of the target paper using the Citation-aware GNN Encoder. Within this encoder, the structural component utilizes CompGAT to extract comparative and co-cited/citing information between papers at each time step. This information is then aggregated evolutionarily across snapshots via type-specific attention readout. Subsequently, aggregated snapshot information from different time steps is interacted in the temporal encoder. By stacking multiple layers of the structural encoder and temporal encoder, it enables CGE to encode complex high-order temporal and structural information effectively. Subsequent to obtaining the target paper’s representation, DPPDCC proceeds to disentangle its citation increment into diffusion, conformity, and contribution values through three corresponding auxiliary tasks. The diffusion module augments the graphs by employing weighted node dropping, facilitating the identification of the diffusion process by contrasting augmented views with the original graph. The conformity module quantifies conformity levels by binning accumulated citations. Finally, the contribution module introduces orthogonal constraints, prompting each module to learn distinct information from diverse perspectives.
3.1. Problem Formulation and Notations
Before presenting our approach, we first introduce the problem formulation and notations. In practice, the citation network comprises diverse entities and relations that evolve dynamically each year. To capture the entity-aware and time-variant characteristics, of each target paper, we construct a dynamic heterogeneous graph from the citation network to predict its potential impact.
3.1.1. Dynamic Heterogeneous Graph
Given a target paper at the time step $t$, its heterogeneous subgraph is defined as $G^{t}=(\nu^{t},\epsilon^{t},\phi,\varphi)$. Here, $\nu^{t}$ denotes the set of nodes surrounding the target paper at the time step $t$. And $\epsilon^{t}$ represents the set of edges among $\nu^{t}$. Each node $\nu$ and edge $\epsilon$ are associated with their type mapping functions $\phi:\nu\rightarrow U$ and $\varphi:\epsilon\rightarrow R$, where $U$ and $R$ denote the types of nodes and edges (see Figure 1 (c)), respectively.
The backbone of the heterogeneous subgraph primarily consists of the cited and citing relations. We collect $k$-hop cited/citing neighbors of the target paper at the specific time $t$. For each $i$-th hop, to maintain the most recent and significant information, we select the $K_{i}$ newest published and then most highly cited papers, for references and citations respectively. To address the potential lack of global information, we incorporate metadata nodes associated with the papers (authors, venue, and publication time). Prior to the observation time point, we consider a total of $T$ such subgraphs, each serving as a snapshot captured at distinct time steps. These subgraphs organized into a sequence collectively constitute the dynamic heterogeneous graph $G$ of the target paper.
3.1.2. Potential Impact Prediction
We assess the potential impact of the given target paper by estimating its citation increment. It is quantified as the increased citation number $\Delta$ years after the observation time point. Given a paper, we try to learn a function $f(\cdot)$ to predict the number of citation increments.
3.2. Initialization
The initial embedding of the paper node is derived from the text embedding obtained by encoding the combination of title and abstract. Metadata node embeddings in each snapshot are the average embeddings over all papers associated with them in the global citation network at the corresponding time step. Specifically, the extraction of text embeddings is accomplished using Sentence-BERT (Reimers and Gurevych, 2019). For $i$-th node at the $t$-th time step, we have:
(1) $\begin{split}&\bm{h}_{ti,paper}^{(0)}=TextEncoder([w_{1},w_{2},...,w_{n^{w}}])\\ &\bm{h}_{ti,u}^{(0)}=avgpool([\bm{h}_{t1,paper}^{(0)},\bm{h}_{t2,paper}^{(0)},...,\bm{h}_{tm_{p},paper}^{(0)}]),\\ &u\in\{author,venue,time\}\end{split}$
where $u$ denotes the node type, $TextEncoder$ denotes the text representation model, and $m_{p}$ is the count of metadata node’s connected papers.
To initialize the 00-th layer representation of the snapshot of the $i$-th target paper at the $t$-th time step $\bm{h}_{ti,s}^{(0)}$, we mainly consider the contextual information from the corresponding time embedding $\bm{h}_{t,time}^{(0)}$ at time step $t$. Since the time embedding only preserves the global features, we utilize a gate module to approach specific information based on the $i$-th target paper embedding $\bm{h}_{i,p}^{(0)}$. The snapshot representation $\bm{h}_{ti,s}^{(l)}$ of $l$-th layer will then be the proxy to aggregate the particular heterogeneous subgraph:
(2) $\begin{split}&\bm{h}_{ti,s}^{(0)}=\bm{a}_{ti}\times\bm{h}_{t,time}^{(0)}\\ &\bm{a}_{ti}=\mathrm{sigmoid}(\bm{W}^{ipt}\bm{h}_{t,time}^{(0)}+\bm{W}^{tgt}\bm{h}_{i,p}^{(0)})\end{split}$
3.3. Citation-aware GNN Encoder
Citation-aware GNN Encoder (CGE) aims to capture distinct structural and temporal features from the dynamic heterogeneous graph, which is facilitated by leveraging knowledge diffusion theories derived from bibliometrics. To be specific, the backbone of CGE consists of modified R-GCNs as the structural encoder and a Transformer Encoder as the temporal encoder. We stack $L$ integrated layers to encode them alternately instead of separate modeling. This allows us to exploit higher-order interactions between space and time. To elaborate, we first employ distinct Citation-aware R-GCNs to model structural information for subgraphs at different time points individually. Afterwards, a Type-specific Attention Snapshot Readout is employed to summarize the information from these subgraphs established on the snapshot representations. Simultaneously, the snapshot information mixer integrates global and temporal information from snapshot representations into the local and structural representations of paper nodes, fostering interactions among various feature spaces. The aggregated snapshot representations are then input into the temporal encoder (Encoder of the Transformer) to capture the intricate temporal relationships between different time points. By integrating structural and temporal encoding, our approach could extract valuable insights from the dynamic heterogeneous graph.
3.3.1. Citation-aware R-GCNs
Citation-aware R-GCNs, as the core component of the structural encoder, are utilized to encode node hidden states $\bm{h}$ within the dynamic heterogeneous graph. Following the manner of R-GCN (Schlichtkrull et al., 2018), we deliver message passing in each relation separately for modeling attentively. Moreover, we replace the original GCNs in R-GCN with more powerful GNN modules to sensitively encode intricate information.
(3) $\begin{split}&\bm{h}_{dst}^{(l+1)}=\underset{r\in\mathcal{R},r_{dst}=dst}{AGG}(f_{r}(g_{r},\bm{h}_{r_{src}}^{(l)},\bm{h}_{r_{dst}}^{(l)}))\\ \end{split}$
where $\mathcal{R}$ is the set of all edges in the graph, $g_{r}$ is the subgraph with relation $r$, and $AGG$ is the relation aggregation function. Here, we adopt sum-pooling as the relation aggregation function for efficiency and performance.
For ”cites” and ”is cited by” edge subgraphs of paper nodes, inspired by Disruption-index (Wu et al., 2019), we introduce co-cited/citing strengths to integrate multi-hop citation information. The Disruption-index categorizes papers as disruptive or developmental based on the metric related to their betweenness centrality. Disruptive papers play a critical role in altering typical citation diffusion. We adapt this subgraph-focused metric to estimate node-pair closeness by considering co-cited/citing strengths. We aim to discern the varying significance of citation behaviors during message passing and distinguish crucial references and citations within ”is cited by” and ”cites” edges for the target paper. On one side, within ”is cited by” edges, when the source (reference) and target (citation) papers share more similar references, the influence of the source paper on the target paper might diminish. From the viewpoint of the target paper, the source paper could be considered as merely developmental, exerting a potentially negative impact on message passing. Moreover, since the information from this source paper is already conveyed through other original disruptive references, the negative co-citing effect might amplify attention toward these disruptive source papers. On the other side, within “cites” edges, if the source (citation) and target (reference) paper have more identical citations, they may share more commonness and are more closely connected, leading to a positive effect. Unlike co-citing relationships, citations are expansive and continuously growing, making it challenging to differentiate disruptive papers based on co-cited relationships alone. Therefore, we employ a normalized score of co-cited strength that considers both similarity and accumulated count to identify highly-related developmental citations. It’s crucial to note that the co-cited/citing strengths are dynamic scores that evolve over time. As new papers emerge, the values of normalized co-cited/citing strengths undergo changes not only due to the direct influence of new citations but also because of alterations in numerical values and maximum values. Leveraging these co-cited/citing strengths can help uncover the dynamics inherent in multi-hop information diffusion. Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2021), to encode comparative and co-cited/citing features between paper nodes. To capture the discrepancy between source and target nodes, we encode the concatenation of their representations. Besides the basic attention scores of GATv2, we incorporate co-cited or co-citing strengths by assessing the normalized similarity of their references or citations in the global citation network. Based on the adjacent matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes. Finally, the attention is a mixed distribution resembling CopyNet (Gu et al., 2016). For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$, we have:
(4) $\begin{split}&\bm{h}_{i}^{(l+1)}=\mathrm{LeakyReLU}\left(\mathrm{layernorm}(\bm{h}_{i}^{(l)}+\bm{h}_{in}^{(l+1)})\right)\\ &\bm{h}_{in}^{(l+1)}=\sum_{j\in\mathcal{N}(i)}\alpha_{ij}^{(l)}\bm{W}_{c}^{(l)}[\bm{W}^{(l)}_{left}\bm{h}_{i}^{(l)}\|\bm{W}^{(l)}_{right}\bm{h}_{j}^{(l)}]\\ &\alpha_{ij}^{(l)}=\lambda\mathrm{softmax_{i}^{c}}(c_{ij})+(1-\lambda)\mathrm{softmax_{i}^{e}}(e_{ij}^{(l)})\\ &e_{ij}^{(l)}=\bm{w}_{a}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{left}\bm{h}_{i}^{(l)}+\bm{W}^{(l)}_{right}\bm{h}_{j}^{(l)}\right)\end{split}$
where $c_{ij}\in\bm{C}$ is the normalized co-cited/citing strength, $\lambda$ is to balance the sum distribution of co-cited/citing strengths and basic attention weights, and $\bm{w}_{a}^{(l)}$is the attention vector.
For other edges in the R-GCN, we apply GIN (Xu et al., 2018) with sum-pooling as the GNN encoder:
(5) $\begin{split}\bm{h}_{i,r_{dst}}^{(l+1)}=f_{\Theta,r}((1+\xi)\bm{h}_{i,r_{dst}}^{(l)}+AGG(\{\bm{h}_{j,r_{src}}^{(l)},&j\in\mathcal{N}(i,r_{src})\})),r\in\mathcal{R}^{normal}\end{split}$
where $\xi$ is the hyper-parameter to decide the retained original information, $\mathcal{R}^{normal}$ is the set including all edges in classic GIN, and $AGG$ is the neighborhood aggregation function of the GIN. In our model, we opt for sum-pooling as the aggregation method, as it can preserve crucial structural information including node degrees.
3.3.2. Type-specific Attention Snapshot Readout
To represent the time-aware information and filter noisy information of a single snapshot $\bm{h}_{t,s}$, we design an attention-based readout mechanism. It extracts the evolved temporal information of different types of paper nodes (reference, citation, and target paper) based on the publication age of the target paper.
(6) $\begin{split}&\bm{h}_{t,s}^{(l+1)}=\bm{h}_{t,s}^{(l)}+\sum_{j\in\mathcal{N}(t)}\hat{\alpha}_{tj}^{(l)}\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\\ &\hat{\alpha}_{tj}^{(l)}=\mathrm{softmax_{t}}(\hat{e}_{tj}^{(l)})\\ &\alpha_{tj}^{(l)}={\bm{w}_{a}}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{s}\bm{h}_{t,s}^{(l)}+\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\right)+{\bm{w}_{t}}^{(l)}\mathrm{LeakyReLU}\left(f^{(l)}_{se}(\omega^{s}_{t})+f^{(l)}_{pe}(\omega^{p}_{j})\right)\end{split}$
where $\mathcal{N}(t)$ denotes the paper nodes within $t$-th 1-hop subgraph, $f^{(l)}_{se}(\omega^{s}_{t})$ and $f^{(l)}_{pe}(\omega^{p}_{j})$ are snapshot and paper type embeddings. The snapshot type is determined by binning the time interval between the time step $t$ and the publication time of the target paper. In this study, we categorize snapshot types based on temporal segments across all datasets: $\leq 5$ years, $5-10$ years, and $>10$ years. To enhance relevance across diverse fields of study, a potential refinement could involve tailoring these segments to align with the half-life of literature aging within specific domains.
3.3.3. Snapshot Information Mixer
To foster interactions between temporal and spatial dimensions, as well as between global and local information, we amalgamate the representations of paper nodes with the updated snapshot representation:
(7) $\begin{split}&\hat{\bm{h}}_{ti,paper}^{(l+1)}=mlp_{m}^{(l)}([\bm{h}_{ti,paper}^{(l+1)}\|\bm{h}_{t,s}^{(l+1)}])\end{split}$
where $mlp_{m}^{(l)}$ integrates and condenses the information from both the current paper node and the snapshot. leveraging the temporal information embedded in the snapshot representations through the mixer, we introduce extra temporal signals, enriching the structural encoder beyond sole reliance on graph structures.
3.3.4. Snapshot Temporal Encoder
To encode the temporal information across various snapshot times, we utilize a 4-layer Transformer Encoder. Compared with classic sequential models like RNN, the self-attention mechanism captures more intricate relationships that extend beyond strict ordinal dependencies:
(8) $\begin{split}&\hat{\bm{H}}_{s}^{(l+1)}=\mathrm{MultiHeadSelfAttn}(\bm{H}_{s}^{(l+1)})\end{split}$
3.3.5. Final Output
We obtain the target paper representation $\bm{o}$ by avg-pooling across the layers and then sum-pooling across the time steps, to maintain high-order temporal information:
(9) $\begin{split}&\tilde{\hat{\bm{H}}}_{s}=avgpool^{L}([\hat{\bm{H}}_{s}^{(1)},\hat{\bm{H}}_{s}^{(2)},\cdots,\hat{\bm{H}}_{s}^{(L)}])\\ &\bm{o}=sumpool^{T}(\tilde{\hat{\bm{H}}}_{s})\end{split}$
3.4. Popularity-aware Disentanglement Module
Based on the encoded representation of the target paper from CGE, we design the Popularity-aware Disentanglement Module (PDM) to disentangle the citation count potentially received from different perspectives, including citation diffusion, collective conformity, and actual contribution. Overall, we sum up the individual predicted values to obtain the total citation increment and apply different auxiliary tasks to help the disentanglement. This approach enables the model to learn distinctive compositions for diverse papers, thereby facilitating the identification and utilization of their authentic contributions. We employ distinct Multilayer Perceptrons (MLPs) as encoders to process information within the target perspectives, and then feed them into another MLP separately to predict the corresponding citation increment $\hat{y}_{i}^{v}$:
(10) $\begin{split}&\hat{\bm{o}}_{i}^{v}=mlp^{v}_{enc}(\bm{o}_{i})\\ &\hat{y}_{i}^{v}=mlp^{v}_{pred}(\hat{\bm{o}}_{i}^{v})\\ \end{split}$
where $mlp^{d}_{enc}$ encodes specific aspects of the information and $mlp^{v}_{pred}$ predicts the citation increment from the target perspective. It is noteworthy that the three dedicated components are separately modeled to disentangle distinct perspectives, and they are trained simultaneously to facilitate comprehensive and concurrent learning.
3.4.1. Diffusion Encoder
The diffusion encoder extracts the popularity influence according to information diffusion, which is represented as diversity affected by spreading nodes’ degrees. We propose a Citation-aware Triplet Graph Contrastive Learning method to extract the trunk of spreading in the dynamic graph. In various snapshots, except for the target paper node, all other paper nodes are selectively omitted with probability, which is weighted by their normalized global citations within the local subgraph. Lowly-cited paper nodes are more likely to be dropped for the positive view $z^{pos}$, as they contribute less to spreading. Conversely, for the negative view $z^{neg}$, highly-cited paper nodes are more likely to be dropped, as they significantly impact the graph semantics. For every sample, we contrast them solely with their augmented views to preserve their unique characteristics from the diffusion perspective. We augment the subgraphs in every training epoch to enhance the robustness of learning. Through this approach, we amplify diffusion influence in the graph structure. Contrasted with original sample $z^{ori}$, the diffusion loss $\mathcal{L}_{dif}$ is:
(11) $\begin{split}&\bm{z}_{i}^{ori\vee pos\vee neg}=norm(mlp(\hat{\bm{o}}_{i}^{dif,ori\vee pos\vee neg}))\\ &\mathcal{L}_{dif}=\frac{1}{N}\sum_{i=1}^{N}{-log\frac{exp(sim(\bm{z}^{ori},\bm{z}^{pos})/\tau)}{\sum^{2}{exp(sim(\bm{z}^{ori},\bm{z}^{neg\vee pos})/\tau)}}}\end{split}$
where $sim(\cdot)$ is the similarity function like dot product and $\tau$ is the temperature parameter.
3.4.2. Conformity Encoder
The conformity encoder extracts the popularity attributed to the reputation of associated entities. Commonly, researchers tend to cite papers with significant accumulated citations instead of other potentially more relevant papers. In our work, we leverage the accumulated citations of the target paper at the predicted time as a signal of conformity. To extract its influence, we divide the predicted samples into equal frequency bins to quantify the Matthew effect. By predicting the binning label $y_{im}^{con}$, we prompt the model to learn the group differences resulting from the accumulated citations. Here, the conformity loss $\mathcal{L}_{con}$ is:
(12) $\begin{split}&\hat{y}_{im}^{con}=mlp(\hat{\bm{o}}_{i}^{con})\\ &\mathcal{L}_{con}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{m=1}^{M}y_{im}^{con}\log(\hat{y}_{im}^{con})\end{split}$
where $m$ is the number of bins for equal frequency binning.
3.4.3. Contribution Encoder
After disentangling the popularity influences, the remaining citation increment can serve as the genuine contribution. To further encourage distinct encoding of various aspects for better approximation of contribution, we apply an orthogonal regularization $\mathcal{L}_{ort}$ for each pair of perspectives:
(13) $\begin{split}&\mathcal{L}_{ort}=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\hat{\bm{o}}_{i}^{v1}*\hat{\bm{o}}_{i}^{v2}\|^{2}}{\|\hat{\bm{o}}_{i}^{v1}\|*\|\hat{\bm{o}}_{i}^{v2}\|}\end{split}$
[…]
3.4.4. Disentanglement Loss
We can obtain the final disentanglement loss $\mathcal{L}_{dis}$ by gathering the losses of all perspectives:
(14) $\begin{split}&\mathcal{L}_{dis}=\mathcal{L}_{dif}+\mathcal{L}_{con}+\mathcal{L}_{ort}\end{split}$
3.5. Training
Our task is to forecast the future increase in citations for the $i$-th paper. Considering that the distribution of citation count is skew distribution, we apply logarithmization to normalize the distribution. Thus, the main loss function $\mathcal{L}_{reg}$ of the potential impact prediction task is the mean squared error (MSE) between the logarithmized target and predicted citation increments: $y_{i}$ and $\hat{y}_{i}^{all}$.
(15) $\begin{split}&\hat{y}_{i}^{all}=\hat{y}_{i}^{dif}+\hat{y}_{i}^{con}+\hat{y}_{i}\\ &\mathcal{L}_{reg}=\frac{1}{N}\sum_{i=1}^{N}{(y_{i}-\hat{y}_{i}^{all})^{2}}\end{split}$
The final loss function $\mathcal{L}$ includes the main regression loss (MSE) and the disentangled learning loss:
(16) $\begin{split}&\mathcal{L}=\mathcal{L}_{reg}+\beta\mathcal{L}_{dis}\end{split}$
where $\beta$ controls the strength of disentanglement tasks.
4. Experimental Settings
In this section, we conduct extensive experiments and analyses on three subsets extracted from the real-world scientific dataset S2AG to validate DPPDCC  [^1] We will make our datasets and codes publicly available on Github.. Particularly, we plan to answer the following research questions (RQs):
$\bullet$ RQ1: Can DPPDCC improve the potential impact prediction task?
$\bullet$ RQ2: What is the role of each component in DPPDCC?
$\bullet$ RQ3: How sensitive is DPPDCC to hyper-parameters?
$\bullet$ RQ4: How does DPPDCC conduct the prediction?
4.1. Datasets
Our dataset is constructed using S2AG (Kinney et al., 2023), a comprehensive repository housing roughly 100 million scientific publications, offering extensive metadata and unique identifiers. To enhance experimental efficiency and accommodate the diverse characteristics across domains, we organize these publications into subsets based on their respective fields of study. To establish the global citation network, we locate English papers with complete metadata (title, abstract, author IDs, and venue ID), or those of high impact despite lacking venue information. From the dataset, we select three fields: computer science (CS), chemistry (CHM), and psychology (PSY). Relevant statistics are summarized in Table 1.
Table 1. Detailed statistics for ground-truth samples, nodes and edges of the citation network, as well as sample information of training, validation, and testing sets.
dataset CS CHM PSY
ground-truth sample (paper) 2,513,197 1,818,138 1,857,277
node paper 1,628,853 1,376,599 1,297,771
author 1,598,925 1,946,073 1,585,595
venue 12,524 8,389 13,775
time 147 187 202
edge cite 11,534,431 10,382,698 13,401,112
write 5,123,460 6,355,630 4,813,135
publish 1,566,442 1,318,158 1,230,440
have 1,628,853 1,376,599 1,297,771
sample train (2012) 183,105 227,304 219,499
val (2014) 224,086 256,403 250,229
test (2017) 300,000 300,000 300,000
qualified pool 1,112,611 1,093,660 1,015,579
To select samples for testing, we target papers published before the test observation point (see Figure 1 (b)). These papers should be written in English, have complete metadata, and include at least one reference, without imposing constraints on the number of citations received, to maintain the integrity of the data distribution. From the eligible pool, we randomly sampled 300,000 papers to compose the test set. Additionally, we gather all the associated metadata for these papers, encompassing titles, abstracts, authors, venues, and publication times. To imitate the practical scenario, the training observation point is set 5 years before the test observation point (equal to the observation time window $T$ and predicted interval $\Delta$) for non-overlapping. The validation observation point precedes the test observation point by 3 years. It is crucial to underscore that though predicted samples in the training set reappear in the test set, their input graphs and predicted values undergo significant variations due to dynamic contextual changes. This intrinsic challenge renders the task both complex and critical. Moreover, we meticulously categorize the papers in the results for thorough and detailed analysis.
4.2. Baseline Models
The baseline models are mainly divided into citation/cascade prediction models, dynamic GNN models, and disentangled GNN models:
Citation/Cascade prediction models: We apply classic and recent citation/cascade prediction models to estimate the potential impacts of papers based on content, citation graphs, or citation cascade graphs.
    SciBERT (Beltagy et al., 2019) is a BERT-based model pretrained on scientific corpus. We fine-tune it to encode paper titles and abstracts, predicting with the [CLS] token for citation prediction.
    HINTS (Jiang et al., 2021) tackles the cold-start problem by employing R-GCN on pseudo meta-data heterogeneous graphs. It then feeds the node (paper) representations into GRU for sequential modeling and employs a stochastic model for predictions. To integrate content information from PLMs, we replace node ID embeddings and adjust the model to encode subgraphs instead of the entire graph, mitigating memory issues.
    MUCas (Chen et al., 2022) first uses time-aware sampling to extract dynamic graph snapshots and generate node embeddings from the graph structure. Then, it utilizes MUG-caps, a hierarchical capsule network, to encode snapshots in different levels (order, node, and graph) and aggregate the final representations with the attention mechanism.
Dynamic GNN models: We employ recent dynamic GNN models to extract paper representations from the dynamic graphs collected the same as our model, taking the potential impact prediction as a graph regression task. Similar to HINTS, we adapt these models to concentrate on target-centric subgraphs rather than the entire graph.
    EvolveGCN (Pareja et al., 2020) integrates RNN structure within GNN modules. It regards graph structures as hidden states within the GRU. It adopts GRU parameters as GCN parameters.
    Dysat (Sankar et al., 2020) incorporates multiple self-attention modules to encode structural and temporal information. It first applies a shared GAT to encode subgraphs separately across different years. Subsequently, a vanilla transformer encoder models the temporal information of the graph representation sequence.
    ROLAND (You et al., 2022) is a novel dynamic heterogeneous GCN model that inherits static GCN methods. It retains the hierarchical information of GCNs’ different layers and then applies a GRU-like updater to iteratively update representations.
Disentangled GNN models: We apply recent disentangled GNN models to extract critical representations from the multi-hop subgraphs.
    DisenGCN (Ma et al., 2019a) partitions the feature vector into K channels and uses neighborhood routing from capsule networks for disentanglement.
    DisenHAN (Wang et al., 2020) adapts DisenGCN to heterogeneous graphs for recommendation. It employs DisenGCNs in different homogeneous graphs for intra-aggregation. It then employs the attention mechanism to aggregate information from different relationships under various channels.
    CAL (Sui et al., 2022) employs attention modules to estimate the causal and trivial masks for structures and attributes. It encodes them with specific GCNs separately and applies causal intervention by randomly swapping the trivial part of the whole embeddings.
    DisC (Fan et al., 2022) first estimates the causal/bias masks and then employs causal/bias-aware loss functions. It further generates counterfactual unbiased samples in the embedding space by swapping the biased part of the representations.
    DIDA (Zhang et al., 2022) discovers invariant patterns within dynamic graphs by generating soft masks using the self-attention mechanism. It then applies random variant pattern swapping in both temporal and spatial dimensions as the causal intervention.
4.3. Implementation Details
All baselines and our proposed model are implemented with PyTorch and DGL/PyG. They are trained on an NVIDIA A800 80GB GPU and optimized with the Adam optimizer (Kingma and Ba, 2014). As for the baselines, we implement them directly using or referring to the official source codes. We develop the models with the training and validation sets and choose the final model with the best main metric on the validation set. Each dataset is divided into previous papers (included in the training set), fresh papers (new additions in the test set), and immediate papers (published at the test time point) for detailed analysis. We adhere to the recommended settings for all baselines. In our model across all datasets, the hop order $k$ is set to 2, corresponding top limits $K_{1}$ and $K_{2}$ are set to 100 and 20, the learning rate is set to 1e-4, the number of layers $L$ is set to 4, the co-cited/citing weight $\lambda$ is set to 0.5, and the categories of equal frequency bins $M$ are set to 5. Additionally, the batch size is set to 64 for CS and CHM, and 32 for PSY, the disentanglement weight $\beta$ is set to 0.5 for CS and PSY, and 1.0 for CHM, and the hidden dimension $d^{h}$ is set to 128 for CS and PSY, and 192 for CHM.
4.4. Evaluation Metric
[…]
(17) $MALE=\frac{1}{N}\sum_{i=1}^{N}{|y_{i}-\hat{y}_{i}|}$
[…]
(18) $R^{2}=1-\frac{\sum_{i=1}^{N}{(y_{i}-\hat{y}_{i})^{2}}}{\sum_{i=1}^{N}{(y_{i}-\frac{1}{N}\sum_{i=1}^{N}y_{i})^{2}}}$
[…]
5. Results
5.1. Performance Comparison (RQ1)
To answer RQ1, we evaluate the performance of DPPDCC and other baselines. The comparison results are reported in Table 2 (Computer Science), 3 (Chemistry), and 4 (Psychology). We have the following observations:
[…]
CS model all prev fresh immed
[…]
C&C SciBERT 0.6847 0.3558 0.6077 0.4424 0.8054 0.1108 0.7860 0.0843
HINTS 0.6799 0.2773 0.6429 0.2028 0.7378 0.2268 0.7387 0.1718
MUCas 0.5079 0.6063 0.4331 0.6757 0.6250 0.4352 0.8261 -0.1079
Dynamic EGCN 0.8157 0.1174 0.8008 0.0531 0.8392 0.0210 0.8387 -0.1460
Dysat 0.6248 0.4355 0.5489 0.5029 0.7437 0.2321 0.8073 0.0089
ROLAND 0.5814 0.5245 0.5166 0.5634 0.6828 0.3764 0.7157 0.2511
Disen DisenGCN 0.8215 0.1428 0.7816 0.1022 0.8841 0.0206 0.8485 0.0236
DisenHAN 0.8138 0.0054 0.7411 0.1003 0.9276 -0.3220 1.0686 -0.8032
CAL 0.7323 0.2298 0.6859 0.2264 0.8050 0.0768 0.7940 0.0427
DisC 0.6687 0.3434 0.6144 0.3908 0.7536 0.1472 0.7900 -0.0735
DIDA 0.5941 0.4703 0.5018 0.5642 0.7387 0.2391 0.8181 -0.0207
ours DPPDCC 0.4473 0.6871 0.3783 0.7236 0.5553 0.5754 0.6035 0.4512
#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
[…]
CHM model all prev fresh immed
[…]
C&C SciBERT 0.5451 0.4648 0.4987 0.4709 0.6903 0.1477 0.6452 0.1653
HINTS 0.6024 0.3795 0.5848 0.3119 0.6574 0.2253 0.6449 0.1633
MUCas 0.4691 0.6168 0.4314 0.6306 0.5869 0.3624 0.8073 -0.3479
Dynamic EGCN 0.6424 0.2684 0.6177 0.1913 0.7196 0.0791 0.6976 0.0035
Dysat 0.5442 0.4573 0.5029 0.4411 0.6733 0.1998 0.6846 0.0582
ROLAND 0.5001 0.5410 0.4592 0.5374 0.6280 0.2942 0.6406 0.1800
Disen DisenGCN 0.6405 0.2735 0.6079 0.2151 0.7427 0.0332 0.6987 0.0387
DisenHAN 0.6132 0.3793 0.5817 0.3683 0.7118 0.0630 0.7313 -0.1466
CAL 0.5864 0.3928 0.5541 0.3707 0.6873 0.1158 0.6535 0.1135
DisC 0.5393 0.4770 0.5037 0.4578 0.6505 0.2387 0.6409 0.1719
DIDA 0.4978 0.5760 0.4642 0.5689 0.6028 0.3587 0.6426 0.1669
ours DPPDCC 0.4355 0.6503 0.4071 0.6329 0.5243 0.5042 0.5802 0.3247
#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
[…]
PSY model all prev fresh immed
[…]
C&C SciBERT 0.6395 0.4308 0.6025 0.5041 0.7402 0.0068 0.7116 -0.1095
HINTS 0.7749 0.1988 0.7799 0.2007 0.7614 -0.0259 0.7064 -0.0804
MUCas 0.5207 0.6175 0.4804 0.6846 0.6306 0.2669 0.7708 -0.3521
Dynamic EGCN 0.7424 0.2305 0.7375 0.2469 0.7560 -0.0388 0.7102 -0.1386
Dysat 0.6861 0.2998 0.6607 0.3342 0.7555 -0.0170 0.7439 -0.2377
ROLAND 0.5643 0.5540 0.4996 0.6601 0.7406 0.0429 0.7926 -0.3765
Disen DisenGCN 0.6557 0.3945 0.6316 0.4364 0.7213 0.0758 0.7081 -0.0846
DisenHAN 0.7763 0.2455 0.8077 0.2187 0.6906 0.1387 0.6476 0.0522
CAL 0.5631 0.5646 0.5469 0.5975 0.6073 0.3256 0.6285 0.1166
DisC 0.5415 0.5915 0.5159 0.6343 0.6113 0.3233 0.6369 0.0963
DIDA 0.5878 0.5118 0.5604 0.5620 0.6624 0.1950 0.6681 -0.0092
ours DPPDCC 0.4463 0.7037 0.4138 0.7454 0.5349 0.4701 0.5633 0.2733
#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
[…]
(2) Both dynamic and heterogeneous graphs are important for our task. On one side, since temporal information within dynamic graphs exposes the evolved trend, it plays a crucial role. Dynamic GNNs possess comparable performance across all datasets for almost all categories of papers. In particular, DIDA, which considers dynamic graphs, stands out among the disentangled GNNs, closely matching MUCas’ performance across all datasets. On the other side, leveraging heterogeneous information can further enhance the performance of dynamic GNNs, as demonstrated by ROLAND. This general dynamic GNN framework achieves competitive performance akin to MUCas by integrating heterogeneous information. However, proper modeling of heterogeneous information is crucial; otherwise, it might lead to performance degradation. For instance, DisenHAN, compared to its backbone DisenGCN, struggles to encode valuable information within heterogeneous graphs, notably within the complex dataset Psychology.
[…]
5.2. Ablation Test (RQ2)
To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules. As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2021), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering. For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP for prediction.
Table 5. Results of ablation studies according to MALE.
model CS CHM PSY
ours 0.4473 0.4355 0.4463
CGE -co-cited/citing 0.4650 0.5014 0.4528
-CompGAT 0.4722 0.4990 0.4582
-readout 0.5116 0.4404 0.4932
PDM -diffusion 0.4593 0.4530 0.4573
-conformity 0.5177 0.4747 0.4580
-orthog 0.4930 0.4388 0.4502
-disen 0.4606 0.4447 0.4670
We observe that: (1) All modules within DPPDCC play pivotal roles, as evidenced by the decrease in performance when any module is removed. Their significance varies across datasets due to diverse graph characteristics. (2) For CGE, the removal of type-specific attention readout may cause a notable performance drop. That could be attributed to the presence of excessive noisy information within multi-hop snapshots, which necessitates the filtering of aggregated information for better concentration. Specially, in chemistry, where subgraphs are sparser, simple avg-pooling might suffice to extract meaningful information. Furthermore, our proposed CompGAT effectively captures distinct comparative information, surpassing the original GATv2 in computer science and psychology. The incorporation of co-cited/citing strengths further enhances multi-hop information beyond the local neighborhood, resulting in improved performance. (3) Eliminating any disentanglement task adversely affects performance. Moreover, an evident decline is observed after removing the whole disentanglement module, especially in Psychology. Conformity rises as the paramount task among all disentanglement tasks. It focuses on the Matthew effect, pervasive and discernible in reality, thus significantly contributing to model performance.
5.3. Hyper-parameters Test (RQ3)
(a) $L$
(b) $\lambda$
(c) $\beta$
Figure 3. Results of hyper-parameters test in terms of MALE.
(a) $L$
(b) $\lambda$
(c) $\beta$
[…]
To address RQ3, we conduct experiments to assess the influence of three critical hyper-parameters in our model design: the stacked layer number $L$, the importance of co-cited/citing strengths $\lambda$, and the weight of the disentanglement tasks $\beta$. To ensure robust experimental outcomes, we focus on tuning these specific hyper-parameters while keeping other hyper-parameters fixed at their optimal values. In the following section, we report the corresponding results when tuning other important hyper-parameters. The experiment results are described in Figure 3 and 4.
5.3.1. The Impact of Model Depth
The number of layers $L$ serves as a pivotal hyperparameter for GNN-based models. Within this study, we explore values of $L$ in the range of 2, 3, 4, 5. For GNN-based models, increasing the number of layers facilitates the capture of higher-order connectivity within the graph during information propagation and aggregation. However, it also introduces the challenge of over-smoothing, a trend we’ve observed. As the layer count $L$ increases, model performance typically exhibits an initial rise followed by a decline—a phenomenon attributed to the interplay between increased connectivity capturing and eventual over-smoothing effects. In DPPDCC, besides GNN’s inherent challenges, the incorporated multi-hop co-cited/citing information and the intricate interaction between time and space dimensions could contribute to this phenomenon. Achieving the optimal hyperparameter $L$ requires finding a delicate balance between the GNN’s depth and its multi-hop/multi-dimensional information, ensuring an equilibrium that maximizes performance.
5.3.2. The Impact of Co-cited/citing Strengths
The weight parameter of co-cited/citing strengths $\lambda$ governs the proportion of co-cited/citing strengths in the sum distributions relative to the original attention scores. Our study investigates optimal values of $\lambda$ within the range 0.25, 0.5, 0.75, 1.0. Setting $\lambda$ to 1.0 essentially disregards attention mechanisms, replacing them with fixed weights and reducing the model to a weighted GCN. Higher weights can potentially dominate significant signals learned by the model, while lower weights may inadequately incorporate additional information. Hence, a moderate value such as 0.5 might represent a suitable setting adaptable across various datasets, balancing the incorporation of additional information without overshadowing the model’s learned signals.
5.3.3. The Impact of Disentanglement Loss
The weight of disentanglement loss $\beta$ controls the importance of the disentanglement tasks relative to the main task. Our study explores optimal values of $\beta$ within the range 0.25, 0.5, 0.75, 1.0. Notably, the disentanglement weight $\beta$ exhibits greater stability in comparison to $L$ and $\lambda$, indicating consistent influence and effectiveness of the disentanglement task as an auxiliary task. Selecting an appropriate $\beta$ value can further enhance overall model performance, underscoring the significance of this parameter in refining the model’s effectiveness.
5.4. Visualization (RQ4)
To answer RQ4, we visualize both CGE and PDM in the test set of Chemistry to delve deeper into our model’s mechanisms. For CGE, we employ linear regressions to uncover relationships among the sum distribution $\alpha$, original attention distribution $e$, and co-cited/citing distribution $c$. This analysis aims to elucidate the interplay and dependencies among these distributions. For PDM, we first calculate correlations between predicted/ground-truth values and the perspective values/proportions. Next, we visualize the comprehensive compositions of disentangled values concerning time, category, and value. Subsequently, we delve into the representation distribution of conformity perspective to validate its rationality. This process offers insights into how different perspectives contribute to predictions.
5.4.1. Weighting Score Visualization
(6.48/16.74)
c
(17.60/17.05)
(40.76/73.79)
(299.40/152.78)
(1.52/4.35)
(e) $\alpha^{(1)}=f(e^{(1)})$
(4.15/4.36)
(f) $\alpha^{(2)}=f(e^{(2)})$
(9.28/18.01)
(g) $\alpha^{(3)}=f(e^{(3)})$
(58.70/36.62)
(h) $\alpha^{(4)}=f(e^{(4)})$
(1.67/4.17)
(i) $\alpha^{(1)}=f(c)$
(4.53/4.28)
(j) $\alpha^{(2)}=f(c)$
(10.67/19.02)
(k) $\alpha^{(3)}=f(c)$
(85.97/40.87)
(l) $\alpha^{(4)}=f(c)$
Figure 5. The outcomes of the linear regression analysis concerning the ”cites”/”is cited by” scores within the field of chemistry. Below the respective sub-figures, we depict the mean Kullback-Leibler (KL) divergences between the target and source distributions, denoted in units of $10^{-4}$.
To effectively visualize the connections between distributions, we gather all scores subsequent to applying softmax based on the target node. These scores are flattened to observe the overarching trend within the latest year. The original attention scores $e$ are averaged across all heads for ease of interpretation. Notably, for the negative effect in the co-citing relationship, we employ reversed values via $1-c$, rendering them positive to reflect the trend. In Figure 5, robustly positive relationships are observed among the sum distribution $\alpha$, original attention distribution $e$, and the distribution of co-cited/citing strengths $c$. It is reasonable to note that the sum distribution closely resembles either of the sourced distributions. When compared to the co-cited/citing distribution, the sum distribution demonstrates greater similarity to the original attention distribution. Intriguingly, within the initial layers of the model, minimal disparities are evident between the original attention scores and the co-cited/citing strengths. This phenomenon is attributed to both the averaging across all heads and the model’s adaptation to the co-cited/citing information during the learning process. Moreover, while the overall trend is evident, a finer examination of the samples reveals that for instances with varying co-cited/citing strengths, their original attention scores tend to be equal, resulting in the displayed horizontal and vertical distribution evident in Figures 5a and 5e. This equal might arise due to the initial state of the attention mechanism in the Graph Attention Network (GAT). However, as the model progresses in depth, noticeable disparities emerge between these distributions in both ”cites” and ”is cited by” edges. In the upper layers, attention scores exhibit a tendency to concentrate on a select few nodes, while in the lower layers, they disperse attention across a wider array of nodes. Co-cited/citing strengths represent 2-hop information conveyed through intermediate nodes, resembling lower-hop information within GNNs. Therefore, continuously incorporating the co-cited/citing distribution across all layers reinforces signals originating from pivotal nodes across a broader spectrum. It can alleviate the over-smoothing issues by preventing the dominance of a select few nodes within aggregated information. This approach serves to explain why a layer count of 4 emerges as optimal—it encapsulates higher-order information beyond the 2-hop range, allowing the co-cited/citing information to complement the original attention scores. This collaboration achieves multi-hop information integration, mitigating both forgetting and overemphasis within the learning process.
5.4.2. Disentanglement Correlation Analysis
Table 6. Results of Disentanglement Correlation in Chemistry
dif con contri dif (%) con (%) contri (%) citations pred real
dif 0.9463 0.9675 0.6177 -0.5847 0.3845 0.2159 0.9947 0.7804
con 0.9463 0.8787 0.5178 -0.3892 0.1889 0.2765 0.9479 0.7501
contri 0.9675 0.8787 0.5182 -0.6342 0.5124 0.1909 0.9828 0.7790
dif (%) 0.6177 0.5178 0.5182 -0.6870 0.2795 0.0651 0.5641 0.3965
con (%) -0.5847 -0.3892 -0.6342 -0.6870 -0.8897 -0.0157 -0.5760 -0.4305
contri (%) 0.3845 0.1889 0.5124 0.2795 -0.8897 -0.0201 0.4067 0.3197
citations 0.2159 0.2765 0.1909 0.0651 -0.0157 -0.0201 0.2226 0.2025
pred 0.9947 0.9479 0.9828 0.5641 -0.5760 0.4067 0.2226 0.7895
real 0.7804 0.7501 0.7790 0.3965 -0.4305 0.3197 0.2025 0.7895
In Table 6, we depict the correlations among all disentangled perspective values/proportions and the predicted/real values, focusing on samples with all positively predicted values within the Chemistry test set. For papers containing negative disentangled values, they always have nearly no increased citations. Remarkably, over 80% of them feature negative because their contribution values are the lowest across all perspectives, demonstrating the rationality of our proposed method. Among the correlations observed among the values of all three perspectives, ”conformity” stands out as particularly distinct. Its correlation with ”contribution” is notably the lowest, falling below 0.9. Additionally, its correlation with the predicted/real values also demonstrates a significant gap (-0.05/-0.03) when compared to other perspectives. Despite this, it emerges as the most closely associated perspective with accumulated citations, implying its significance in representing information linked to a paper’s reputation. The predicted increment appears as the most influential predictor of the real increment. Its aggregated value outperforms individual perspectives, highlighting the effectiveness of the proposed disentangled learning approach. Moreover, the real increment demonstrates minimal correlation with the paper’s accumulated citations. This revelation underscores two key points: (1) Citation increment proves to be an effective metric in portraying a paper’s potential impact, with accumulated citation accounting for only a fraction of the future increment. (2) The DPPDCC model adeptly captures the trend in a paper’s citation increment, validating its substantial potential and reliability. In our calculations of the proportions of all three perspectives within predicted values, a notable finding emerges: conformity diffusion stands out as the only perspective with numerous negative correlations. This observation suggests that in the field of Chemistry, papers solely characterized by accumulated citations tend to attract fewer new citations. Moreover, it shows an almost negative correlation with the proportion of contribution (-0.89), indicating their mutual exclusivity. Consequently, conformity and contribution have learned entirely distinct facets of citation increments. Overall, the diffusion perspective can be seen as the most pivotal factor for prediction, as it showcases the highest correlations in both values and proportions. This prominence aligns with the intuitive connection between citation increments and the information diffusion process. Conversely, the secondary perspective, contribution, likely delves into factors beyond the graph structure, aiming to uncover additional crucial elements. As a complement to the numerical citation increment, it offers a more objective evaluation, potentially aiding in the identification of valuable papers by encompassing diverse and complementary information.
5.4.3. Disentangled Composition Visualization
(a) time trend
(b) group detail
(c) value bin
Figure 6. Visualization of disentangled value proportions in Chemistry. (a) displays the trend that evolved with the publication time. (b) demonstrates the detailed composition of papers categorized into previous, fresh, and immediate ones. (c) is binning the samples based on the predicted values.
We visualize the disentanglement composition with all positive values in the test set of Chemistry. For the stacked plot associated with time in Figure 6a, a clear trend emerges: the contribution proportion gradually increases, while the conformity proportion undergoes a significant drop. This trend is sensible since old papers tend to accumulate substantial citations, thus attracting more researchers to cite them due to inertia. Additionally, the diffusion proportion appears to be more stable across different times, as it represents the potential hidden properties within diffusion. This stability becomes more pronounced when papers are categorized into previous, fresh, and immediate ones in Figure 6b. Diffusion proportions remain steady in both mean and standard deviation. Moreover, contribution emerges as the most critical perspective, with its dominant proportion. It further holds particular significance for fresh and immediate papers. For immediate papers receiving almost no citations, both the diffusion and conformity perspectives struggle to offer predictive insights. However, the contribution enables the model to anticipate the potential of samples in advance by analyzing their citation context. This capability enhances the model’s robustness, particularly for challenging samples close to cold-start scenarios. Furthermore, in Figure 6c, when we bin the samples based on the predicted logarithmic values, we observe a descending trend for conformity and a growing trend for contribution. Notably, within the lowest bin group, the conformity proportion surpasses the contribution proportion for the first time. The alignment between time and impact groups can unveil an overall trend: in dynamic contexts, previous papers tend to attract fewer new citations and may primarily owe to their conformity, whereas recent and immediate papers are more likely to garner increased citations owing to their genuine contributions.
5.4.4. Disentangled Conformity Representation Visualization
(a) total
(b) prev
(c) fresh
(d) immed
Figure 7. Results of T-SNE visualization for representations from conformity perspective in Chemistry. Various colors correspond to the binning labels assigned to the samples.
To validate that disentangled perspectives learn distinct aspects of the features, we visualize the representations extracted from the conformity encoder with T-SNE employed by scikit-learn (Pedregosa et al., 2011), as this perspective is the only one with available labels. Initially, the T-SNE algorithm is executed on the complete test sets of each category. Subsequently, 3000 points are randomly sampled to generate the figures. The points are color-coded based on the binning labels of the cumulative citations at the test time point. Figure 7 illustrates a consistent and structured distribution of points across all figures, indicating our model’s effective capture of group discrepancies across various binning labels in all categories. Across different categories, variations in the distribution of binning labels are evident due to differences in cumulative citations over time. Fresh Papers published within the last five years, particularly immediate ones, tend to fall under the lowest accumulated citation label (red), dominating a significant area. Nevertheless, discernible boundaries persist between distinct binning labels within these categories. This suggests the model’s capacity to generalize and transfer learned group differences to unseen samples, showcasing its potential applicability in real-world scenarios.
5.5. Case Study (RQ4)
Table 7. Detailed meta-data, disentangled values/proportions, and potential impact of the selected cases.
id meta disentangled values disentangled proportions potential impact
pub time refs citations dif con contri dif con contri pred real
242481 2012 28 43 1.2272 0.8977 1.4735 34.10% 24.95% 40.95% 3.5984 3.0445
15676 2001 19 103 1.2151 0.9443 1.4569 33.60% 26.11% 40.29% 3.6163 3.0445
374215 2015 60 26 1.1949 0.8409 1.6318 32.58% 22.93% 44.49% 3.6675 3.2189
(a) 242481-$c$
(b) 242481-$e$
(c) 242481-$\alpha$
(d) 15676-$c$
(e) 15676-$e$
(f) 15676-$\alpha$
(g) 374215-$c$
(h) 374215-$e$
(i) 374215-$\alpha$
Figure 8. Weighted paper ”cites” subgraphs of selected cases organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ in Chemistry. we filter the edges within 2-hop graphs by retaining only the top 30% accumulated weights of each node and preserve the direct citation or reference nodes of the target paper to attain these sparser graphs. We only label the target paper and its directly linked highly-cited citations.
To delve into the intricate workings of our proposed model in addressing RQ4, we select three representative papers for each perspective. These papers should be co-cited or citing papers with similar topics, reflecting comparable predicted and real potential impact. Our approach involves contrasting their metadata, disentangled values, perspective proportions, and predicted/real potential impact. In this section, we visualize graph structures organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ within ”cites” edge subgraphs of the selected papers. To maintain clarity, we focus on graphs extracted from the latest year of the snapshots and the final layer of DPPDCC. We then filter the edges in 2-hop graphs based on the top 30% accumulated weights, resulting in sparser graphs. To create target-centric connected graphs, we retain only the direct citations and references of the target paper. Highly-cited papers, representing papers that provide the top 80% of citations within their respective fields of the whole datasets, are highlighted by filling in red within the graphs. Our labeling strategy encompasses the target paper and its directly connected highly-cited citations.
As outlined in Table 7, 242481, 15676, and 374215 respectively embody the diffusion, conformity, and contribution perspectives. Despite their near-identical predicted and real increments, their disentangled value compositions differ from one another. Notably, 374215 stands as a recent paper not encountered during the training phase, while 242481 and 15673 are earlier publications, each with varying years since publication and distinct accumulated citations. These papers exhibit perspective proportions aligning with all trends discussed in Section 5.4.3. As depicted in Figure 8, $c$ subgraphs manifest as the densest, while $e$ and $\alpha$ subgraphs appear considerably sparser, featuring fewer interconnected nodes. In the higher layers of DPPDCC, the model tends to concentrate on specific nodes, resulting in sparsity in these graphs. The introduced co-cited strengths $c$ serve as a complementary factor to the original attention scores $e$, aiding in identifying valuable citations from highly-cited papers. Notably, $c$ consistently incorporates highly-cited papers overlooked in $e$ subgraphs. For instance, Figures 8b and 8e consider only one highly-cited paper each (532725 and 91549). However, in Figures 8c and 8f, three additional highly-cited papers (559976, 532726, 180237 and 26005, 582958, 193949) - initially presented in Figures 8a and 8d - are taken into account.
For 242481, within the test set, it stands as a recent publication with limited citations due to its relatively short existence. Despite its lower accumulated citations compared to 15676, 242481 has garnered significant attention from four highly-cited papers, equaling the number received by 15676. Consequently, its average influential citation surpasses that of 15676, leading to a higher proportion of the diffusion perspective in its incremental source. Furthermore, since subgraphs depicted in Figure 8 are based on the ”cites” edge, only edges from valuable citations are retained. In contrast to 15676, 242481 boasts numerous connections from references, indicating its significance as an important citation from the perspective of its references. Consequently, it emerges as a pivotal element in the information diffusion process, contributing to its high diffusion values. For 15767, being the oldest paper with a publication span of over 10 years by the test time point (2017), its conformity value stands out as the highest among all cases, contributing significantly to a larger proportion of its incremental impact. Despite its extensive citation count, considering the evolving context of the citation network, its significance has diminished. Over time, citations toward this paper tend to lean more towards its established reputation rather than other factors. For 374215, it is the most recent paper with limited citations. Because it has neither highly-cited citations nor enough accumulated citations, compared to either 242481 or 15767, both its diffusion and conformity perspectives are low in values and proportions. Conversely, its contribution perspective appears as the most pivotal, showcasing the largest values and proportion. Benefiting from its contribution, 374215 successfully surpasses the other two papers in predicted/real potential impact. Interestingly, beyond the factors highlighted in the graph structures or accumulated citation counts, something intrinsic to its contribution stands out, as delineated by DPPDCC. This factor distinguishes itself from mere popularity, emphasizing DPPDCC’s substantial potential for practical application and impact.
Indeed, our proposed DPPDCC effectively disentangles citation increments derived from various factors for the target paper. This disentanglement, reflected in the values and proportions assigned to distinct perspectives, provides more comparable metrics for identifying valuable papers. In future work, we aim to extend this framework and assess its practical application in uncovering novel insights within real-world contexts. This ongoing exploration will enable us to delve deeper into new discoveries and applications facilitated by our disentanglement approach.
6. Conclusion
In this study, we propose a novel model named DPPDCC that disentangles the potential impacts of papers into diffusion, conformity, and contribution values. To better align with real-world conditions, we reformulate the problem by partitioning the datasets based on the observation time point, without excluding lowly-cited papers. Further, DPPDCC comprises two principal components: Citation-aware GNN Encoder (CGE) and Popularity-aware Disentanglement Module (PDM). Given the dynamic heterogeneous graph of a target paper, CGE employs CompGAT to encode the comparative and co-cited/citing information between papers and utilizes Type-specific Snapshot Readout to aggregate the information from each snapshot in an evolutionary manner. Subsequently, PDM identifies amplifying effects in information diffusion and addresses the Matthew effect stemming from accumulated citations from a collective conformity perspective. Finally, it constrains each perspective to address distinct aspects for preserving the genuine contribution. Experimental results on three datasets demonstrate that DPPDCC significantly outperforms alternative approaches for previous, fresh, and immediate papers. Disentanglement visualizations and a detailed case study provide additional evidence supporting DPPDCC’s reliable predictive capabilities.
In our future endeavors, we intend to develop extensive datasets spanning diverse fields of study to simulate practical scenarios. This initiative is geared towards significantly enhancing the practicality and relevance of our approach. Additionally, we plan to incorporate additional factors that account for citation intent, thereby enabling a more precise evaluation of the genuine contributions made by academic papers. Furthermore, we will conduct a thorough analysis of the model’s performance in identifying novel papers within specific domains. Importantly, we’ll validate these findings by comparing results with relevant studies, elucidating the practical implications of our proposed methodology.

### v3

@@
 For example, John et al.
-(John et al., 2018) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style.
+(John et al., 2019) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style.
 This separation allows precise control over content and style in text generation.
@@
 Leveraging these co-cited/citing strengths can help uncover the dynamics inherent in multi-hop information diffusion.
-Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2021), to encode comparative and co-cited/citing features between paper nodes.
+Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2022), to encode comparative and co-cited/citing features between paper nodes.
 To capture the discrepancy between source and target nodes, we encode the concatenation of their representations.
@@
 where $c_{ij}\in\bm{C}$ is the normalized co-cited/citing strength, $\lambda$ is to balance the sum distribution of co-cited/citing strengths and basic attention weights, and $\bm{w}_{a}^{(l)}$is the attention vector.
-For other edges in the R-GCN, we apply GIN (Xu et al., 2018) with sum-pooling as the GNN encoder:
+For other edges in the R-GCN, we apply GIN (Xu et al., 2019) with sum-pooling as the GNN encoder:
 (5) $\begin{split}\bm{h}_{i,r_{dst}}^{(l+1)}=f_{\Theta,r}((1+\xi)\bm{h}_{i,r_{dst}}^{(l)}+AGG(\{\bm{h}_{j,r_{src}}^{(l)},&j\in\mathcal{N}(i,r_{src})\})),r\in\mathcal{R}^{normal}\end{split}$
@@
 4.4. Evaluation Metric
-[…]
+In our evaluation, we utilize two metrics: MALE (Mean Absolute Logarithmic Error) and LogR2 (Logarithmic R-squared).
+MALE calculates the absolute error between the target and predicted values post-logarithm transformation, reflecting the direct predictive capability:
 (17) $MALE=\frac{1}{N}\sum_{i=1}^{N}{|y_{i}-\hat{y}_{i}|}$
-[…]
+Regarding LogR2, it evaluates the coefficient of determination, calculated as the proportion of the dependent variable’s variation predictable from the independent variable, following logarithm transformation for both variables:
 (18) $R^{2}=1-\frac{\sum_{i=1}^{N}{(y_{i}-\hat{y}_{i})^{2}}}{\sum_{i=1}^{N}{(y_{i}-\frac{1}{N}\sum_{i=1}^{N}y_{i})^{2}}}$
-[…]
+LogR2 assumes a negative value when the model’s predictions fall short of the data’s mean value, indicating an inability to capture the underlying trend within the dataset.
+In our problem context, mirroring practical scenarios within dynamic contexts that require extrapolative abilities, encountering negative values is both frequent and expected due to inherent complexities.
+Consequently, by integrating MALE and LogR2, we holistically gauge the model’s accuracy in direct predictions for individual samples and its aptness in capturing broader trends across groups.
 5. Results
@@
 We have the following observations:
-[…]
+Table 2.
+Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Computer Science.
+We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total.
+The best results are shown in bold, and the second results are underlined.
+Significant improvements over best baseline results are marked with .
 CS model all prev fresh immed
@@
 ours DPPDCC 0.4473 0.6871 0.3783 0.7236 0.5553 0.5754 0.6035 0.4512
-#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
-[…]
+#improve (%) 11.93 13.32 12.65 7.09 11.15 32.22 15.68 79.67
+Table 3.
+Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Chemistry.
+We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total.
+The best results are shown in bold, and the second results are underlined.
+Significant improvements over best baseline results are marked with .
 CHM model all prev fresh immed
@@
 ours DPPDCC 0.4355 0.6503 0.4071 0.6329 0.5243 0.5042 0.5802 0.3247
-#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
-[…]
+#improve (%) 7.16 5.44 5.64 0.36 10.67 39.13 9.42 80.39
+Table 4.
+Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Psychology.
+We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total.
+The best results are shown in bold, and the second results are underlined.
+Significant improvements over best baseline results are marked with .
 PSY model all prev fresh immed
@@
 ours DPPDCC 0.4463 0.7037 0.4138 0.7454 0.5349 0.4701 0.5633 0.2733
-#improve (%) $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$ $*{}^{*}$
-[…]
+#improve (%) 14.29 13.96 13.86 8.87 11.92 44.38 10.38 134.40
+(1) All models perform well on previous papers presented in the training set but experience degradation when dealing with fresh and immediate papers in the test set, particularly in terms of the LogR2.
+Many models fail to figure out the overall trend of immediate papers, which results in negative values in LogR2.
+Even the best baseline, MUCas, struggles with immediate papers due to limited graph structures, as it is a cascade prediction model designed for papers with sufficient citations.
+However, dynamic GNN models can effectively encode multi-hop information to complement the lack of cascade structures.
+Disentangled GNNs, in particular, effectively uncover crucial insights within both multi-hop graph structures and node representations, notably beneficial for datasets like Psychology, which are denser.
+Moreover, models such as SciBERT and HINTS leverage prior knowledge, like content information or citation distribution, to tackle these challenges.
+Comparatively, while SciBERT outperforms HINTS in fitting the increments of previous papers (SciBERT ¿ HINTS), it faces greater difficulty with fresh papers (SciBERT ¡ HINTS).
+This implies that graph structures exhibit better generality and transferability to unseen samples, while shared content information remains robust for immediate papers with limited graph structures, both alleviating issues related to data distribution shifts.
 (2) Both dynamic and heterogeneous graphs are important for our task.
@@
 For instance, DisenHAN, compared to its backbone DisenGCN, struggles to encode valuable information within heterogeneous graphs, notably within the complex dataset Psychology.
-[…]
-[…]
+(3) While Disentanglement methods offer performance improvements, their adaptation to dynamic heterogeneous graphs is essential to achieve optimal flexibility.
+Notably, DIDA exhibits superior outcomes than ROLAND in LogR2 for Chemistry, even when relying solely on homogeneous graphs.
+Moreover, within the densest dataset, Psychology, Disentangled GNN models showcase impressive performance.
+Among them, CAL outperforms both MUCas and ROLAND in predicting fresh and immediate papers.
+However, in other scenarios, apart from DIDA, these methods face challenges in competing with task-specific models and dynamic GNNs.
+(4) Our proposed model surpasses all baselines, by effectively modeling the dynamic heterogeneous graph and incorporating task-adjusted disentanglement.
+This leads to an average improvement of about 10% in MALE across all categories, along with a substantial advancement of over 30% in LogR2 for fresh and immediate papers.
+First, compared with traditional citation/cascade prediction models, it exhibits superior generalization and practicability through the fusion of content and context information within citation networks, as well as the disentanglement power.
+Content information aids in discerning the impact of previous papers and fortifies robustness towards immediate papers.
+Moreover, its adeptness in handling fresh and immediate papers stems from the rich information embedded within graph structures.
+Furthermore, we consider the unique properties of citation networks to encode complex high-order structural and temporal information, thereby outperforming common dynamic models.
+The proposed CGE facilitates the identification of influential papers related to the target paper based on comparative and multi-hop information.
+More importantly, our model introduces interpretability from disentanglement models without compromising performance and further adapts to the task for enhanced improvement.
+It excels in predicting precise citation increments and approaches the genuine contribution of papers by disentangling values derived from information diffusion and collective conformity, thus providing insights into the papers’ authentic impact.
 5.2. Ablation Test (RQ2)
 To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules.
-As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2021), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering.
+As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2022), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering.
 For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP for prediction.
@@
 (c) $\beta$
-[…]
+Figure 4. Results of hyper-parameters test in terms of LogR2.
 To address RQ3, we conduct experiments to assess the influence of three critical hyper-parameters in our model design: the stacked layer number $L$, the importance of co-cited/citing strengths $\lambda$, and the weight of the disentanglement tasks $\beta$.

### v4

@@
-Disentangling the Potential Impacts of Papers into Diffusion, Conformity, and Contribution Values
+Predicting Scientific Impact Through Diffusion, Conformity, and Contribution Disentanglement
 Abstract.
-The potential impact of an academic paper is determined by various factors, including its popularity and contribution.
-Existing models usually estimate original citation counts based on static graphs and fail to differentiate values from nuanced perspectives.
-In this study, we propose a novel graph neural network to Disentangle the Potential impacts of Papers into Diffusion, Conformity, and Contribution values (called DPPDCC).
-Given a target paper, DPPDCC encodes temporal and structural features within the constructed dynamic heterogeneous graph.
-Particularly, to capture the knowledge flow, we emphasize the importance of comparative and co-cited/citing information between papers and aggregate snapshots evolutionarily.
-To unravel popularity, we contrast augmented graphs to extract the essence of diffusion and predict the accumulated citation binning to model conformity.
-We further apply orthogonal constraints to encourage distinct modeling of each perspective and preserve the inherent value of contribution.
-To evaluate models’ generalization for papers published at various times, we reformulate the problem by partitioning data based on specific time points to mirror real-world conditions.
-Extensive experimental results on three datasets demonstrate that DPPDCC significantly outperforms baselines for previously, freshly, and immediately published papers.
-Further analyses confirm its robust capabilities.
-We will make our datasets and codes publicly available.
-Impact Prediction, Hierarchical and Heterogeneous Graph, Graph Neural Network, Disentangled Representation Learning, Contrastive Learning
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Theory of computation Dynamic graph algorithmsccs: Mathematics of computing Graph algorithms
+The scientific impact of academic papers is influenced by intricate factors such as dynamic popularity and inherent contribution.
+Existing models typically rely on static graphs for citation count estimation, failing to differentiate among its sources.
+In contrast, we propose distinguishing effects derived from various factors and predicting citation increments as estimated potential impacts within the dynamic context.
+In this research, we introduce a novel model, DPPDCC, which Disentangles the Potential impacts of Papers into Diffusion, Conformity, and Contribution values.
+It encodes temporal and structural features within dynamic heterogeneous graphs derived from the citation networks and applies various auxiliary tasks for disentanglement.
+By emphasizing comparative and co-cited/citing information and aggregating snapshots evolutionarily, DPPDCC captures knowledge flow within the citation network.
+Afterwards, popularity is outlined by contrasting augmented graphs to extract the essence of citation diffusion and predicting citation accumulation bins for quantitative conformity modeling.
+Orthogonal constraints ensure distinct modeling of each perspective, preserving the contribution value.
+To gauge generalization across publication times and replicate the realistic dynamic context, we partition data based on specific time points and retain all samples without strict filtering.
+Extensive experiments on three datasets validate DPPDCC’s superiority over baselines for papers published previously, freshly, and immediately, with further analyses confirming its robustness.
+Our codes and supplementary materials can be found at Github.
+Impact Prediction, Citation Network, Dynamic Heterogeneous Graph, Graph Neural Network, Contrastive Learning
+copyright: acmlicensedjournalyear: 2024conference: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management; October 21–25, 2024; Boise, ID, USAbooktitle: Proceedings of the 33rd ACM International Conference on Information and Knowledge Management (CIKM ’24), October 21–25, 2024, Boise, ID, USAdoi: 10.1145/3627673.3679546isbn: 979-8-4007-0436-9/24/10ccs: Information systems Data mining
 1. Introduction
-In paper retrieval and recommendation, ranking papers based on their scientific impact can aid researchers to delve into intricate research efforts.
+In paper retrieval and recommendation, ranking papers based on their scientific impact can aid researchers in delving into intricate research efforts.
 This is particularly significant considering the exponential annual growth in the number of published papers (Lo et al., 2020; Chu and Evans, 2021; Xue et al., 2023).
-Given the inherent difficulties in quantifying the precise numerical value of scientific impact, citation count is regularly employed as a rough approximation (Evans and Reimer, 2009; Sinatra et al., 2016; Jiang et al., 2021).
+Given the inherent difficulties in quantifying the precise numerical value of scientific impact, citation count is regularly employed as a rough approximation (Evans and Reimer, 2009; Sinatra et al., 2016; Jiang et al., 2021; Geng et al., 2022; Yang and Han, 2023).
 Nevertheless, the current ranking depends on the cumulative number of citations received to date, which merely represents the impact within the prior research environment.
 Therefore, it becomes imperative to cultivate a sophisticated model that can assess the potential impact of a research paper, placing significant emphasis on the future (Wang et al., 2013).
-At a given point in time (e.g., at present), estimating the anticipated increment in citations can effectively facilitate a fair comparison of the potential impact between long-published papers and newly published ones.
-Actually, the citations received by papers are not solely derived from the influence of their contributions, but are also affected by a multitude of other factors (Bornmann and Daniel, 2008; Case and Higgins, 2000).
+At a given point in time (e.g., at present), estimating the anticipated increment in citations can effectively serve as a better approximate measure of the potential impact, allowing for a fair comparison between long-published papers and newly-published ones.
+Actually, the citations received by papers are not solely derived from the influence of their contributions but are also affected by a multitude of other factors (Bornmann and Daniel, 2008; Case and Higgins, 2000).
 Hence, rather than directly forecasting original citation counts, it is more effective to disentangle citation increments received from extrinsic factors and the actual contributing factor, making impact prediction more practical and interpretable.
@@
 More precisely, when referring to the popularity factor of a paper, we are considering its citation diffusion patterns and the collective conformity degree within the citation network.
-Regarding citation diffusion, we posit that influential papers assume an amplifying effect in shaping the process of information dissemination within the citation network (Page et al., 1998; Wu et al., 2019).
+Regarding citation diffusion, we posit that influential papers assume an amplifying effect in shaping the information dissemination process within the citation network (Page et al., 1998; Wu et al., 2019).
 This leads to increased visibility of papers connected to them.
 In other words, papers cited by highly-cited papers will receive significantly more citations than others.
-The statistics of the dataset provides empirical support.
+The statistics of the datasets provide empirical support.
 For instance, in the field of Computer Science, the former papers will receive an average of 1.5475 citations (after applying a logarithmic transformation), while the latter ones will only receive an average of 0.8580 citations.
@@
 This phenomenon is substantiated by the future disparity (2.1917 vs.
-0.8009) in papers with different accumulated citations, which aligns with the concept of Matthew effect in the academic domain (Allison and Stewart, 1974).
-Therefore, by separating out the influence of the popularity, i.e., citation diffusion and collective conformity, we can gain a closer approximation to the genuine contributing impacts of papers.
-In essence, we aim to disentangle the citation increment into diffusion, conformity, and contribution values (as shown in Figure 1 (a)).
+0.8009) in papers with different accumulated citations, which aligns with the concept of the Matthew effect in the academic domain (Allison and Stewart, 1974).
+Therefore, by separating out the influence of the popularity, i.e., citation diffusion, and collective conformity, we can better approximate the genuine contributing impacts of papers.
+In essence, we aim to disentangle the citation increment into diffusion, conformity, and contribution values (as shown in Figure 1a).
 Figure 1.
 (a) We disentangle the citation increment into diffusion, conformity, and contribution values for better interpretability.
-(b) To imitate the evaluation of future dynamics of the citation context, we split the dataset by the observation time point and categorize the samples into previous, fresh, and immediate papers in terms of their publication time.
-(c) We represent the heterogeneous subgraph schema with all metadata, where each type of relation is bidirectional.
-Typically, conventional research (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022; Xu et al., 2022) on citation prediction imposes stringent constraints to select favorable samples by removing lowly-cited papers.
-Moreover, the division of these data into training, validation, and testing is carried out randomly, disregarding the chronological order of paper publications (Ruan et al., 2020; Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023).
-However, this problem setting deviates from real-world scenarios, possibly causing a substantial distribution shift.
-Hence, the model may face challenges in handling recently published papers.
-Random data partitioning disregards new citation network updates, hampering the model’s practicality and robustness in upcoming real-world test.
-This limitation impedes the exploration of the potential value of papers reactivated in the evolved context (Van Raan, 2004).
-In contrast, our work emulates reality by performing a novel problem setting.
-We include all papers without strict filtering to maintain consistency with the data distribution of real-world citation network.
-As illustrated in Figure 1 (b), we partition the dataset based on particular observation time points.
-This can introduce fresh papers that are not part of the training set (previous papers) into the test set.
-We further consider immediate papers (published at the test time point) that closely align with cold-start scenarios to validate the robustness of our model.
-Thus, papers in the test set can be divided into previous, fresh, and immediate papers, facilitating a more practical evaluation.
+(b) To imitate the evaluation of future dynamics within the citation context, we split the dataset by the observation time point and categorize samples into previous, fresh, and immediate papers in terms of publication time.
+(c) We represent the heterogeneous subgraph schema with all metadata, where each relation type is bidirectional.
+Typically, conventional research (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022b; Xu et al., 2022; He et al., 2023) on citation prediction imposes strict constraints to select favorable samples by removing lowly-cited papers.
+Additionally, their division of data into training, validation, and testing sets is typically performed randomly, ignoring the chronological order of paper publications (Ruan et al., 2020; Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023).
+Moreover, some studies focus solely on the most recently published papers, neglecting the value of those published previously (Jiang et al., 2021; Geng et al., 2022; Yang and Han, 2023).
+These problem settings deviate from real-world scenarios, potentially causing a substantial distribution shift.
+Consequently, models may struggle to handle papers published at different times.
+Overall, the random data partitioning setting disregards new updates to the citation network, hampering the model’s practicality and robustness in real-world applications, while the static problem setting impedes the exploration of the potential value of papers reactivated in the evolving context (Van Raan, 2004).
+In contrast, our work emulates reality by revising the dynamic problem setting of previous studies (He et al., 2023).
+We include all papers without strict filtering to maintain consistency with the real-world distribution of citation network data.
+As illustrated in Figure 1b, we partition the dataset based on specific observation time points, ensuring distinct training and testing windows.
+This methodology introduces fresh papers—excluded from the training set (previous papers) —as part of the test set, serving both as samples and context.
+We further consider immediate papers (published at the test time point) that closely align with cold-start scenarios, facilitating a more practical evaluation.
 Notably, previous papers have different contexts and future increments at the train and test time points.
-By employing this approach, our model is incentivized to fit the dynamic citation context, thereby enhancing its generalization.
-Existing graph disentanglement studies usually aim at decomposing abstractive hidden factors from either static or homogeneous graphs (see Section 2) (Ma et al., 2019a; Yang et al., 2020; Wang et al., 2020).
-Instead, our focus lies in disentangling the categorical characteristics of citation behaviors within dynamic and heterogeneous citation graphs.
-In this study, tailored to each target paper, we construct a Dynamic Heterogeneous Graph from the citation network.
-The graph consists of a sequence of heterogeneous snapshots (see Figure 1 (c)).
-This graph retains both the content of the paper and its dynamic citation context.
-Based on the proposed graph, we devise a model that could Disentangle the Potential impacts of Papers into Diffusion, Conformity, and Contribution Values (called DPPDCC).
-This model utilizes Citation-aware Relational Graph Convolutional Networks (R-GCNs) to capture the structural features of the input sequence of snapshots.
-The outputs are then passed through a Transformer Encoder, which allows for the incorporation of temporal information.
-Specially, we design a novel module in modified R-GCN to integrate the comparative and co-cited/citing information between paper nodes.
-Furthermore, we propose a Type-specific Attention Readout mechanism to aggregate information from each snapshot effectively.
-Upon acquiring the final representation of the target paper, we proceed to disentangle the predicted citation increment into diffusion, conformity, and contribution values individually.
-For the diffusion, we employ Triplet Citation-aware Graph Contrastive Learning to capture the popularity influence inherent in the information diffusion process within the citation network.
-For the conformity, we introduce an auxiliary task to classify bins of accumulated citations.
-Additionally, we impose an orthogonal constraint on the representations of disentangled perspectives in pair, guaranteeing that they encode distinct features without overlap.
-Extensive experiments are conducted on three fields of the real-world scientific dataset S2AG (Kinney et al., 2023).
-The results illustrate the superior performance of DPPDCC compared to baseline models across previous, fresh, and immediate papers.
-Ablation tests further highlight the significance of the proposed components.
-Additionally, visualizations offer insights into the model’s interpretability.
+Adopting this strategy incentivizes our model to fit the dynamic citation context, thereby enhancing its generalization.
+Existing graph disentanglement studies typically aim at implicitly decomposing abstractive hidden factors from either static or homogeneous graphs (see Section 2) (Ma et al., 2019a; Yang et al., 2020; Wang et al., 2020).
+In contrast, our focus is explicitly disentangling citation behaviors’ categorical characteristics within dynamic and heterogeneous citation graphs.
+For each target paper, we construct a Dynamic Heterogeneous Graph from the citation network, consisting of a sequence of heterogeneous snapshots (see Figure 1c), which retains both the content of the paper and its dynamic citation context.
+Based on this graph, we devise a model named DPPDCC that Disentangles the Potential impacts of Papers into Diffusion, Conformity, and Contribution values.
+The model first employs a Citation-aware Graph Neural Network Encoder (CGE) to extract distinctive structural and temporal information alternately.
+The Popularity-aware Disentanglement Module (PDM) then separates the values derived from different factors through three simultaneous auxiliary tasks.
+Specifically, the model utilizes Citation-aware Relational Graph Convolutional Networks (R-GCNs) to capture the structural features of the input snapshot sequence and a Transformer Encoder to incorporate temporal information.
+A novel message-passing module within the R-GCN, CompGAT, is designed to integrate comparative and co-cited/citing information between papers.
+Additionally, the Type-specific Attention Snapshot Readout is proposed to aggregate information from each snapshot evolutionarily.
+Upon acquiring the final representation of the target paper, we disentangle the predicted increment into diffusion, conformity, and contribution values.
+For diffusion, we employ Triplet Citation-aware Graph Contrastive Learning to capture the popularity influence inherent in the information diffusion process within the citation network.
+For conformity, we introduce an auxiliary task to classify bins of accumulated citations.
+Furthermore, orthogonal constraints are imposed on the representations of disentangled perspectives in pairs, ensuring they encode distinct features without overlap.
+Extensive experiments conducted on three fields of the real-world scientific dataset S2AG (Kinney et al., 2023) demonstrate DPPDCC’s superior performance compared to baseline models across previous, fresh, and immediate papers.
+Ablation tests further highlight the significance of the proposed components, and visualizations provide insights into the model’s interpretability.
 To sum up, the contributions of this research are as follows:
 $\bullet$ We revise the problem formulation of impact prediction to better align with real-world scenarios.
-Without filtering the predicted samples by their citation counts, we partition the datasets based on specific time points to maintain the dynamics of the citation context.
+Without filtering predicted samples by their citation counts, we partition the datasets based on specific time points to maintain the dynamics of the citation context.
 Moreover, we employ disentangled representation learning to extract the contributing impacts of papers.
-$\bullet$ We propose a novel model called DPPDCC that first encodes comparative and co-cited/citing information evolutionarily within the Dynamic Heterogeneous Graph through Citation-aware Graph Neural Network Encoder for potential impact prediction.
+$\bullet$ We propose a novel model, DPPDCC, which first encodes comparative and co-cited/citing information evolutionarily within the Dynamic Heterogeneous Graph through a Citation-aware Graph Neural Network Encoder.
 It further disentangles the citation increment into diffusion, conformity, and contribution values.
-$\bullet$ Experimental results show significantly superior to existing baselines for previous, fresh, and immediate papers. Further analyses illustrate that our model can reasonably make predictions.
+$\bullet$ Experimental results show that DPPDCC notably outperforms existing baselines for previous, fresh, and immediate papers, and further analyses confirm our model’s reasonable prediction ability.
 2. Related Work
-We review three lines of related work: Citation/Cascade Prediction, Dynamic and Heterogeneous Graph Neural Network (GNN), and Disentangled Representation Learning.
-2.1. Citation/Cascade prediction
-Citation prediction is a vital sub-task of automated academic evaluation as it enables the estimation of potential impact.
-It is closely examined within cascade prediction tasks, which share analogous graph structures and objectives (Zhou et al., 2021).
-The primary goal of cascade prediction is to anticipate the popularity of a post based on the interactions between users.
-Similarly, in the context of citation networks, the paper can serve as the post, while the authors can serve as the users.
-Citation count prediction places a focus on the content and distinctive features of individual papers, heavily relying on comprehensive feature extraction techniques.
-In contrast, cascade prediction delves into the intricate relationships between various entities, including papers and authors, directly exploiting the information of graph structure.
-Their approaches can be broadly classified into three categories: stochastic models, feature-based models, and deep learning models.
-Stochastic models can predict future citation counts by fitting the curve of citation counts (Glänzel and Schubert, 1995), following Zipf-Mandelbrot’s law (Silagadze, 1997).
-Recently, machine learning models have shown promising results by utilizing manually extracted features from various meta-data (Yan et al., 2011; Yu et al., 2014; Ruan et al., 2020; Guille et al., 2013; Tatar et al., 2014).
-More presently, deep neural networks have dominated this task by applying advanced Natural Language Processing (NLP) and Computer Vision (CV) methods to extract abstractive representations from paper content (Abrishami and Aliakbary, 2019; Huang et al., 2022; Xue et al., 2023).
-For cascade prediction, graph embeddings, sequence models, and GNN models are employed to extract structural information from the underlying graph, and then temporal information is encoded with sequence models (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022; Xu et al., 2022).
-However, existing models for citation and cascade prediction exhibit sub-optimal performance.
-They fail to fully exploit the valuable information presented in paper content and scientific context within the citation network simultaneously.
-Furthermore, their strict data selection and random data splitting strategies may hinder their practicality.
-Many valuable samples will be filtered and the temporal information within dynamic context will be neglected.
-Thus, models in this setting are more prone to face distribution shift issues.
-In contrast, our study proposes to specifically revise these settings and utilize both content and dynamic context for impact estimation.
-We also emphasize retaining genuine contributions to bolster the robustness of our models.
-2.2. Dynamic/Heterogeneous Graph Neural Networks
+We review three lines of related work: Citation/Cascade Prediction, Dynamic and Heterogeneous Graph Neural Networks (GNNs), and Disentangled Representation Learning.
+Citation prediction is a vital sub-task of automatic academic evaluation as it enables the estimation of scientific impact.
+It is closely examined within cascade prediction, which shares analogous graph structures and objectives (Zhou et al., 2021).
+Their approaches can be classified into three categories: stochastic, feature-based, and deep learning models.
+Stochastic models predict future citation counts by fitting citation curves (Glänzel and Schubert, 1995), following Zipf-Mandelbrot’s law (Silagadze, 1997).
+Recently, machine learning models have demonstrated promising results using manually extracted features from various metadata (e.g.
+content, author, venue) (Yan et al., 2011; Yu et al., 2014; Ruan et al., 2020; Dong et al., 2015).
+More presently, deep neural networks have dominated by applying advanced Natural Language Processing (NLP) and Computer Vision (CV) methods to extract abstractive representations from paper content (Abrishami and Aliakbary, 2019; Huang et al., 2022; Cohan et al., 2020; Xue et al., 2023).
+For cascade prediction, graph embeddings, sequence models, and GNN models extract structural information from the graph, with temporal information encoded by sequence models (Li et al., 2017; Cao et al., 2017; Xu et al., 2022; Cheng et al., 2023).
+However, most existing models exhibit suboptimal performance, failing to simultaneously exploit the valuable information presented in paper content and the scientific context within the citation network.
+Furthermore, their strict data selection and static random data splitting strategies may hinder their practicality.
+In contrast, our study revises the problem setting of impact prediction in the dynamic context (He et al., 2023; Yan et al., 2024).
+We ensure non-overlapping training and testing observation windows, retaining the real data distribution without sample filtering.
+We also utilize both paper content and dynamic context, and further introduce disentangled representation learning to provide a more practical impact estimation.
 GNNs (Sperduti and Starita, 1997; Gori et al., 2005; Scarselli et al., 2008; Wu et al., 2020) are widely applied to handle non-Euclidean data like graphs.
-Recently, research interest has surged in dynamic graphs, which incorporate temporal information, as well as in heterogeneous graphs, which involve multiple node types or edge types.
-These types of graphs are more prevalent in practical scenarios and encapsulate more intricate information, compared to static homogeneous graphs, driving the formulation and refinement of diverse methodologies within the academic realm.
-The dynamic graph can be divided into multiple snapshots at different time points.
-Previous models like DGCN (Manessi et al., 2020), Dysat (Sankar et al., 2020), and ROLAND (You et al., 2022) typically encode the snapshots with static GCN as the structural encoder, followed by sequential models like Recurrent Neural Networks (RNN) or Transformer Encoder.
-Moreover, certain models incorporate GCN into RNN models by configuring the parameters within RNN cells as GCN parameters (Pareja et al., 2020).
-Within heterogeneous graphs, on one side, R-GCN serves as a prominent model with its separate message-passing architecture (Schlichtkrull et al., 2018).
+Dynamic graphs incorporate temporal information, while heterogeneous graphs involve multiple types of nodes or edges.
+Dynamic graphs can be divided into snapshots at different time points.
+Previous models like DGCN (Manessi et al., 2020), Dysat (Sankar et al., 2020), and ROLAND (You et al., 2022) typically encode the snapshots with static GCN(s) as the structural encoder, followed by sequential models like RNN or Transformer Encoder.
+In heterogeneous graphs, R-GCN serves as a prominent model (Schlichtkrull et al., 2018).
 It conducts independent message passing within different relations and aggregates them to update node representations.
-On the other side, the relation-learning method transforms heterogeneous graphs into homogeneous ones, while retaining node and edge type indicators to preserve heterogeneous information (Hu et al., 2020; You et al., 2022).
-An example of this technique is HGT (Hu et al., 2020), which utilizes multi-head attention modules with specific parameters tailored to different node and edge types, effectively modeling the complexity of heterogeneous graphs.
-Citation network can be regarded as dynamic and heterogeneous graphs, as it evolves over time and encompasses diverse entities such as papers, authors, and venues.
-Addressing the complexities within this network involves two primary challenges: (1) How to encode the temporal target-centric information within a single snapshot?
-(2) How to model distinct characteristics of information diffusion in citation networks?
-To overcome these challenges, our study proposes a Citation-aware GNN Encoder that effectively models both dynamic and heterogeneous graphs of target papers.
-These graphs are constructed from the citation network on an annual basis and incorporate multiple metadata nodes beyond papers themselves.
-(1) To encode the temporal target-centric information within a single snapshot, we employ a type-aware attention readout.
-It aggregates snapshot information considering paper type and publication time.
-(2) To model distinct characteristics of information diffusion in citation networks, we propose a novel GCN module CompGAT.
-This module focuses on capturing comparisons and co-cited/citing information between papers, drawing inspiration from bibliometrics theories (Wu et al., 2019).
-2.3. Disentangled Representation Learning
-Disentangled representation learning, aiming to discern and separate fundamental explanatory factors (Bengio et al., 2013), stands as a pivotal method within deep learning.
-It strives to produce resilient, manageable, and interpretable representations.
-Existing research predominantly focuses on the field of CV, often employing methods like Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) to disentangle relevant latent factors (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021).
-Researchers explore techniques to align distinct representations with specific factors through explicit and implicit guidance, employing approaches like direct labeling, loss penalties, and regularization.
-For instance, DR-GAN (Tran et al., 2017) utilizes an adversarial network to generate diverse facial poses, aiming to disentangle pose information from facial attributes.
-In NLP, disentangled representation learning finds applications in various tasks like generation.
-For example, John et al.
-(John et al., 2019) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style.
-This separation allows precise control over content and style in text generation.
-Recently, there has been a growing interest in applying disentangled representation learning to graph learning tasks.
-Given that graph data encompasses both structural and attributed information, numerous new methodologies have emerged to address this complexity.
-In the feature space, inspired by capsule networks, approaches have emerged to partition node features into multiple hidden channels (Ma et al., 2019a).
-Some methods introduce additional constraints to enable effective separation (Liu et al., 2020).
-Moving into the structural space, subsequent models strive to factorize the input graphs into multiple subgraphs, to facilitate distinct message-passing strategies (Yang et al., 2020; Li et al., 2021).
-Recent models have expanded their scope to isolate causal and biased information presented in both the graph structure and node features, with a particular emphasis on causal effects (Sui et al., 2022; Fan et al., 2022).
-These models commonly employ techniques to generate masks within the adjacency matrix and node features.
-This process aims to identify the crucial elements of the graph within the structural and attributed spaces.
-In addition, certain studies extend disentangled GNNs to specialized scenarios such as recommendation systems, handling intricate graph structures in heterogeneous or dynamic graphs (Wang et al., 2020; Zhang et al., 2022; Wen et al., 2022).
-Besides the masking strategy considering solely the graph structures and node features, they integrate classical methods like VAE and adversarial attack learning from CV and NLP to handle task-specific information from a broader perspective.
-In contrast to previous studies, our approach acknowledges that the potential impact within the citation network is influenced not only by the paper’s contribution but also by factors related to its popularity.
-These popularity factors encompass the amplifying effect owing to the highly-cited paper nodes in information diffusion and the manifestation of the Matthew effect through collective conformity.
-Based on the dynamic heterogeneous graph, we disentangle the potential impact into diffusion, conformity, and contribution values, seeking a more comprehensive estimation.
-Specifically, we devise dedicated auxiliary tasks aimed at extracting values associated with these respective factors.
+Our study proposes a Citation-aware GNN Encoder that effectively models both dynamic and heterogeneous graphs of target papers constructed from the citation network.
+It utilizes modified R-GCNs as structural encoders and a multi-layer transformer as the temporal encoder.
+Additionally, we incorporate bibliometric theory to model distinct features within the citation network.
+Disentangled representation learning aims to identify and disentangle the underlying explanatory factors (Bengio et al., 2013).
+The existing efforts primarily focus on CV and NLP, utilizing Variational Auto-encoders (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021).
+In graph learning tasks inspired by capsule networks, they partition node features into multiple hidden channels for implicit disentangled modeling (Ma et al., 2019a).
+Subsequent models strive to factorize the input graphs to facilitate distinct message passing (Yang et al., 2020).
+Recent models expand their scope to isolate causal and biased information presented in the graph structure and node features, with a particular emphasis on causal effects (Sui et al., 2022; Fan et al., 2022).
+In addition, certain studies extend the disentangled GNNs to specialized scenarios like recommendations and healthcare, addressing more complex graph structures involving either heterogeneous or dynamic graphs (Wang et al., 2020; Zhang et al., 2022; Wen et al., 2022; Zhang et al., 2023b, a).
+In contrast to previous studies, our approach acknowledges that the potential impact within the citation network is influenced not only by the paper’s contribution but also by popularity factors.
+These popularity factors encompass the amplifying effect owing to the highly-cited paper nodes in information diffusion and the manifestation of the Matthew effect as collective conformity.
+Building upon previous disentangled studies in recommendations (Zheng et al., 2021; Chen et al., 2022a), we regard the final predicted values as mixed results of multiple factors and introduce auxiliary tasks for separate encoding of different perspectives.
+Leveraging the dynamic heterogeneous graph, we disentangle the potential impact into diffusion, conformity, and contribution values, aiming for a more rational estimation through explicit modeling.
 3. Methodology
-In this section, we present a detailed exposition of our proposed model, DPPDCC.
-Initially, we formulate the constructed dynamic heterogeneous and the potential impact prediction task.
-Following this, we provide an in-depth elucidation of DPPDCC’s core components: the Citation-aware GNN Encoder (CGE) and the Popularity-aware Disentanglement Module (PDM), as depicted in Figure 2.
-The CGE integrates multiple layers of structural and temporal encoders, aimed at capturing intricate high-order temporal and structural information within the network.
-Concurrently, the PDM comprises three distinct modules, each targeting a distinct perspective of citation increment, all trained in parallel.
-Lastly, we delineate the training and optimization strategies employed for DPPDCC.
 Figure 2.
 The overall architecture of DPPDCC.
-Initially, DPPDCC processes the dynamic heterogeneous graph of the target paper using the Citation-aware GNN Encoder.
-Within this encoder, the structural component utilizes CompGAT to extract comparative and co-cited/citing information between papers at each time step.
-This information is then aggregated evolutionarily across snapshots via type-specific attention readout.
-Subsequently, aggregated snapshot information from different time steps is interacted in the temporal encoder.
-By stacking multiple layers of the structural encoder and temporal encoder, it enables CGE to encode complex high-order temporal and structural information effectively.
-Subsequent to obtaining the target paper’s representation, DPPDCC proceeds to disentangle its citation increment into diffusion, conformity, and contribution values through three corresponding auxiliary tasks.
-The diffusion module augments the graphs by employing weighted node dropping, facilitating the identification of the diffusion process by contrasting augmented views with the original graph.
-The conformity module quantifies conformity levels by binning accumulated citations.
-Finally, the contribution module introduces orthogonal constraints, prompting each module to learn distinct information from diverse perspectives.
+DPPDCC first encodes the Dynamic Heterogeneous Graph of the target paper with the Citation-aware GNN Encoder.
+Based on the encoded representation of the target paper, DPPDCC then disentangles its citation increment into diffusion, conformity, and contribution values through three corresponding auxiliary tasks.
+In this section, we introduce our proposed model DPPDCC in detail.
+The overall framework of DPPDCC is depicted in Figure 2, which can be divided into two parts: Citation-aware GNN Encoder (CGE) and Popularity-aware Disentanglement Module (PDM).
 3.1. Problem Formulation and Notations
-Before presenting our approach, we first introduce the problem formulation and notations.
 In practice, the citation network comprises diverse entities and relations that evolve dynamically each year.
-To capture the entity-aware and time-variant characteristics, of each target paper, we construct a dynamic heterogeneous graph from the citation network to predict its potential impact.
+To capture the entity-aware and time-variant characteristics, we construct a dynamic heterogeneous graph from the citation network for each target paper to predict its potential impact.
 3.1.1. Dynamic Heterogeneous Graph
@@
 And $\epsilon^{t}$ represents the set of edges among $\nu^{t}$.
-Each node $\nu$ and edge $\epsilon$ are associated with their type mapping functions $\phi:\nu\rightarrow U$ and $\varphi:\epsilon\rightarrow R$, where $U$ and $R$ denote the types of nodes and edges (see Figure 1 (c)), respectively.
+Each node $\nu$ and edge $\epsilon$ are associated with their type mapping functions $\phi:\nu\rightarrow U$ and $\varphi:\epsilon\rightarrow R$, where $U$ and $R$ denote the types of nodes and edges shared across all times (see Figure 1c), respectively.
 The backbone of the heterogeneous subgraph primarily consists of the cited and citing relations.
 We collect $k$-hop cited/citing neighbors of the target paper at the specific time $t$.
-For each $i$-th hop, to maintain the most recent and significant information, we select the $K_{i}$ newest published and then most highly cited papers, for references and citations respectively.
-To address the potential lack of global information, we incorporate metadata nodes associated with the papers (authors, venue, and publication time).
+For each node in $i$-th hop, to maintain the most recent and significant information, we select the $K_{i}$ newest published and then most highly cited papers as neighbors.
+To address the potential lack of global information, we incorporate metadata nodes associated with the papers (see Figure 1c).
 Prior to the observation time point, we consider a total of $T$ such subgraphs, each serving as a snapshot captured at distinct time steps.
@@
 3.1.2. Potential Impact Prediction
-We assess the potential impact of the given target paper by estimating its citation increment.
-It is quantified as the increased citation number $\Delta$ years after the observation time point.
-Given a paper, we try to learn a function $f(\cdot)$ to predict the number of citation increments.
+We estimate the potential impact of the target paper by predicting its citation increment, defined as the increase in citation count $\Delta$ years after the observation time point.
+By disentangling the sources of this increment, our model can more accurately assess the genuine potential impact, thereby providing a closer approximation to the real scientific significance.
 3.2. Initialization
@@
 Specifically, the extraction of text embeddings is accomplished using Sentence-BERT (Reimers and Gurevych, 2019).
-For $i$-th node at the $t$-th time step, we have:
-(1) $\begin{split}&\bm{h}_{ti,paper}^{(0)}=TextEncoder([w_{1},w_{2},...,w_{n^{w}}])\\ &\bm{h}_{ti,u}^{(0)}=avgpool([\bm{h}_{t1,paper}^{(0)},\bm{h}_{t2,paper}^{(0)},...,\bm{h}_{tm_{p},paper}^{(0)}]),\\ &u\in\{author,venue,time\}\end{split}$
-where $u$ denotes the node type, $TextEncoder$ denotes the text representation model, and $m_{p}$ is the count of metadata node’s connected papers.
-To initialize the 00-th layer representation of the snapshot of the $i$-th target paper at the $t$-th time step $\bm{h}_{ti,s}^{(0)}$, we mainly consider the contextual information from the corresponding time embedding $\bm{h}_{t,time}^{(0)}$ at time step $t$.
-Since the time embedding only preserves the global features, we utilize a gate module to approach specific information based on the $i$-th target paper embedding $\bm{h}_{i,p}^{(0)}$.
-The snapshot representation $\bm{h}_{ti,s}^{(l)}$ of $l$-th layer will then be the proxy to aggregate the particular heterogeneous subgraph:
-(2) $\begin{split}&\bm{h}_{ti,s}^{(0)}=\bm{a}_{ti}\times\bm{h}_{t,time}^{(0)}\\ &\bm{a}_{ti}=\mathrm{sigmoid}(\bm{W}^{ipt}\bm{h}_{t,time}^{(0)}+\bm{W}^{tgt}\bm{h}_{i,p}^{(0)})\end{split}$
+To initialize the 00-th layer representation of the snapshot of the $i$-th target paper at the $t$-th time step $\bm{h}_{ti,s}^{(0)}$, we mainly take into account the contextual information from the corresponding time embedding $\bm{h}_{t,time}^{(0)}$ of time step $t$.
+Since the time embedding only preserves the global features, we utilize a gate module to approach specific information from the $i$-th target paper embedding $\bm{h}_{i,p}^{(0)}$.
+The snapshot representation $\bm{h}_{ti,s}^{(l)}$ of $l$-th layer will then be the proxy to aggregate the corresponding heterogeneous subgraph:
+(1) $\begin{split}&\bm{h}_{ti,s}^{(0)}=\bm{a}_{ti}\times\bm{h}_{t,time}^{(0)}\\ &\bm{a}_{ti}=\mathrm{sigmoid}(\bm{W}^{input}\bm{h}_{t,time}^{(0)}+\bm{W}^{target}\bm{h}_{i,p}^{(0)})\end{split}$
 3.3. Citation-aware GNN Encoder
-Citation-aware GNN Encoder (CGE) aims to capture distinct structural and temporal features from the dynamic heterogeneous graph, which is facilitated by leveraging knowledge diffusion theories derived from bibliometrics.
-To be specific, the backbone of CGE consists of modified R-GCNs as the structural encoder and a Transformer Encoder as the temporal encoder.
+Citation-aware GNN Encoder (CGE) aims to capture distinct structural and temporal features from the dynamic heterogeneous graph.
+To be specific, the backbone of CGE consists of modified R-GCNs (Schlichtkrull et al., 2018) as the structural encoder and a Transformer Encoder (Vaswani et al., 2017) as the temporal encoder.
 We stack $L$ integrated layers to encode them alternately instead of separate modeling.
 This allows us to exploit higher-order interactions between space and time.
-To elaborate, we first employ distinct Citation-aware R-GCNs to model structural information for subgraphs at different time points individually.
-Afterwards, a Type-specific Attention Snapshot Readout is employed to summarize the information from these subgraphs established on the snapshot representations.
-Simultaneously, the snapshot information mixer integrates global and temporal information from snapshot representations into the local and structural representations of paper nodes, fostering interactions among various feature spaces.
-The aggregated snapshot representations are then input into the temporal encoder (Encoder of the Transformer) to capture the intricate temporal relationships between different time points.
+To elaborate, for subgraphs at different time points, we first utilize distinct Citation-aware R-GCNs to model their structural information respectively.
+Afterwards, the Type-specific Attention Snapshot Readout is employed to summarize the information from these subgraphs based on the snapshot representations.
+These aggregated representations are then fed into the temporal encoder to capture the intricate temporal relationships between different time points.
 By integrating structural and temporal encoding, our approach could extract valuable insights from the dynamic heterogeneous graph.
 3.3.1. Citation-aware R-GCNs
-Citation-aware R-GCNs, as the core component of the structural encoder, are utilized to encode node hidden states $\bm{h}$ within the dynamic heterogeneous graph.
-Following the manner of R-GCN (Schlichtkrull et al., 2018), we deliver message passing in each relation separately for modeling attentively.
-Moreover, we replace the original GCNs in R-GCN with more powerful GNN modules to sensitively encode intricate information.
+Citation-aware R-GCNs are utilized to encode node hidden states $\bm{h}$ within the dynamic heterogeneous graph for structural modeling:
 (2) $\begin{split}&\bm{h}_{dst}^{(l+1)}=\underset{r\in\mathcal{R},r_{dst}=dst}{AGG}(f_{r}(g_{r},\bm{h}_{r_{src}}^{(l)},\bm{h}_{r_{dst}}^{(l)}))\\ \end{split}$
-where $\mathcal{R}$ is the set of all edges in the graph, $g_{r}$ is the subgraph with relation $r$, and $AGG$ is the relation aggregation function.
-Here, we adopt sum-pooling as the relation aggregation function for efficiency and performance.
-For ”cites” and ”is cited by” edge subgraphs of paper nodes, inspired by Disruption-index (Wu et al., 2019), we introduce co-cited/citing strengths to integrate multi-hop citation information.
-The Disruption-index categorizes papers as disruptive or developmental based on the metric related to their betweenness centrality.
-Disruptive papers play a critical role in altering typical citation diffusion.
-We adapt this subgraph-focused metric to estimate node-pair closeness by considering co-cited/citing strengths.
-We aim to discern the varying significance of citation behaviors during message passing and distinguish crucial references and citations within ”is cited by” and ”cites” edges for the target paper.
-On one side, within ”is cited by” edges, when the source (reference) and target (citation) papers share more similar references, the influence of the source paper on the target paper might diminish.
-From the viewpoint of the target paper, the source paper could be considered as merely developmental, exerting a potentially negative impact on message passing.
-Moreover, since the information from this source paper is already conveyed through other original disruptive references, the negative co-citing effect might amplify attention toward these disruptive source papers.
-On the other side, within “cites” edges, if the source (citation) and target (reference) paper have more identical citations, they may share more commonness and are more closely connected, leading to a positive effect.
-Unlike co-citing relationships, citations are expansive and continuously growing, making it challenging to differentiate disruptive papers based on co-cited relationships alone.
-Therefore, we employ a normalized score of co-cited strength that considers both similarity and accumulated count to identify highly-related developmental citations.
-It’s crucial to note that the co-cited/citing strengths are dynamic scores that evolve over time.
-As new papers emerge, the values of normalized co-cited/citing strengths undergo changes not only due to the direct influence of new citations but also because of alterations in numerical values and maximum values.
-Leveraging these co-cited/citing strengths can help uncover the dynamics inherent in multi-hop information diffusion.
-Overall, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2022), to encode comparative and co-cited/citing features between paper nodes.
+where $\mathcal{R}$ is the set of all edges in the graph, $f_{r}$ is the message passing method applied to relation $r$, $g_{r}$ is the subgraph with relation $r$, $\bm{h}_{r_{src/dst}}$ is the hidden representation of the source/target node of relation $r$, and $AGG$ is the relation aggregation function.
+Motivated by the Disruption-index (Wu et al., 2019), in ”is cited by” edges, if the source and target paper have more identical references, the source paper may be less influenced by the target paper, resulting in a negative effect.
+Conversely, in ”cites” edges, if the source and target paper have more identical citations, they may share more commonality and are more closely connected, leading to a positive effect.
+Thus, we introduce CompGAT, a novel module based on GATv2 (Brody et al., 2022), to encode comparative and co-cited/citing features between paper nodes.
 To capture the discrepancy between source and target nodes, we encode the concatenation of their representations.
 Besides the basic attention scores of GATv2, we incorporate co-cited or co-citing strengths by assessing the normalized similarity of their references or citations in the global citation network.
-Based on the adjacent matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes.
+Based on the adjacency matrix $\bm{A}$, it is calculated as $\bm{C}=\mathrm{norm}(\bm{A}\bm{A}^{T})$, which represents 2-hop relationships connected with intermediate paper nodes.
 Finally, the attention is a mixed distribution resembling CopyNet (Gu et al., 2016).
-For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$, we have:
+For $i$-th paper node in $(l+1)$-th layer $\bm{h}_{i}^{(l+1)}$:
 (3) $\begin{split}&\bm{h}_{i}^{(l+1)}=\mathrm{LeakyReLU}\left(\mathrm{layernorm}(\bm{h}_{i}^{(l)}+\bm{h}_{in}^{(l+1)})\right)\\ &\bm{h}_{in}^{(l+1)}=\sum_{j\in\mathcal{N}(i)}\alpha_{ij}^{(l)}\bm{W}_{c}^{(l)}[\bm{W}^{(l)}_{left}\bm{h}_{i}^{(l)}\|\bm{W}^{(l)}_{right}\bm{h}_{j}^{(l)}]\\ &\alpha_{ij}^{(l)}=\lambda\mathrm{softmax_{i}^{c}}(c_{ij})+(1-\lambda)\mathrm{softmax_{i}^{e}}(e_{ij}^{(l)})\\ &e_{ij}^{(l)}=\bm{w}_{a}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{left}\bm{h}_{i}^{(l)}+\bm{W}^{(l)}_{right}\bm{h}_{j}^{(l)}\right)\end{split}$
 where $c_{ij}\in\bm{C}$ is the normalized co-cited/citing strength, $\lambda$ is to balance the sum distribution of co-cited/citing strengths and basic attention weights, and $\bm{w}_{a}^{(l)}$is the attention vector.
-For other edges in the R-GCN, we apply GIN (Xu et al., 2019) with sum-pooling as the GNN encoder:
-(5) $\begin{split}\bm{h}_{i,r_{dst}}^{(l+1)}=f_{\Theta,r}((1+\xi)\bm{h}_{i,r_{dst}}^{(l)}+AGG(\{\bm{h}_{j,r_{src}}^{(l)},&j\in\mathcal{N}(i,r_{src})\})),r\in\mathcal{R}^{normal}\end{split}$
-where $\xi$ is the hyper-parameter to decide the retained original information, $\mathcal{R}^{normal}$ is the set including all edges in classic GIN, and $AGG$ is the neighborhood aggregation function of the GIN.
-In our model, we opt for sum-pooling as the aggregation method, as it can preserve crucial structural information including node degrees.
+Additionally, we apply GIN (Xu et al., 2019) with sum-pooling for other edges in the R-GCN.
 3.3.2. Type-specific Attention Snapshot Readout
-To represent the time-aware information and filter noisy information of a single snapshot $\bm{h}_{t,s}$, we design an attention-based readout mechanism.
+To represent the time-aware information of a single snapshot $\bm{h}_{t,s}$, we design an attention-based readout mechanism.
 It extracts the evolved temporal information of different types of paper nodes (reference, citation, and target paper) based on the publication age of the target paper.
-(6) $\begin{split}&\bm{h}_{t,s}^{(l+1)}=\bm{h}_{t,s}^{(l)}+\sum_{j\in\mathcal{N}(t)}\hat{\alpha}_{tj}^{(l)}\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\\ &\hat{\alpha}_{tj}^{(l)}=\mathrm{softmax_{t}}(\hat{e}_{tj}^{(l)})\\ &\alpha_{tj}^{(l)}={\bm{w}_{a}}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{s}\bm{h}_{t,s}^{(l)}+\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\right)+{\bm{w}_{t}}^{(l)}\mathrm{LeakyReLU}\left(f^{(l)}_{se}(\omega^{s}_{t})+f^{(l)}_{pe}(\omega^{p}_{j})\right)\end{split}$
+(4) $\begin{split}&\bm{h}_{t,s}^{(l+1)}=\bm{h}_{t,s}^{(l)}+\sum_{j\in\mathcal{N}(t)}\alpha_{tj}^{(l)}\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\\ &\alpha_{tj}^{(l)}=\mathrm{softmax_{t}}(\hat{e}_{tj}^{(l)})\\ &\hat{e}_{tj}^{(l)}={\bm{w}_{a}}^{(l)}\mathrm{LeakyReLU}\left(\bm{W}^{(l)}_{s}\bm{h}_{t,s}^{(l)}+\bm{W}^{(l)}_{p}\bm{h}_{j,p}^{(l)}\right)+e_{tj}^{(l)}\\ &e_{tj}^{(l)}={\bm{w}_{t}}^{(l)}\mathrm{LeakyReLU}\left(f^{(l)}_{se}(\omega^{s}_{t})+f^{(l)}_{pe}(\omega^{p}_{j})\right)\end{split}$
 where $\mathcal{N}(t)$ denotes the paper nodes within $t$-th 1-hop subgraph, $f^{(l)}_{se}(\omega^{s}_{t})$ and $f^{(l)}_{pe}(\omega^{p}_{j})$ are snapshot and paper type embeddings.
-The snapshot type is determined by binning the time interval between the time step $t$ and the publication time of the target paper.
-In this study, we categorize snapshot types based on temporal segments across all datasets: $\leq 5$ years, $5-10$ years, and $>10$ years.
-To enhance relevance across diverse fields of study, a potential refinement could involve tailoring these segments to align with the half-life of literature aging within specific domains.
-3.3.3. Snapshot Information Mixer
-To foster interactions between temporal and spatial dimensions, as well as between global and local information, we amalgamate the representations of paper nodes with the updated snapshot representation:
-(7) $\begin{split}&\hat{\bm{h}}_{ti,paper}^{(l+1)}=mlp_{m}^{(l)}([\bm{h}_{ti,paper}^{(l+1)}\|\bm{h}_{t,s}^{(l+1)}])\end{split}$
-where $mlp_{m}^{(l)}$ integrates and condenses the information from both the current paper node and the snapshot.
-leveraging the temporal information embedded in the snapshot representations through the mixer, we introduce extra temporal signals, enriching the structural encoder beyond sole reliance on graph structures.
+The snapshot type is determined by binning the time interval between time step $t$ and the publication time of the target paper.
 3.3.3. Snapshot Temporal Encoder
-To encode the temporal information across various snapshot times, we utilize a 4-layer Transformer Encoder.
-Compared with classic sequential models like RNN, the self-attention mechanism captures more intricate relationships that extend beyond strict ordinal dependencies:
+To encode the temporal information within snapshot representations across various times $\bm{H}_{s}^{(l+1)}$, we utilize a multi-layer Transformer Encoder (Vaswani et al., 2017).
+Compared with classic sequential models, the self-attention mechanism captures more intricate relationships beyond strict ordinal dependencies:
 (5) $\begin{split}&\hat{\bm{H}}_{s}^{(l+1)}=\mathrm{MultiHeadSelfAttn}(\bm{H}_{s}^{(l+1)})\end{split}$
 3.3.4. Final Output
-We obtain the target paper representation $\bm{o}$ by avg-pooling across the layers and then sum-pooling across the time steps, to maintain high-order temporal information:
+We obtain the target representation $\bm{o}$ by avg-pooling across the layers and then sum-pooling across the time steps to maintain high-order temporal information:
 (6) $\begin{split}&\tilde{\hat{\bm{H}}}_{s}=avgpool^{L}([\hat{\bm{H}}_{s}^{(1)},\hat{\bm{H}}_{s}^{(2)},\cdots,\hat{\bm{H}}_{s}^{(L)}])\\ &\bm{o}=sumpool^{T}(\tilde{\hat{\bm{H}}}_{s})\end{split}$
 3.4. Popularity-aware Disentanglement Module
-Based on the encoded representation of the target paper from CGE, we design the Popularity-aware Disentanglement Module (PDM) to disentangle the citation count potentially received from different perspectives, including citation diffusion, collective conformity, and actual contribution.
-Overall, we sum up the individual predicted values to obtain the total citation increment and apply different auxiliary tasks to help the disentanglement.
-This approach enables the model to learn distinctive compositions for diverse papers, thereby facilitating the identification and utilization of their authentic contributions.
-We employ distinct Multilayer Perceptrons (MLPs) as encoders to process information within the target perspectives, and then feed them into another MLP separately to predict the corresponding citation increment $\hat{y}_{i}^{v}$:
+We design the Popularity-aware Disentanglement Module (PDM) to disentangle the citation count potentially received from different perspectives, including citation diffusion, collective conformity, and actual contribution.
+Overall, we train the three components simultaneously and sum up their predicted values to obtain the original increment.
+We further apply different auxiliary tasks to help the disentanglement.
+This approach enables the model to learn distinctive compositions for diverse papers, thereby facilitating the identification of their authentic contributions.
+Specifically, we utilize separate Multilayer Perceptrons (MLPs) as encoders and predictors to handle information within the target perspectives.
+For the target perspective $v$, we begin by deriving its representation $\hat{\bm{o}}_{i}^{v}$ from the original target representation $\bm{o}_{i}$ using its encoder $mlp^{v}_{enc}$.
+Subsequently, we input this representation into its corresponding predictor $mlp^{v}_{pred}$ to anticipate its increment $\hat{y}_{i}^{v}$:
 (7) $\begin{split}&\hat{\bm{o}}_{i}^{v}=mlp^{v}_{enc}(\bm{o}_{i})\\ &\hat{y}_{i}^{v}=mlp^{v}_{pred}(\hat{\bm{o}}_{i}^{v})\\ \end{split}$
-where $mlp^{d}_{enc}$ encodes specific aspects of the information and $mlp^{v}_{pred}$ predicts the citation increment from the target perspective.
-It is noteworthy that the three dedicated components are separately modeled to disentangle distinct perspectives, and they are trained simultaneously to facilitate comprehensive and concurrent learning.
 3.4.1. Diffusion Encoder
-The diffusion encoder extracts the popularity influence according to information diffusion, which is represented as diversity affected by spreading nodes’ degrees.
-We propose a Citation-aware Triplet Graph Contrastive Learning method to extract the trunk of spreading in the dynamic graph.
-In various snapshots, except for the target paper node, all other paper nodes are selectively omitted with probability, which is weighted by their normalized global citations within the local subgraph.
-Lowly-cited paper nodes are more likely to be dropped for the positive view $z^{pos}$, as they contribute less to spreading.
-Conversely, for the negative view $z^{neg}$, highly-cited paper nodes are more likely to be dropped, as they significantly impact the graph semantics.
-For every sample, we contrast them solely with their augmented views to preserve their unique characteristics from the diffusion perspective.
-We augment the subgraphs in every training epoch to enhance the robustness of learning.
-Through this approach, we amplify diffusion influence in the graph structure.
+The diffusion encoder captures the influence of popularity based on information diffusion, which is characterized by diversity affected by the degrees of spreading nodes.
+To extract the trunk of spreading, we propose a Citation-aware Triplet Graph Contrastive Learning method.
+We employ citation/degree-aware node-dropping (Xu et al., 2022; He et al., 2023) as the graph augmentation technique.
+In different snapshots, apart from the target paper node, all other paper nodes are selectively omitted with probability, which is weighted by their normalized global citations within the local subgraph.
+Lowly-cited paper nodes are excluded for the positive view $z^{pos}$ since they contribute less to spreading.
+Conversely, for the negative view $z^{neg}$, highly-cited paper nodes are omitted, given their significant impact on the graph semantics.
+This approach amplifies the influence of diffusion in the graph structure.
 Contrasted with original sample $z^{ori}$, the diffusion loss $\mathcal{L}_{dif}$ is:
@@
 (9) $\begin{split}&\hat{y}_{im}^{con}=mlp(\hat{\bm{o}}_{i}^{con})\\ &\mathcal{L}_{con}=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{i}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{m=1}^{M}y_{im}^{con}\log(\hat{y}_{im}^{con})\end{split}$
-where $m$ is the number of bins for equal frequency binning.
 3.4.3. Contribution Encoder
 After disentangling the popularity influences, the remaining citation increment can serve as the genuine contribution.
-To further encourage distinct encoding of various aspects for better approximation of contribution, we apply an orthogonal regularization $\mathcal{L}_{ort}$ for each pair of perspectives:
+To further encourage distinct encoding of various aspects for better approximation of its contribution, we apply the orthogonal regularization $\mathcal{L}_{ort}$ for each pair of perspectives:
 (10) $\begin{split}&\mathcal{L}_{ort}=\frac{1}{N}\sum_{i=1}^{N}\frac{\|\hat{\bm{o}}_{i}^{v1}*\hat{\bm{o}}_{i}^{v2}\|^{2}}{\|\hat{\bm{o}}_{i}^{v1}\|*\|\hat{\bm{o}}_{i}^{v2}\|}\end{split}$
@@
 3.4.4. Disentanglement Loss
-We can obtain the final disentanglement loss $\mathcal{L}_{dis}$ by gathering the losses of all perspectives:
+We can obtain the final disentanglement loss $\mathcal{L}_{dis}$ by gathering the losses of different perspectives:
 (11) $\begin{split}&\mathcal{L}_{dis}=\mathcal{L}_{dif}+\mathcal{L}_{con}+\mathcal{L}_{ort}\end{split}$
@@
 Considering that the distribution of citation count is skew distribution, we apply logarithmization to normalize the distribution.
-Thus, the main loss function $\mathcal{L}_{reg}$ of the potential impact prediction task is the mean squared error (MSE) between the logarithmized target and predicted citation increments: $y_{i}$ and $\hat{y}_{i}^{all}$.
+Thus, the primary loss function $\mathcal{L}_{reg}$ of the potential impact prediction task is the mean squared error (MSE) between logarithmized target and predicted citation increments: $y_{i}$ and $\hat{y}_{i}^{all}$.
 (12) $\begin{split}&\hat{y}_{i}^{all}=\hat{y}_{i}^{dif}+\hat{y}_{i}^{con}+\hat{y}_{i}\\ &\mathcal{L}_{reg}=\frac{1}{N}\sum_{i=1}^{N}{(y_{i}-\hat{y}_{i}^{all})^{2}}\end{split}$
-The final loss function $\mathcal{L}$ includes the main regression loss (MSE) and the disentangled learning loss:
+The final loss function $\mathcal{L}$ includes the main regression loss and the disentangled learning loss:
 (13) $\begin{split}&\mathcal{L}=\mathcal{L}_{reg}+\beta\mathcal{L}_{dis}\end{split}$
 where $\beta$ controls the strength of disentanglement tasks.
-4. Experimental Settings
-In this section, we conduct extensive experiments and analyses on three subsets extracted from the real-world scientific dataset S2AG to validate DPPDCC  [^1] We will make our datasets and codes publicly available on Github..
-Particularly, we plan to answer the following research questions (RQs):
-$\bullet$ RQ1: Can DPPDCC improve the potential impact prediction task?
+4. Experiment
+In this section, we conduct extensive experiments and analyses on three subsets extracted from the real-world dataset S2AG to validate DPPDCC, aiming to answer the following research questions (RQs):
+$\bullet$ RQ1: Can DPPDCC improve the impact prediction task?
 $\bullet$ RQ2: What is the role of each component in DPPDCC?
@@
 $\bullet$ RQ4: How does DPPDCC conduct the prediction?
-4.1. Datasets
-Our dataset is constructed using S2AG (Kinney et al., 2023), a comprehensive repository housing roughly 100 million scientific publications, offering extensive metadata and unique identifiers.
-To enhance experimental efficiency and accommodate the diverse characteristics across domains, we organize these publications into subsets based on their respective fields of study.
-To establish the global citation network, we locate English papers with complete metadata (title, abstract, author IDs, and venue ID), or those of high impact despite lacking venue information.
-From the dataset, we select three fields: computer science (CS), chemistry (CHM), and psychology (PSY).
-Relevant statistics are summarized in Table 1.
-Table 1. Detailed statistics for ground-truth samples, nodes and edges of the citation network, as well as sample information of training, validation, and testing sets.
+4.1. Experimental Settings
+4.1.1. Datasets
+Table 1. Statistics for nodes and edges of the last citation network, as well as sample information of all sets.
 dataset CS CHM PSY
-ground-truth sample (paper) 2,513,197 1,818,138 1,857,277
 node paper 1,628,853 1,376,599 1,297,771
@@
 qualified pool 1,112,611 1,093,660 1,015,579
-To select samples for testing, we target papers published before the test observation point (see Figure 1 (b)).
-These papers should be written in English, have complete metadata, and include at least one reference, without imposing constraints on the number of citations received, to maintain the integrity of the data distribution.
+To establish the global citation network, we locate papers from S2AG (Kinney et al., 2023) that possess complete metadata or are high-impact despite no venue.
+To select samples for testing, we target papers published prior to the test observation point (see Figure 1b).
+These papers should have complete metadata and contain at least one reference, without any restrictions on the citations they have received.
+We randomly sample 300,000 papers from the qualified pool to form the test set.
+Additionally, we gather all the metadata associated with these papers.
+To imitate the practical scenario, the training observation point is set 5 years before the test observation point (equal to the observation time window $T$ and predicted interval $\Delta$ for non-overlapping).
+The validation observation point precedes the test observation point by 3 years.
+Given the data quality and the disciplinary differences, we select three fields from the dataset: computer science (CS), chemistry (CHM), and psychology (PSY).
+The relevant statistics are shown in Table 1.
+Table 2.
+Performance comparison of DPPDCC with baselines in MALE and LogR2.
+We divide the results into four categories: previous, fresh, and immediate papers, as well as the total.
+The best, second-best and third-best results are in red, blue, and green, respectively.
+Significant improvements over best baseline results (with a 95% confidence interval) are marked with .
+Dataset Computer Science Chemsitry Psychology
+MALE($\downarrow$) total prev fresh immed total prev fresh immed total prev fresh immed
+SciBERT 0.6904 0.6110 0.8148 0.7996 0.5868 0.5384 0.7380 0.6841 0.6131 0.5776 0.7098 0.6828
+SPECTER2 0.7482 0.6780 0.8583 0.8226 0.5832 0.5314 0.7450 0.6957 0.5935 0.5604 0.6837 0.6527
+HINTS 0.9022 0.9139 0.8838 0.9025 0.8265 0.7952 0.9246 0.9314 0.8529 0.8510 0.8582 0.8214
+CasSampling 0.5061 0.4249 0.6334 0.8535 0.4679 0.4266 0.5971 0.8389 0.5196 0.4757 0.6392 0.8066
+H2CGL 0.4579 0.3870 0.5688 0.6455 0.4337 0.3982 0.5447 0.5820 0.4727 0.4415 0.5581 0.5848
+EGCN 0.8051 0.7865 0.8341 0.8322 0.6381 0.6135 0.7150 0.6894 0.7339 0.7307 0.7427 0.6978
+Dysat 0.6297 0.5571 0.7434 0.8129 0.5412 0.5036 0.6588 0.6774 0.6207 0.5874 0.7115 0.7138
+ROLAND 0.5978 0.5290 0.7056 0.7384 0.5437 0.4951 0.6955 0.6858 0.5521 0.5003 0.6934 0.7280
+DisenGCN 0.8321 0.7729 0.9249 0.9000 0.6343 0.6022 0.7347 0.6945 0.6689 0.6466 0.7296 0.7145
+DisenHAN 0.8221 0.7483 0.9379 1.0828 0.8652 0.9072 0.7341 0.6896 0.8766 0.8358 0.9877 0.9758
+CAL 0.7094 0.6613 0.7848 0.7997 0.5733 0.5424 0.6698 0.6404 0.5535 0.5335 0.6081 0.6324
+DisC 0.6733 0.6143 0.7656 0.7953 0.5609 0.5233 0.6784 0.6613 0.5592 0.5336 0.6289 0.6505
+DIDA 0.6203 0.5403 0.7457 0.8096 0.5108 0.4707 0.6360 0.6757 0.5723 0.5389 0.6634 0.6765
+DyTed 0.6904 0.6110 0.8148 0.7996 0.5386 0.4960 0.6719 0.6784 0.5838 0.5451 0.6894 0.6986
+SILD 0.5836 0.5061 0.7051 0.7738 0.5009 0.4679 0.6040 0.6347 0.5448 0.5122 0.6335 0.6566
+DPPDCC 0.4432 0.3724 0.5540 0.6232 0.4264 0.3949 0.5250 0.5688 0.4587 0.4287 0.5405 0.5699
+#improve (%) 3.21 3.78 2.60 3.45 1.69 0.84 3.62 2.27 2.96 2.88 3.14 2.55
+LogR2($\uparrow$) total prev fresh immed total prev fresh immed total prev fresh immed
+SciBERT 0.3370 0.4255 0.0859 0.0540 0.3887 0.3976 0.0209 0.0546 0.4761 0.5458 0.0779 -0.0366
+SPECTER2 0.2337 0.2944 -0.0024 0.0125 0.3906 0.4101 -0.0067 0.0120 0.5082 0.5714 0.1425 0.0449
+HINTS -0.1227 -0.2590 -0.1743 -0.3801 -0.1283 -0.1891 -0.5866 -0.8466 -0.0198 -0.0021 -0.3622 -0.5247
+CasSampling 0.6112 0.7002 0.4153 -0.1788 0.6127 0.6331 0.3371 -0.4418 0.6200 0.6945 0.2431 -0.4616
+H2CGL 0.6812 0.7268 0.5563 0.3846 0.6573 0.6582 0.4626 0.3064 0.6836 0.7315 0.4217 0.2209
+EGCN 0.1296 0.0704 0.0290 -0.1268 0.2826 0.2080 0.0940 0.0350 0.2530 0.2654 0.0042 -0.0896
+Dysat 0.4242 0.4812 0.2320 -0.0057 0.4705 0.4499 0.2326 0.0784 0.4354 0.4881 0.0882 -0.1431
+ROLAND 0.4911 0.5337 0.3314 0.2044 0.4495 0.4518 0.1344 0.0530 0.5718 0.6572 0.1419 -0.1849
+DisenGCN 0.1026 0.1055 -0.0849 -0.1153 0.2897 0.2344 0.0502 0.0456 0.3746 0.4154 0.0549 -0.1024
+DisenHAN -0.0488 0.0421 -0.3820 -0.8766 -0.1110 -0.3875 0.0577 0.0335 -0.1186 -0.0173 -0.7948 -1.1430
+CAL 0.1859 0.2790 -0.1024 -0.2403 0.4166 0.3867 0.1753 0.1569 0.5783 0.6163 0.3238 0.1045
+DisC 0.3496 0.3926 0.1605 -0.0066 0.4259 0.4086 0.1538 0.0835 0.5606 0.6039 0.2820 0.0589
+DIDA 0.4449 0.5228 0.2295 0.0043 0.5354 0.5318 0.2855 0.0847 0.5328 0.5887 0.2005 -0.0299
+DyTed 0.3370 0.4255 0.0859 0.0540 0.4602 0.4454 0.2001 0.0821 0.5029 0.5654 0.1382 -0.0967
+SILD 0.4977 0.5809 0.2862 0.0503 0.5629 0.5539 0.3440 0.1766 0.5889 0.6444 0.2733 0.0260
+DPPDCC 0.6938 0.7349 0.5775 0.4236 0.6667 0.6577 0.5056 0.3458 0.6969 0.7380 0.4633 0.2594
+#improve (%) 1.85 1.11 3.80 10.13 1.43 9.29 12.84 1.94 0.89 9.87 17.44
+4.1.2. Baseline models
+The baseline models are mainly divided into citation/cascade prediction models, dynamic GNN models, and disentangled GNN models:
+Citation/Cascade prediction models: We apply classic and recent citation/cascade prediction models to estimate the potential impacts of papers based on content, citation graphs, or citation cascade graphs.
+SciBERT (Beltagy et al., 2019) is fine-tuned to encode the title and abstract of the paper.
+SPECTER2 (Cohan et al., 2020) focuses on document-level embeddings enhanced by a citation-aware pretrained task.
+HINTS (Jiang et al., 2021) addresses the cold-start problem by applying GCN on pseudo meta-data graphs, using a stochastic model to make predictions.
+CasSampling employs an importance-aware sampling technique to extract crucial segments of cascade graphs while preserving the temporal signals of the overall graphs.
+H2CGL (He et al., 2023) models citation network dynamics hierarchically and integrates contrastive learning for representation refinement.
+It shares a dynamic setup and a heterogeneous GNN framework similar to ours.
+Dynamic GNN models: We employ recent dynamic GNN models to extract paper representations from the dynamic graphs collected similar to our model, treating potential impact prediction as a graph regression task.
+EvolveGCN (Pareja et al., 2020) integrates the RNN structure within GNN modules to facilitate sequential dynamic updates.
+Dysat (Sankar et al., 2020) incorporates multiple self-attention modules to encode structural and temporal information separately.
+ROLAND (You et al., 2022) is a novel dynamic heterogeneous GCN model that inherits static GCN methods hierarchically and homogeneously encodes heterogeneous information using type indicators.
+Disentangled GNN models: We apply recent disentangled GNN models to extract critical representations from the multi-hop subgraphs.
+DisenGCN (Ma et al., 2019a) leverages neighborhood routing from capsule networks for disentanglement.
+DisenHAN (Wang et al., 2020) builds upon DisenGCN and is adapted to heterogeneous graphs for the recommendation.
+CAL (Sui et al., 2022) employs attention modules to estimate the causal and trivial masks for structures and attributes and encodes them with separate GCNs.
+DisC (Fan et al., 2022) estimates the causal/bias masks and further employs causal/bias-aware loss functions as well as counterfactual unbiased training.
+DIDA (Zhang et al., 2022) discovers invariant patterns within dynamic graphs and applies random variant pattern swapping as the causal intervention.
+DyTed (Zhang et al., 2023a) devises a framework combining contrastive learning and adversarial learning to disentangle time-invariant and time-varying representations proficiently.
+SILD (Zhang et al., 2023b) expands upon DIDA in the spectral domain.
+4.1.3. Implementation details
+All models are implemented with PyTorch (Paszke et al., 2019) and DGL (Wang et al., 2019)/PyG (Fey and Lenssen, 2019).
+Training is conducted on an NVIDIA A800 80GB GPU and optimized with the Adam optimizer (Kingma and Ba, 2014).
+MALE (Mean Absolute Logarithmic Error) and LogR2 (Logarithmic R-squared) are employed for evaluation.
+We run the experiments with three random seeds and report the averaged results.
+Models are developed using both training and validation sets, with the final model selected based on the best MALE on the validation set.
+Each dataset is divided into previous, fresh, and immediate papers for detailed analysis.
+We adhere to the recommended settings in the original papers/codes for all baselines.
+In our model, the hop order $k$ is set to 2, corresponding top limits $K_{1}$ and $K_{2}$ are set to 100 and 20, the number of layers $L$ is set to 4, and the co-cited/citing weight $\lambda$ and the disentanglement weight $\beta$ are set to 0.5.
+4.2. Performance Comparison (RQ1)
+To answer RQ1, we evaluate the performance of DPPDCC and other baselines. From Table 2, we have the following observations:
+(1) All models perform well on previous papers presented in the training set but experience degradation when dealing with fresh and immediate papers in the test set, particularly in terms of the LogR2.
+A substantial disparity is evident between the dynamic citation prediction models (H2CGL and ours) and other models.
+Even the best static model, CasSampling, struggles with immediate papers due to limited graph structures, tailored as it is a cascade prediction model designed for papers with sufficient citations.
+However, dynamic GNNs excel in encoding multi-hop information, compensating for the absence of cascade structures.
+Additionally, models like SciBERT, SPECTER2, and H2CGL can leverage prior knowledge, such as content information or bibliometric theory, to mitigate these challenges.
+(2) Both dynamic and heterogeneous graphs are important for our task.
+On one side, temporal information within dynamic graphs reveals the evolving trends, underscoring its critical role.
+DIDA, DyTed, and SILD, which consider dynamic graphs, stand out among all disentangled GNNs.
+On the other side, leveraging heterogeneous information can bolster the performance of dynamic GNNs, as demonstrated by ROLAND.
+However, proper modeling of heterogeneous information is essential; otherwise, it may result in performance degradation, as evidenced by DisenHAN compared to its backbone, DisenGCN.
+The top-performing baseline, H2CGL, hierarchically encodes both dynamic and heterogeneous graphs, effectively adapting to the dynamic setting and achieving the best performance among baselines.
+(3) While disentanglement methods yield performance enhancements, their adaptation to dynamic heterogeneous graphs is crucial for achieving optimal flexibility.
+Remarkably, SILD achieves superior outcomes compared to ROLAND in Chemistry, even when utilizing only homogeneous graphs.
+Other dynamic disentangled GNNs also showcase comparable performance, with DyTed outperforming its backbone, Dysat.
+Additionally, in the densest dataset, Psychology, disentangled GNNs demonstrate commendable performance, with CAL surpassing both CasSampling and ROLAND for fresh and immediate papers.
+However, in other scenarios, aside from dynamic disentangled GNNs, other methods struggle to compete with task-specific models and dynamic GNNs.
+(4) Our proposed model surpasses all baselines by effectively modeling the dynamic heterogeneous graph and integrating task-adjusted disentanglement.
+This results in an average improvement of approximately 3% in MALE across all categories, alongside a significant advancement of over 10% in LogR2 for fresh and immediate papers.
+Initially, compared to traditional citation/cascade prediction models, it demonstrates superior generalization and practicability through the fusion of content and context information within citation networks, coupled with its disentanglement capabilities.
+Furthermore, we leverage the unique properties of citation networks to encode complex high-order structural and temporal information, thereby outperforming common dynamic GNNs.
+Crucially, our model introduces interpretability from disentanglement models without sacrificing performance and further adapts to the task for enhanced improvement, thus surpassing the best dynamic citation prediction model, H2CGL.
+4.3. Ablation Test (RQ2)
+To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules.
+For CGE, (-co-weight) removes the proposed co-cited/citing strengths, (-CompGAT) replaces the CompGAT with GATv2 (Brody et al., 2022), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes.
+For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP.
+Table 3. Results of ablation studies in all datasets.
+model CS CHM PSY
+MALE LogR2 MALE LogR2 MALE LogR2
+complete 0.4432 0.6938 0.4264 0.6667 0.4587 0.6969
+-co-weight 0.4517 0.6855 0.4623 0.6106 0.4625 0.6953
+-CompGAT 0.4620 0.6842 0.4532 0.6202 0.4604 0.6921
+-readout 0.4982 0.6432 0.4404 0.6564 0.4867 0.6720
+-diffusion 0.4632 0.6801 0.4531 0.6231 0.4621 0.6914
+-conformity 0.4655 0.6814 0.4767 0.5799 0.4646 0.6935
+-orthog 0.4576 0.6870 0.4560 0.6168 0.4597 0.6958
+-disen 0.4597 0.6937 0.4696 0.5990 0.4830 0.6738
+From Table 3, we observe that: (1) Since removing any module leads to a performance decrease, all modules play crucial roles in DPPDCC.
+(2) For CGE, the removal of type-specific attention snapshot readout may cause a notable performance drop.
+That could be attributed to the presence of excessive noisy information within multi-hop snapshots, which necessitates the filtering of aggregated information for better concentration.
+CompGAT effectively captures distinct comparative information, and the incorporation of co-cited/citing strengths further complements the multi-hop information beyond the local neighborhood.
+(3) Eliminating any disentanglement task adversely affects performance.
+Moreover, an evident decline is observed after removing the whole disentanglement module.
+Conformity rises as the paramount task among all disentanglement tasks.
+It focuses on the Matthew effect, which is widely pervasive and readily discernible in reality.
+4.4. Hype-parameters Test (RQ3)
+(a) $L$
+(b) $\lambda$
+(c) $\beta$
+Figure 3. Results of hyper-parameters test in terms of MALE.
+To address RQ3, we aim to explore the model’s sensitivity to essential hyper-parameters, such as $L$, $\lambda$, and $\beta$, which respectively control the order/hop information of CGE, the importance of co-cited/citing strengths, and the weight of the disentanglement tasks relative to the main task.
+The results about MALE with varying settings are depicted in Figure 3.
+We observe that: (1) When the layer number $L$ continues to increase, the performance will first increase and then decrease.
+Since we stack the structural and temporal encoders simultaneously, their interaction is more complicated and may also face the over-smoothing problem.
+(2) The co-cited/citing weight $\lambda$ and disentanglement weight $\beta$ demonstrate greater stability compared to $L$, especially in Psychology.
+Selecting the proper hyper-parameters can further improve the performance.
+4.5. Disentanglement Visualization (RQ4)
+(a) time trend
+(b) group detail
+(c) value bin
+Figure 4.
+Visualization of disentangled value proportions in CS.
+(a) displays the trend that evolved with the publication time.
+(b) demonstrates the detailed composition of papers categorized into previous, fresh, and immediate ones.
+(c) is binning the samples based on the predicted values.
+To address RQ4, we provide visualizations of the disentanglement composition and the conformity representation distribution. Additional information can be found in the supplementary materials.
+4.5.1. Disentanglement Composition Visualization
+In Figure 4, we visualize samples with all positive values in the test set of Computer Science.
+Papers containing negative disentangled values typically exhibit nearly no increase in citations.
+Remarkably, these papers consistently feature negative values because their contribution values are the lowest across all perspectives, demonstrating our model’s rationality.
+In the stacked plot associated with time in Figure 4a, a clear trend emerges: the contribution proportion gradually increases, while the conformity proportion undergoes a significant drop.
+This trend is logical, as older papers tend to accumulate substantial citations, thereby attracting more researchers to cite them due to inertia.
+Additionally, the diffusion proportion appears more stable over time, as it represents potential hidden properties within diffusion.
+This stability becomes more pronounced when papers are categorized into previous, fresh, and immediate ones, as shown in Figure 4b.
+Diffusion proportions remain steady in both mean and standard deviation.
+Moreover, contribution emerges as the most critical perspective, with its dominant and variable proportion.
+It holds particular significance for fresh and immediate papers.
+For immediate papers receiving almost no citations, both the diffusion and conformity perspectives struggle to offer predictive insights.
+However, the contribution can distinguish their potential in advance based on the citation context.
+Furthermore, in Figure 4c, when we bin the samples based on the predicted values, we observe a descending trend for conformity and a growing trend for contribution.
+Notably, conformity is on par with contribution within the lowest bin group, while diffusion accounts for the largest proportion.
+However, for papers with higher increments, contribution plays the most significant role and surpasses all other perspectives.
+4.5.2. Disentanglement Representation Visualization
+(a) total
+(b) prev
+(c) fresh
+(d) immed
+Figure 5. Results of T-SNE visualization for representations from conformity perspective in CS. Various colors correspond to the binning labels assigned to the samples.
+To validate that the disentangled perspectives learn distinct aspects of the features, we visualize the representations extracted from the conformity encoder using T-SNE (Pedregosa et al., 2011).
+Figure 5 illustrates a consistent and structured distribution of points across all figures.
+This indicates our model effectively captures group discrepancies across various binning labels for papers published at different times.
+Variations in the distribution of binning labels are evident across different categories due to differences in cumulative citations over time.
+Fresh papers published within the last five years, particularly immediate ones, tend to fall under the lowest accumulated citation label (red), dominating a significant area.
+Despite this, discernible boundaries persist between distinct binning labels within all these categories.
+This suggests the model’s capacity to generalize and transfer learned group differences to unseen samples, showcasing its potential applicability in real-world scenarios.
+5. Conclusion
+In this study, we propose a novel model named DPPDCC that disentangles the potential impacts of papers into diffusion, conformity, and contribution values.
+To better align with real-world conditions, we reformulate the problem by partitioning the datasets based on the observation time points without excluding lowly-cited papers.
+Further, DPPDCC comprises two principal components: Citation-aware GNN Encoder (CGE) and Popularity-aware Disentanglement Module (PDM).
+Given the dynamic heterogeneous graph of a target paper, CGE employs CompGAT to encode the comparative and co-cited/citing information between papers and utilizes Type-specific Attention Snapshot Readout to aggregate the information from each snapshot evolutionarily.
+Subsequently, PDM identifies amplifying effects in information diffusion and addresses the Matthew effect stemming from accumulated citations from a collective conformity perspective.
+Finally, it constrains each perspective to address distinct aspects to preserve genuine contributions.
+Experimental results on three datasets demonstrate that DPPDCC significantly outperforms alternative approaches for previous, fresh, and immediate papers.
+Disentanglement visualization further demonstrates that DPPDCC can reliably perform predictions.
+Acknowledgements.
+This work is supported by the National Natural Science Foundation of China (72204087, 72234005, 72104212), the Shanghai Planning Office of Philosophy and Social Science Youth Project (2022ET0001), ”Chen Guang” project supported by Shanghai Municipal Education Commission and Shanghai Education Development Foundation (23CGA28), Shanghai Pujiang Program (23PJC030), the Fundamental Research Funds for the Central Universities, 2024 Innovation Evaluation Open Fund, Fudan University (CXPJ2024006).
+We also appreciate the constructive comments from the reviewers.
+Appendix A Extended Related Work
+A.1. Citation/Cascade prediction
+Citation prediction is a vital sub-task of automatic academic evaluation as it enables the estimation of potential impact.
+It is closely examined within cascade prediction tasks, which share analogous graph structures and objectives (Zhou et al., 2021).
+The primary goal of cascade prediction is to anticipate a post’s popularity based on user interactions.
+Similarly, in the context of citation networks, the paper can serve as the post, while the authors can serve as the users.
+Citation count prediction focuses on individual papers’ content and distinctive features, heavily relying on comprehensive feature extraction techniques.
+In contrast, cascade prediction delves into the intricate relationships between various entities, including papers and authors, directly exploiting graph structure information.
+Their approaches can be broadly classified into three categories: stochastic models, feature-based models, and deep learning models.
+Stochastic models can predict future citation counts by fitting the curve of citation counts (Glänzel and Schubert, 1995), following Zipf-Mandelbrot’s law (Silagadze, 1997).
+Recently, machine learning models have shown promising results by utilizing manually extracted features from various meta-data (e.g.
+content, author, venue) (Yan et al., 2011; Yu et al., 2014; Ruan et al., 2020; Dong et al., 2015; Guille et al., 2013; Tatar et al., 2014).
+More presently, deep neural networks have dominated this task by applying advanced Natural Language Processing (NLP) and Computer Vision (CV) methods to extract abstractive representations from paper content (Abrishami and Aliakbary, 2019; Huang et al., 2022; Cohan et al., 2020; Xue et al., 2023).
+For cascade prediction, graph embeddings, sequence models, and GNN models are employed to extract structural information from the underlying graph, and then temporal information is encoded with sequence models (Li et al., 2017; Cao et al., 2017; Yang et al., 2022; Chen et al., 2022b; Xu et al., 2022; Cheng et al., 2023).
+However, most existing models for citation and cascade prediction exhibit sub-optimal performance.
+They fail to fully exploit the valuable information presented in paper content and scientific context within the citation network simultaneously.
+Furthermore, their strict data selection and random data splitting strategies may hinder their practicality.
+Many valuable samples will be filtered, and the temporal information within the dynamic context will be neglected.
+Thus, models in this setting are more prone to face distribution shift issues.
+In contrast, our study revises the problem setting of impact prediction in the dynamic context (He et al., 2023; Yan et al., 2024).
+We ensure non-overlapping training and testing observation windows, retaining the real data distribution without sample filtering.
+We also utilize both content and dynamic context, introducing disentangled representation learning to provide a more practical impact estimation.
+A.2. Dynamic/Heterogeneous Graph Neural Networks
+GNNs (Sperduti and Starita, 1997; Gori et al., 2005; Scarselli et al., 2008; Wu et al., 2020) are widely applied to handle non-Euclidean data like graphs.
+Recently, research interest has surged in dynamic graphs, which incorporate temporal information, as well as in heterogeneous graphs, which involve multiple node types or edge types.
+These types of graphs are more prevalent in practical scenarios and encapsulate more intricate information compared to static homogeneous graphs, driving the formulation and refinement of diverse methodologies within the academic realm.
+The dynamic graph can be divided into multiple snapshots at different time points.
+Previous models like DGCN (Manessi et al., 2020), Dysat (Sankar et al., 2020), and ROLAND (You et al., 2022) typically encode the snapshots with static GCN as the structural encoder, followed by sequential models like Recurrent Neural Networks (RNN) or Transformer Encoder.
+Moreover, certain models incorporate GCN into RNN models by configuring the parameters within RNN cells as GCN parameters (Pareja et al., 2020).
+Within heterogeneous graphs, on one side, R-GCN serves as a prominent model with its separate message-passing architecture (Schlichtkrull et al., 2018).
+It conducts independent message passing within different relations and aggregates them to update node representations.
+On the other side, the relation-learning method transforms heterogeneous graphs into homogeneous ones, while retaining node and edge type indicators to preserve heterogeneous information (Hu et al., 2020; You et al., 2022).
+An example of this technique is HGT (Hu et al., 2020), which utilizes multi-head attention modules with specific parameters tailored to different node and edge types, effectively modeling the complexity of heterogeneous graphs.
+Citation networks can be regarded as dynamic and heterogeneous graphs as they evolve over time and encompass diverse entities such as papers, authors, and venues.
+Addressing the complexities within this network involves two primary challenges: (1) How to encode the temporal target-centric information within a single snapshot?
+(2) How to model distinct characteristics of information diffusion in citation networks?
+To overcome these challenges, our study proposes a Citation-aware GNN Encoder that effectively models both dynamic and heterogeneous graphs of target papers.
+These graphs are constructed from the citation network annually and incorporate multiple metadata nodes beyond the papers themselves.
+(1) To encode the temporal target-centric information within a single snapshot, we employ a type-aware attention readout.
+It aggregates snapshot information considering the paper type and publication time.
+(2) To model distinct characteristics of information diffusion in citation networks, we propose a novel GCN module CompGAT.
+This module focuses on capturing comparisons and co-cited/citing information between papers, drawing inspiration from bibliometrics theories (Wu et al., 2019).
+A.3. Disentangled Representation Learning
+Disentangled representation learning, aiming to discern and separate fundamental explanatory factors (Bengio et al., 2013), stands as a pivotal method within deep learning.
+It strives to produce resilient, manageable, and interpretable representations.
+Existing research predominantly focuses on the field of CV, often employing methods like Variational Autoencoders (VAEs) and Generative Adversarial Networks (GANs) to disentangle relevant latent factors (Ma et al., 2019b; Locatello et al., 2019; Cai et al., 2019; Denton et al., 2017; Lee et al., 2021).
+Researchers explore techniques to align distinct representations with specific factors through explicit and implicit guidance, employing approaches like direct labeling, loss penalties, and regularization.
+For instance, DR-GAN (Tran et al., 2017) utilizes an adversarial network to generate diverse facial poses, aiming to disentangle pose information from facial attributes.
+In NLP, disentangled representation learning finds applications in various tasks like generation.
+For example, John et al.
+(John et al., 2019) leverage VAE to disentangle text semantics from its writing style, enabling independent manipulation of text semantics and style.
+This separation allows precise control over content and style in text generation.
+Recently, there has been a growing interest in applying disentangled representation learning to graph learning tasks.
+Given that graph data encompasses both structural and attributed information, numerous new methodologies have emerged to address this complexity.
+In the feature space, approaches inspired by capsule networks have emerged to partition node features into multiple hidden channels for implicit disentangled modeling (Ma et al., 2019a).
+Some methods introduce additional constraints to enable effective separation (Liu et al., 2020).
+Moving into the structural space, subsequent models strive to factorize the input graphs into multiple subgraphs to facilitate distinct message-passing strategies (Yang et al., 2020; Li et al., 2021).
+Recent models have expanded their scope to isolate causal and biased information presented in both the graph structure and node features, with a particular emphasis on causal effects (Sui et al., 2022; Fan et al., 2022).
+These models commonly employ techniques to generate masks within the adjacency matrix and node features.
+This process aims to identify the crucial elements of the graph within the structural and attributed spaces.
+In addition, certain studies extend disentangled GNNs to specialized scenarios such as recommendation systems, handling intricate graph structures in heterogeneous or dynamic graphs (Wang et al., 2020; Zhang et al., 2022; Wen et al., 2022; Zhang et al., 2023b, a).
+Besides the masking strategy considering solely the graph structures and node features, they integrate classical methods like VAE and adversarial learning from CV and NLP to handle task-specific information from a broader perspective.
+In contrast to previous studies, our approach acknowledges that the potential impact within the citation network is influenced not only by the paper’s contribution but also by factors related to its popularity.
+These popularity factors encompass the amplifying effect owing to the highly-cited paper nodes in information diffusion and the manifestation of the Matthew effect through collective conformity.
+Following previous disentangled studies in recommendations (Zheng et al., 2021; Chen et al., 2022a), we consider the final predicted values as mixed results of multiple factors and introduce auxiliary tasks for separate encoding of different perspectives.
+Based on the dynamic heterogeneous graph, we disentangle the potential impact into diffusion, conformity, and contribution values, seeking a more reasonable estimation in explicit modeling.
+Specifically, we devise dedicated auxiliary tasks to extract values associated with these respective factors.
+Appendix B Notation Table
+Table 4. The complete notation table of our paper.
+Notation Description
+$N$ All predicted samples
+$G$ Input dynamic heterogeneous graph
+$t$ The $t$-th time point of the whole time window
+$G^{t}$ The $t$-th subgraph of the input graph
+$\epsilon^{t}$, $\nu^{t}$ Set of nodes and edges at $t$-th time point
+$\phi$, $\varphi$ Type map function of nodes and edges
+$U$, $R$ Set of node types and edge types
+$p$ The paper node type
+$\bm{h}_{ti,u}^{(l)}$ The $i$-th node hidden representation of type $u$ at $t$-th time step in $l$-th layer
+$\bm{H}_{s}$ Snapshot representation matrix containing all $h_{ti,s}$ of the $i$-th sample
+$AGG$ Aggregation functions of update representations from multiple relationships (edges)
+$f_{r}$ Message passing method (module) of the edge $r$
+$g_{r}$ Subgraph of the edge $r$
+$\bm{h}_{r_{src}}$, $\bm{h}_{r_{dst}}$ The source and target node representations of the edge r
+$\bm{W}$ The model parameters
+$\bm{C}$ Normalized co-cited/citing strength matrix
+$\bm{o}_{i}$ The final encoded representation of the $i$-th target sample before disentanglement
+$\bm{o}_{i}^{dif}$, $\bm{o}_{i}^{con}$ The representation of the $i$-th sample’s diffusion and conformity view
+$\bm{z}_{i}^{ori\vee pos\vee neg}$ The $i$-th sample’s projected representation of the original, positive and negative view
+$\mathcal{L}_{dif}$ Objective to disentangle value of diffusion view
+$\mathcal{L}_{con}$ Objective to disentangle value of conformity view
+$\mathcal{L}_{ort}$ Orthogonal constraint to ensure the non-overlapping information of disentangled views
+$\mathcal{L}_{reg}$ Objective to predict the citation increments as the main task
+$L$ The number of stacked graph encoder
+$\lambda$ The importance of co-cited/citing strengths
+$\beta$ The weight of the whole disentanglement task
+Appendix C Complexity Analysis
+The primary complexity lies in the graph encoder component of our model, CGE.
+Within each layer of CGE, multiple R-GCNs operate on snapshots captured at different time points, alongside a snapshot representation updater and a multi-layer transformer encoder.
+For edges apart from ”cites” and ”is cited by” in R-GCNs, a simple MLP ($W^{u}\in\mathbb{R}^{d\times d}$) serves as learnable parameters for GIN, with a complexity of $\mathcal{O}\left(LT(|R|-2)(|\nu^{t}_{r_{src}}|d^{2}+|\epsilon^{t}_{r}|d)\right)$.
+For ”cites” and ”is cited by” edges, the backbone of CompGAT, GATv2, incurs a complexity of $\mathcal{O}\left(LT(|\nu^{t}_{p}|d^{2}+|\epsilon^{t}_{p}|d)\right)$.
+Due to the modeling of concatenated source and target representations, it further requires $\mathcal{O}\left(LT|\epsilon^{t}_{p}|d^{2})\right)$ time, resulting in an overall complexity of $\mathcal{O}\left(LT(|\nu^{t}_{p}|+|\epsilon^{t}_{p}|)d^{2}\right)$.
+The snapshot representation updater, functioning as a cross-attention module, operates with a complexity of $\mathcal{O}\left(LT(|\nu^{t}_{p}|d+d^{2})\right)$, while the temporal encoder exhibits a complexity of $\mathcal{O}\left(L(T^{2}d+Td^{2})\right)$.
+These modules can be readily disregarded compared to R-GCNs.
+For PDE, which solely models the final target representation using multiple MLPs ($\mathcal{O}\left(d^{2}\right)$), its contribution can also be omitted.
+Thus, the resulting complexity of DPPDCC is $\mathcal{O}\left(LT(|\nu^{t}|d^{2}+|\epsilon^{t}_{p}|d^{2}+|\epsilon^{t}|d)\right)$.
+The final time complexity depends on the sampling strategy for graph size and the sparsity of paper node subgraphs.
+Since citation graphs are sparse and only a small portion of the entire graphs is sampled, computations are confined mainly within the limited paper subgraphs.
+Therefore, DPPDCC is comparable to other dynamic heterogeneous GNNs.
+Appendix D Experimental Details
+D.1. Datasets
+Table 5. Detailed statistics for ground-truth samples, nodes, and edges of the citation network, as well as sample information of training, validation, and testing sets.
+dataset CS CHM PSY
+ground-truth sample 2,513,197 1,818,138 1,857,277
+node paper 1,628,853 1,376,599 1,297,771
+author 1,598,925 1,946,073 1,585,595
+venue 12,524 8,389 13,775
+time 147 187 202
+edge cite 11,534,431 10,382,698 13,401,112
+write 5,123,460 6,355,630 4,813,135
+publish 1,566,442 1,318,158 1,230,440
+have 1,628,853 1,376,599 1,297,771
+sample train (2012) 183,105 227,304 219,499
+val (2014) 224,086 256,403 250,229
+test (2017) 300,000 300,000 300,000
+qualified pool 1,112,611 1,093,660 1,015,579
+Our dataset is constructed using S2AG (Kinney et al., 2023), a comprehensive repository housing roughly 100 million scientific publications, offering extensive metadata and unique identifiers.
+To enhance experimental efficiency and accommodate the diverse characteristics across domains, we organize these publications into subsets based on their respective fields of study.
+To establish the global citation network, we locate English papers with complete metadata (title, abstract, author IDs, and venue ID) or those of high impact despite lacking venue information.
+From the dataset, considering the data quality and the disciplinary differences, we select three fields: computer science (CS), chemistry (CHM), and psychology (PSY).
+Their total number of papers and the completeness of their metadata are the highest among all fields.
+Additionally, the differences in citation patterns resulting from varying research paradigms, such as variations in citation half-lives (Mendoza, 2021), render these fields appropriate representative samples for demonstrating citation differences across disciplines.
+Relevant statistics are summarized in Table 5.
+To select samples for testing, we target papers published before the test observation point.
+These papers should be written in English, have complete metadata, and include at least one reference without imposing constraints on the number of citations received to maintain the integrity of the data distribution.
 From the eligible pool, we randomly sampled 300,000 papers to compose the test set.
@@
 Moreover, we meticulously categorize the papers in the results for thorough and detailed analysis.
-4.2. Baseline Models
+Table 6. Detailed statistics of single input heterogeneous subgraph for DPPDCC.
+dataset CS CHM PSY
+phase train val test train val test train val test
+nodes paper 204 222 257 149 187 254 361 401 460
+author 446 501 613 505 646 907 1015 1173 1431
+venue 72 80 92 35 41 50 92 104 123
+time 23 24 26 22 24 27 32 34 37
+edges cites 740 827 1016 480 662 1036 1607 1808 2085
+writes 599 674 827 673 864 1218 1388 1598 1933
+publishes 165 183 216 135 168 228 311 350 408
+In Table 6, we present detailed statistics of the input data for DPPDCC to demonstrate the sufficiency of information within multi-hop heterogeneous subgraphs.
+For simplicity, we average the node and edge counts by time and sample.
+It is evident that the Psychology dataset is the densest, while the Computer Science and Chemistry datasets are comparable.
+D.2. Descriptive Statistics
+To validate the rationale behind the disentangled factors used in our work, we visualize the distribution of citation increments between papers potentially influenced by corresponding popularity factors and other normal papers.
+For the information diffusion effect, it is posited that documents cited by highly-cited papers are more likely to attract attention from other papers during dissemination, thereby accruing more citations.
+Conversely, for the collective conformity effect, it is argued that the cumulative citations received by a paper can reflect its partial reputation and consequently increase its likelihood of being cited.
+Highly-cited papers mentioned in both contexts are classified using a Pareto distribution method: papers providing over 80% of the total cumulative citations are deemed highly-cited.
+The papers cited by these highly-cited papers represent focused samples of the information diffusion effect, while the highly-cited papers themselves serve as focused samples of the collective conformity effect.
+(1) In Figure 6, the significant deviation among different categories is evident, with papers cited by highly-cited ones generally receiving more citations.
+Notably, the CS dataset exhibits the most distinct pattern, with the left peak nearly disappearing, resulting in an overall distribution closer to a normal distribution.
+The model aims to extract these diffusion-related factors to model the information diffusion effect independently.
+(2) Figure 7 illustrates the difference in attracting new citations between papers with accumulated citations and those with fewer citations, indicating that the collective conformity effect is more pronounced than the information diffusion effect.
+Notably, the distribution of focused samples in the CS dataset has essentially normalized, while other datasets also display significant peak differences.
+The model is expected to separately model the conformity effect, enabling it to focus more on aspects unrelated to existing cumulative citations and align more closely with contribution-related information.
+(a) CS
+(b) CHM
+(c) PSY
+Figure 6. Citation increment distribution of normal samples and focused samples cited by highly-cited papers for diffusion perspective.
+(a) CS
+(b) CHM
+(c) PSY
+Figure 7. Citation increment distribution of normal samples and focused highly-cited samples for conformity perspective.
+D.3. Baseline Models
 The baseline models are mainly divided into citation/cascade prediction models, dynamic GNN models, and disentangled GNN models:
@@
     SciBERT (Beltagy et al., 2019) is a BERT-based model pretrained on scientific corpus. We fine-tune it to encode paper titles and abstracts, predicting with the [CLS] token for citation prediction.
+    SPECTER2 (Cohan et al., 2020) generates document-level embeddings of scientific documents by pretraining a BERT-based model on relatedness signals extracted from the citation graph.
+Instead of fine-tuning the model’s adapters, we fully fine-tune the basic SPECTER2 for better performance.
     HINTS (Jiang et al., 2021) tackles the cold-start problem by employing R-GCN on pseudo meta-data heterogeneous graphs.
@@
 To integrate content information from PLMs, we replace node ID embeddings and adjust the model to encode subgraphs instead of the entire graph, mitigating memory issues.
-    MUCas (Chen et al., 2022) first uses time-aware sampling to extract dynamic graph snapshots and generate node embeddings from the graph structure.
-Then, it utilizes MUG-caps, a hierarchical capsule network, to encode snapshots in different levels (order, node, and graph) and aggregate the final representations with the attention mechanism.
-Dynamic GNN models: We employ recent dynamic GNN models to extract paper representations from the dynamic graphs collected the same as our model, taking the potential impact prediction as a graph regression task.
+    CasSampling (Cheng et al., 2023) employs an importance-aware sampling method to extract critical components of cascade graphs based on outdegree, retaining the time series of the global graphs as node features.
+Subsequently, an attention aggregator is used to encode the global-level time flow.
+    H2CGL (Cheng et al., 2023), the state-of-the-art citation prediction model, shares a similar dynamic setting and heterogeneous GNN framework with our approach.
+It encodes the dynamics of the citation network hierarchically and designs message-passing modules that consider the unique characteristics of citation networks.
+Additionally, it incorporates graph contrastive learning to refine representations, encouraging the model to encode potential citation information unrelated to specific topics.
+Dynamic GNN models: We employ recent dynamic GNN models to extract paper representations from the dynamic graphs collected in the same way as our model, taking the potential impact prediction as a graph regression task.
 Similar to HINTS, we adapt these models to concentrate on target-centric subgraphs rather than the entire graph.
@@
 It then applies random variant pattern swapping in both temporal and spatial dimensions as the causal intervention.
-4.3. Implementation Details
-All baselines and our proposed model are implemented with PyTorch and DGL/PyG.
+    DyTed (Zhang et al., 2023a) designs a temporal-clips contrastive learning task along with a structural contrastive learning task to effectively disentangle time-invariant and time-varying representations.
+It further employs an adversarial learning framework with a disentanglement-aware discriminator for enhancement.
+EGCN and Dysat are utilized as the backbone dynamic GNNs of DyTed, with the best performance reported using Dysat.
+    SILD (Zhang et al., 2023b) addresses distribution shifts in both dynamic and spectral domains.
+It builds upon DIDA with spectral transformation and proposes a disentangled spectrum mask to capture and utilize invariant and variant spectral patterns.
+D.4. Implementation Details
+All baselines and our proposed model are implemented using PyTorch and DGL (Wang et al., 2019)/PyG (Fey and Lenssen, 2019).
 They are trained on an NVIDIA A800 80GB GPU and optimized with the Adam optimizer (Kingma and Ba, 2014).
-As for the baselines, we implement them directly using or referring to the official source codes.
-We develop the models with the training and validation sets and choose the final model with the best main metric on the validation set.
+For the baselines, we directly use or refer to the official source codes.
+Experiments are conducted with three random seeds, and the averaged results are reported.
+Models are developed using the training and validation sets, with the final model selected based on the best main metric on the validation set.
 Each dataset is divided into previous papers (included in the training set), fresh papers (new additions in the test set), and immediate papers (published at the test time point) for detailed analysis.
-We adhere to the recommended settings for all baselines.
-In our model across all datasets, the hop order $k$ is set to 2, corresponding top limits $K_{1}$ and $K_{2}$ are set to 100 and 20, the learning rate is set to 1e-4, the number of layers $L$ is set to 4, the co-cited/citing weight $\lambda$ is set to 0.5, and the categories of equal frequency bins $M$ are set to 5.
-Additionally, the batch size is set to 64 for CS and CHM, and 32 for PSY, the disentanglement weight $\beta$ is set to 0.5 for CS and PSY, and 1.0 for CHM, and the hidden dimension $d^{h}$ is set to 128 for CS and PSY, and 192 for CHM.
-4.4. Evaluation Metric
+We follow the recommended settings from the original papers or open-source codes for all baselines.
+In our model across all datasets, the hop order $k$ is set to 2, corresponding top limits $K_{1}$ and $K_{2}$ are set to 100 and 20, the learning rate is set to 1e-4, the number of layers $L$ is set to 4, the co-cited/citing weight $\lambda$ and the disentanglement weight $\beta$ are set to 0.5, and the categories of equal frequency bins $M$ are set to 5.
+Additionally, the batch size is set to 64 for CS and CHM, and 32 for PSY, and the hidden dimension $d$ is set to 128 for CS and PSY, and 192 for CHM.
+D.5. Evaluation Metric
 In our evaluation, we utilize two metrics: MALE (Mean Absolute Logarithmic Error) and LogR2 (Logarithmic R-squared).
@@
 Regarding LogR2, it evaluates the coefficient of determination, calculated as the proportion of the dependent variable’s variation predictable from the independent variable, following logarithm transformation for both variables:
-(18) $R^{2}=1-\frac{\sum_{i=1}^{N}{(y_{i}-\hat{y}_{i})^{2}}}{\sum_{i=1}^{N}{(y_{i}-\frac{1}{N}\sum_{i=1}^{N}y_{i})^{2}}}$
+(15) $LogR^{2}=1-\frac{\sum_{i=1}^{N}{(y_{i}-\hat{y}_{i})^{2}}}{\sum_{i=1}^{N}{(y_{i}-\frac{1}{N}\sum_{i=1}^{N}y_{i})^{2}}}$
 LogR2 assumes a negative value when the model’s predictions fall short of the data’s mean value, indicating an inability to capture the underlying trend within the dataset.
@@
 Consequently, by integrating MALE and LogR2, we holistically gauge the model’s accuracy in direct predictions for individual samples and its aptness in capturing broader trends across groups.
-5. Results
-5.1. Performance Comparison (RQ1)
-To answer RQ1, we evaluate the performance of DPPDCC and other baselines.
-The comparison results are reported in Table 2 (Computer Science), 3 (Chemistry), and 4 (Psychology).
-We have the following observations:
-Table 2.
-Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Computer Science.
-We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total.
-The best results are shown in bold, and the second results are underlined.
-Significant improvements over best baseline results are marked with .
-CS model all prev fresh immed
-[…]
-C&C SciBERT 0.6847 0.3558 0.6077 0.4424 0.8054 0.1108 0.7860 0.0843
-HINTS 0.6799 0.2773 0.6429 0.2028 0.7378 0.2268 0.7387 0.1718
-MUCas 0.5079 0.6063 0.4331 0.6757 0.6250 0.4352 0.8261 -0.1079
-Dynamic EGCN 0.8157 0.1174 0.8008 0.0531 0.8392 0.0210 0.8387 -0.1460
-Dysat 0.6248 0.4355 0.5489 0.5029 0.7437 0.2321 0.8073 0.0089
-ROLAND 0.5814 0.5245 0.5166 0.5634 0.6828 0.3764 0.7157 0.2511
-Disen DisenGCN 0.8215 0.1428 0.7816 0.1022 0.8841 0.0206 0.8485 0.0236
-DisenHAN 0.8138 0.0054 0.7411 0.1003 0.9276 -0.3220 1.0686 -0.8032
-CAL 0.7323 0.2298 0.6859 0.2264 0.8050 0.0768 0.7940 0.0427
-DisC 0.6687 0.3434 0.6144 0.3908 0.7536 0.1472 0.7900 -0.0735
-DIDA 0.5941 0.4703 0.5018 0.5642 0.7387 0.2391 0.8181 -0.0207
-ours DPPDCC 0.4473 0.6871 0.3783 0.7236 0.5553 0.5754 0.6035 0.4512
-#improve (%) 11.93 13.32 12.65 7.09 11.15 32.22 15.68 79.67
-Table 3.
-Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Chemistry.
-We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total.
-The best results are shown in bold, and the second results are underlined.
-Significant improvements over best baseline results are marked with .
-CHM model all prev fresh immed
-[…]
-C&C SciBERT 0.5451 0.4648 0.4987 0.4709 0.6903 0.1477 0.6452 0.1653
-HINTS 0.6024 0.3795 0.5848 0.3119 0.6574 0.2253 0.6449 0.1633
-MUCas 0.4691 0.6168 0.4314 0.6306 0.5869 0.3624 0.8073 -0.3479
-Dynamic EGCN 0.6424 0.2684 0.6177 0.1913 0.7196 0.0791 0.6976 0.0035
-Dysat 0.5442 0.4573 0.5029 0.4411 0.6733 0.1998 0.6846 0.0582
-ROLAND 0.5001 0.5410 0.4592 0.5374 0.6280 0.2942 0.6406 0.1800
-Disen DisenGCN 0.6405 0.2735 0.6079 0.2151 0.7427 0.0332 0.6987 0.0387
-DisenHAN 0.6132 0.3793 0.5817 0.3683 0.7118 0.0630 0.7313 -0.1466
-CAL 0.5864 0.3928 0.5541 0.3707 0.6873 0.1158 0.6535 0.1135
-DisC 0.5393 0.4770 0.5037 0.4578 0.6505 0.2387 0.6409 0.1719
-DIDA 0.4978 0.5760 0.4642 0.5689 0.6028 0.3587 0.6426 0.1669
-ours DPPDCC 0.4355 0.6503 0.4071 0.6329 0.5243 0.5042 0.5802 0.3247
-#improve (%) 7.16 5.44 5.64 0.36 10.67 39.13 9.42 80.39
-Table 4.
-Experimental results of performance comparison of DPPDCC with baselines in MALE and LogR2 within Psychology.
-We divide the results into four categories: previously, freshly, and immediately-published papers, as well as the total.
-The best results are shown in bold, and the second results are underlined.
-Significant improvements over best baseline results are marked with .
-PSY model all prev fresh immed
-[…]
-C&C SciBERT 0.6395 0.4308 0.6025 0.5041 0.7402 0.0068 0.7116 -0.1095
-HINTS 0.7749 0.1988 0.7799 0.2007 0.7614 -0.0259 0.7064 -0.0804
-MUCas 0.5207 0.6175 0.4804 0.6846 0.6306 0.2669 0.7708 -0.3521
-Dynamic EGCN 0.7424 0.2305 0.7375 0.2469 0.7560 -0.0388 0.7102 -0.1386
-Dysat 0.6861 0.2998 0.6607 0.3342 0.7555 -0.0170 0.7439 -0.2377
-ROLAND 0.5643 0.5540 0.4996 0.6601 0.7406 0.0429 0.7926 -0.3765
-Disen DisenGCN 0.6557 0.3945 0.6316 0.4364 0.7213 0.0758 0.7081 -0.0846
-DisenHAN 0.7763 0.2455 0.8077 0.2187 0.6906 0.1387 0.6476 0.0522
-CAL 0.5631 0.5646 0.5469 0.5975 0.6073 0.3256 0.6285 0.1166
-DisC 0.5415 0.5915 0.5159 0.6343 0.6113 0.3233 0.6369 0.0963
-DIDA 0.5878 0.5118 0.5604 0.5620 0.6624 0.1950 0.6681 -0.0092
-ours DPPDCC 0.4463 0.7037 0.4138 0.7454 0.5349 0.4701 0.5633 0.2733
-#improve (%) 14.29 13.96 13.86 8.87 11.92 44.38 10.38 134.40
-(1) All models perform well on previous papers presented in the training set but experience degradation when dealing with fresh and immediate papers in the test set, particularly in terms of the LogR2.
-Many models fail to figure out the overall trend of immediate papers, which results in negative values in LogR2.
-Even the best baseline, MUCas, struggles with immediate papers due to limited graph structures, as it is a cascade prediction model designed for papers with sufficient citations.
-However, dynamic GNN models can effectively encode multi-hop information to complement the lack of cascade structures.
-Disentangled GNNs, in particular, effectively uncover crucial insights within both multi-hop graph structures and node representations, notably beneficial for datasets like Psychology, which are denser.
-Moreover, models such as SciBERT and HINTS leverage prior knowledge, like content information or citation distribution, to tackle these challenges.
-Comparatively, while SciBERT outperforms HINTS in fitting the increments of previous papers (SciBERT ¿ HINTS), it faces greater difficulty with fresh papers (SciBERT ¡ HINTS).
-This implies that graph structures exhibit better generality and transferability to unseen samples, while shared content information remains robust for immediate papers with limited graph structures, both alleviating issues related to data distribution shifts.
-(2) Both dynamic and heterogeneous graphs are important for our task.
-On one side, since temporal information within dynamic graphs exposes the evolved trend, it plays a crucial role.
-Dynamic GNNs possess comparable performance across all datasets for almost all categories of papers.
-In particular, DIDA, which considers dynamic graphs, stands out among the disentangled GNNs, closely matching MUCas’ performance across all datasets.
-On the other side, leveraging heterogeneous information can further enhance the performance of dynamic GNNs, as demonstrated by ROLAND.
-This general dynamic GNN framework achieves competitive performance akin to MUCas by integrating heterogeneous information.
-However, proper modeling of heterogeneous information is crucial; otherwise, it might lead to performance degradation.
-For instance, DisenHAN, compared to its backbone DisenGCN, struggles to encode valuable information within heterogeneous graphs, notably within the complex dataset Psychology.
-(3) While Disentanglement methods offer performance improvements, their adaptation to dynamic heterogeneous graphs is essential to achieve optimal flexibility.
-Notably, DIDA exhibits superior outcomes than ROLAND in LogR2 for Chemistry, even when relying solely on homogeneous graphs.
-Moreover, within the densest dataset, Psychology, Disentangled GNN models showcase impressive performance.
-Among them, CAL outperforms both MUCas and ROLAND in predicting fresh and immediate papers.
-However, in other scenarios, apart from DIDA, these methods face challenges in competing with task-specific models and dynamic GNNs.
-(4) Our proposed model surpasses all baselines, by effectively modeling the dynamic heterogeneous graph and incorporating task-adjusted disentanglement.
-This leads to an average improvement of about 10% in MALE across all categories, along with a substantial advancement of over 30% in LogR2 for fresh and immediate papers.
-First, compared with traditional citation/cascade prediction models, it exhibits superior generalization and practicability through the fusion of content and context information within citation networks, as well as the disentanglement power.
-Content information aids in discerning the impact of previous papers and fortifies robustness towards immediate papers.
-Moreover, its adeptness in handling fresh and immediate papers stems from the rich information embedded within graph structures.
-Furthermore, we consider the unique properties of citation networks to encode complex high-order structural and temporal information, thereby outperforming common dynamic models.
-The proposed CGE facilitates the identification of influential papers related to the target paper based on comparative and multi-hop information.
-More importantly, our model introduces interpretability from disentanglement models without compromising performance and further adapts to the task for enhanced improvement.
-It excels in predicting precise citation increments and approaches the genuine contribution of papers by disentangling values derived from information diffusion and collective conformity, thus providing insights into the papers’ authentic impact.
-5.2. Ablation Test (RQ2)
-To answer RQ2, ablation tests are performed on both CGE and PDM, highlighting the significance of various modules.
-As demonstrated in Table 5, for CGE, (-co-cited/citing) removes the proposed co-cited/citing strengths and only uses the original attention scores, (-CompGAT) replaces the CompGAT with original GATv2 (Brody et al., 2022), and (-readout) simply obtains the snapshot representation through avg-pooling over all paper nodes without filtering.
-For PDM, (-diffusion) removes the diffusion disentanglement, (-conformity) omits the conformity disentanglement, (-orthog) disregards the orthogonal constraints, and (-disen) eliminates all perspectives and replaces with a single MLP for prediction.
-Table 5. Results of ablation studies according to MALE.
+Appendix E Complete Experimental Results
+Table 7. The complete experimental results of total samples with standard deviations.
 model CS CHM PSY
-ours 0.4473 0.4355 0.4463
-CGE -co-cited/citing 0.4650 0.5014 0.4528
--CompGAT 0.4722 0.4990 0.4582
--readout 0.5116 0.4404 0.4932
-PDM -diffusion 0.4593 0.4530 0.4573
--conformity 0.5177 0.4747 0.4580
--orthog 0.4930 0.4388 0.4502
--disen 0.4606 0.4447 0.4670
-We observe that: (1) All modules within DPPDCC play pivotal roles, as evidenced by the decrease in performance when any module is removed.
-Their significance varies across datasets due to diverse graph characteristics.
-(2) For CGE, the removal of type-specific attention readout may cause a notable performance drop.
-That could be attributed to the presence of excessive noisy information within multi-hop snapshots, which necessitates the filtering of aggregated information for better concentration.
-Specially, in chemistry, where subgraphs are sparser, simple avg-pooling might suffice to extract meaningful information.
-Furthermore, our proposed CompGAT effectively captures distinct comparative information, surpassing the original GATv2 in computer science and psychology.
-The incorporation of co-cited/citing strengths further enhances multi-hop information beyond the local neighborhood, resulting in improved performance.
-(3) Eliminating any disentanglement task adversely affects performance.
-Moreover, an evident decline is observed after removing the whole disentanglement module, especially in Psychology.
-Conformity rises as the paramount task among all disentanglement tasks.
-It focuses on the Matthew effect, pervasive and discernible in reality, thus significantly contributing to model performance.
-5.3. Hyper-parameters Test (RQ3)
-(a) $L$
-(b) $\lambda$
-(c) $\beta$
-Figure 3. Results of hyper-parameters test in terms of MALE.
-(a) $L$
-(b) $\lambda$
-(c) $\beta$
-Figure 4. Results of hyper-parameters test in terms of LogR2.
-To address RQ3, we conduct experiments to assess the influence of three critical hyper-parameters in our model design: the stacked layer number $L$, the importance of co-cited/citing strengths $\lambda$, and the weight of the disentanglement tasks $\beta$.
-To ensure robust experimental outcomes, we focus on tuning these specific hyper-parameters while keeping other hyper-parameters fixed at their optimal values.
-In the following section, we report the corresponding results when tuning other important hyper-parameters.
-The experiment results are described in Figure 3 and 4.
-5.3.1. The Impact of Model Depth
-The number of layers $L$ serves as a pivotal hyperparameter for GNN-based models.
-Within this study, we explore values of $L$ in the range of 2, 3, 4, 5.
-For GNN-based models, increasing the number of layers facilitates the capture of higher-order connectivity within the graph during information propagation and aggregation.
-However, it also introduces the challenge of over-smoothing, a trend we’ve observed.
-As the layer count $L$ increases, model performance typically exhibits an initial rise followed by a decline—a phenomenon attributed to the interplay between increased connectivity capturing and eventual over-smoothing effects.
-In DPPDCC, besides GNN’s inherent challenges, the incorporated multi-hop co-cited/citing information and the intricate interaction between time and space dimensions could contribute to this phenomenon.
-Achieving the optimal hyperparameter $L$ requires finding a delicate balance between the GNN’s depth and its multi-hop/multi-dimensional information, ensuring an equilibrium that maximizes performance.
-5.3.2. The Impact of Co-cited/citing Strengths
-The weight parameter of co-cited/citing strengths $\lambda$ governs the proportion of co-cited/citing strengths in the sum distributions relative to the original attention scores.
-Our study investigates optimal values of $\lambda$ within the range 0.25, 0.5, 0.75, 1.0.
-Setting $\lambda$ to 1.0 essentially disregards attention mechanisms, replacing them with fixed weights and reducing the model to a weighted GCN.
-Higher weights can potentially dominate significant signals learned by the model, while lower weights may inadequately incorporate additional information.
-Hence, a moderate value such as 0.5 might represent a suitable setting adaptable across various datasets, balancing the incorporation of additional information without overshadowing the model’s learned signals.
-5.3.3. The Impact of Disentanglement Loss
-The weight of disentanglement loss $\beta$ controls the importance of the disentanglement tasks relative to the main task.
-Our study explores optimal values of $\beta$ within the range 0.25, 0.5, 0.75, 1.0.
-Notably, the disentanglement weight $\beta$ exhibits greater stability in comparison to $L$ and $\lambda$, indicating consistent influence and effectiveness of the disentanglement task as an auxiliary task.
-Selecting an appropriate $\beta$ value can further enhance overall model performance, underscoring the significance of this parameter in refining the model’s effectiveness.
-5.4. Visualization (RQ4)
-To answer RQ4, we visualize both CGE and PDM in the test set of Chemistry to delve deeper into our model’s mechanisms.
-For CGE, we employ linear regressions to uncover relationships among the sum distribution $\alpha$, original attention distribution $e$, and co-cited/citing distribution $c$.
-This analysis aims to elucidate the interplay and dependencies among these distributions.
-For PDM, we first calculate correlations between predicted/ground-truth values and the perspective values/proportions.
-Next, we visualize the comprehensive compositions of disentangled values concerning time, category, and value.
-Subsequently, we delve into the representation distribution of conformity perspective to validate its rationality.
-This process offers insights into how different perspectives contribute to predictions.
-5.4.1. Weighting Score Visualization
-(6.48/16.74)
+MALE↓ LogR2↑ MALE↓ LogR2↑ MALE↓ LogR2↑
+SciBERT 0.6904±.007 0.3370±.016 0.5868±.030 0.3887±.054 0.6131±.019 0.4761±.034
+SPECTER2 0.7482±.027 0.2337±.058 0.5832±.013 0.3906±.032 0.5935±.020 0.5082±.031
+HINTS 0.9022±.108 -0.1227±.231 0.8265±.074 -0.1283±.160 0.8529±.081 -0.0198±.204
+CasSampling 0.5061±.003 0.6112±.007 0.4679±.002 0.6127±.003 0.5196±.002 0.6200±.006
+H2CGL 0.4579±.007 0.6812±.014 0.4337±.003 0.6573±.010 0.4727±.005 0.6836±.006
+EGCN 0.8051±.009 0.1296±.009 0.6381±.007 0.2826±.020 0.7339±.012 0.2530±.021
+Dysat 0.6297±.006 0.4242±.015 0.5412±.009 0.4705±.017 0.6207±.047 0.4354±.096
+ROLAND 0.5978±.033 0.4911±.048 0.5437±.037 0.4495±.073 0.5521±.014 0.5718±.017
+DisenGCN 0.8321±.010 0.1026±.030 0.6343±.006 0.2897±.016 0.6689±.011 0.3746±.018
+DisenHAN 0.8221±.006 -0.0488±.038 0.8652±.181 -0.1110±.355 0.8766±.191 -0.1186±.552
+CAL 0.7094±.018 0.1859±.222 0.5733±.015 0.4166±.026 0.5535±.009 0.5783±.013
+DisC 0.6733±.006 0.3496±.013 0.5609±.031 0.4259±.065 0.5592±.021 0.5606±.035
+DIDA 0.6203±.034 0.4449±.045 0.5108±.016 0.5354±.035 0.5723±.022 0.5328±.034
+DyTed (EGCN) 0.8101±.008 0.1235±.008 0.6411±.007 0.2465±.006 0.7896±.009 0.1478±.025
+DyTed (Dysat) 0.6820±.035 0.2426±.138 0.5386±.010 0.4602±.013 0.5838±.001 0.5029±.001
+SILD 0.5836±.011 0.4977±.031 0.5009±.003 0.5629±.004 0.5448±.009 0.5889±.016
+DPPDCC 0.4432±.005 0.6938±.008 0.4264±.002 0.6667±.005 0.4587±.002 0.6969±.002
+#improve (%) 3.21 1.85 1.69 1.43 2.96 1.94
+We run the experiments with three random seeds and report both the averaged results and standard deviations of all models.
+As shown in Figure 7, our proposed model DPPDCC consistently outperforms others with minimal variation, highlighting its capacity and stability.
+Appendix F Complexity Experiment
+Table 8. The input data, parameters, and inference time of selected baselines and DPPDCC.
+Type model Input Data CS CHM PSY
+Dynamic Hetero params inference params inference params inference
+PLM SciBERT $\times$ $\times$ 111.1M 2.91s 111.1M 2.92s 111.1M 2.92s
+Cascade CasSampling $\times$ $\times$ 0.2M 0.02s 0.2M 0.02s 0.2M 0.02s
+Dynamic Dysat $\times$ 0.2M 0.66s 0.2M 0.76s 0.2M 1.24s
+ROLAND 0.5M 4.76s 0.5M 6.59s 0.5M 8.71s
+Disentangled CAL $\times$ $\times$ 0.3M 0.07s 0.3M 0.07s 0.3M 0.11s
+DIDA $\times$ 1.4M 2.26s 1.4M 2.55s 1.4M 4.89s
+DyTed $\times$ 2.1M 0.79s 2.1M 0.78s 2.1M 1.37s
+SILD $\times$ 2.2M 1.31s 2.2M 1.42s 2.2M 2.40s
+ours DPPDCC 14.3M 4.29s 31.0M 6.06s 14.3M 7.67s
+Our model’s complexity is primarily influenced by the R-GCN architecture, which requires distinct message-passing.
+During training, proposed contrastive learning involves additional encoding of multi-view graphs, but this step can be omitted during inference.
+Notably, our model converges quickly, requiring only 5 epochs compared to 20 epochs for other baselines, ensuring comparable training time.
+Given that we train and test on datasets with approximately 0.2 million samples, it ensures comparable coverage.
+Figure 8 demonstrates that our model has comparable space occupancy and time complexity to the baseline models.
+Appendix G Extended Visualizations
+G.1. Weighting Score Visualization
+(8.56/21.00)
 c
-(17.60/17.05)
-(40.76/73.79)
-(299.40/152.78)
-(1.52/4.35)
+(8.97/184.96)
+(19.23/106.80)
+(373.88/240.24)
+(1.98/5.42)
 (e) $\alpha^{(1)}=f(e^{(1)})$
-(4.15/4.36)
+(2.06/43.91)
 (f) $\alpha^{(2)}=f(e^{(2)})$
-(9.28/18.01)
+(4.46/26.08)
 (g) $\alpha^{(3)}=f(e^{(3)})$
-(58.70/36.62)
+($\inf$/50.61)
 (h) $\alpha^{(4)}=f(e^{(4)})$
-(1.67/4.17)
+(2.21/5.26)
 (i) $\alpha^{(1)}=f(c)$
-(4.53/4.28)
+(2.32/49.07)
 (j) $\alpha^{(2)}=f(c)$
-(10.67/19.02)
+(4.98/27.61)
 (k) $\alpha^{(3)}=f(c)$
-(85.97/40.87)
+(110.59/69.58)
 (l) $\alpha^{(4)}=f(c)$
 Figure 8.
-The outcomes of the linear regression analysis concerning the ”cites”/”is cited by” scores within the field of chemistry.
+The outcomes of the linear regression analysis concerning the ”cites”/”is cited by” scores within the field of computer science.
 Below the respective sub-figures, we depict the mean Kullback-Leibler (KL) divergences between the target and source distributions, denoted in units of $10^{-4}$.
-To effectively visualize the connections between distributions, we gather all scores subsequent to applying softmax based on the target node.
+To effectively visualize the connections between distributions, we gather all scores after applying softmax based on the target node.
 These scores are flattened to observe the overarching trend within the latest year.
 The original attention scores $e$ are averaged across all heads for ease of interpretation.
-Notably, for the negative effect in the co-citing relationship, we employ reversed values via $1-c$, rendering them positive to reflect the trend.
+Notably, we employ reversed values via $1-c$ for the negative effect in the co-citing relationship, rendering them positive to reflect the trend.
 In Figure 8, robustly positive relationships are observed among the sum distribution $\alpha$, original attention distribution $e$, and the distribution of co-cited/citing strengths $c$.
 It is reasonable to note that the sum distribution closely resembles either of the sourced distributions.
-When compared to the co-cited/citing distribution, the sum distribution demonstrates greater similarity to the original attention distribution.
+Compared to the co-cited/citing distribution, the sum distribution demonstrates greater similarity to the original attention distribution.
 Intriguingly, within the initial layers of the model, minimal disparities are evident between the original attention scores and the co-cited/citing strengths.
 This phenomenon is attributed to both the averaging across all heads and the model’s adaptation to the co-cited/citing information during the learning process.
-Moreover, while the overall trend is evident, a finer examination of the samples reveals that for instances with varying co-cited/citing strengths, their original attention scores tend to be equal, resulting in the displayed horizontal and vertical distribution evident in Figures 5a and 5e.
-This equal might arise due to the initial state of the attention mechanism in the Graph Attention Network (GAT).
-However, as the model progresses in depth, noticeable disparities emerge between these distributions in both ”cites” and ”is cited by” edges.
-In the upper layers, attention scores exhibit a tendency to concentrate on a select few nodes, while in the lower layers, they disperse attention across a wider array of nodes.
-Co-cited/citing strengths represent 2-hop information conveyed through intermediate nodes, resembling lower-hop information within GNNs.
-Therefore, continuously incorporating the co-cited/citing distribution across all layers reinforces signals originating from pivotal nodes across a broader spectrum.
+Moreover, while the overall trend is evident, a finer examination of the samples reveals that for instances with varying co-cited/citing strengths, their original attention scores tend to be equal, resulting in the displayed horizontal and vertical distribution evident in Figures 8a and 8e.
+This equal might arise due to the initial state of the attention mechanism in the GAT.
+However, as the model progresses in-depth, noticeable disparities emerge between these distributions in both ”cites” and ”is cited by” edges.
+In the upper layers, attention scores tend to concentrate on a select few nodes, while in the lower layers, they disperse attention across a wider array of nodes.
+Furthermore, co-cited/citing strengths represent 2-hop information conveyed through intermediate nodes, resembling lower-hop information within GNNs.
+Therefore, continuously incorporating the co-cited/citing distribution across all layers reinforces signals from pivotal nodes across a broader spectrum.
 It can alleviate the over-smoothing issues by preventing the dominance of a select few nodes within aggregated information.
@@
 This collaboration achieves multi-hop information integration, mitigating both forgetting and overemphasis within the learning process.
-5.4.2. Disentanglement Correlation Analysis
-Table 6. Results of Disentanglement Correlation in Chemistry
-dif con contri dif (%) con (%) contri (%) citations pred real
-dif 0.9463 0.9675 0.6177 -0.5847 0.3845 0.2159 0.9947 0.7804
-con 0.9463 0.8787 0.5178 -0.3892 0.1889 0.2765 0.9479 0.7501
-contri 0.9675 0.8787 0.5182 -0.6342 0.5124 0.1909 0.9828 0.7790
-dif (%) 0.6177 0.5178 0.5182 -0.6870 0.2795 0.0651 0.5641 0.3965
-con (%) -0.5847 -0.3892 -0.6342 -0.6870 -0.8897 -0.0157 -0.5760 -0.4305
-contri (%) 0.3845 0.1889 0.5124 0.2795 -0.8897 -0.0201 0.4067 0.3197
-citations 0.2159 0.2765 0.1909 0.0651 -0.0157 -0.0201 0.2226 0.2025
-pred 0.9947 0.9479 0.9828 0.5641 -0.5760 0.4067 0.2226 0.7895
-real 0.7804 0.7501 0.7790 0.3965 -0.4305 0.3197 0.2025 0.7895
-In Table 6, we depict the correlations among all disentangled perspective values/proportions and the predicted/real values, focusing on samples with all positively predicted values within the Chemistry test set.
-For papers containing negative disentangled values, they always have nearly no increased citations.
-Remarkably, over 80% of them feature negative because their contribution values are the lowest across all perspectives, demonstrating the rationality of our proposed method.
-Among the correlations observed among the values of all three perspectives, ”conformity” stands out as particularly distinct.
-Its correlation with ”contribution” is notably the lowest, falling below 0.9.
-Additionally, its correlation with the predicted/real values also demonstrates a significant gap (-0.05/-0.03) when compared to other perspectives.
-Despite this, it emerges as the most closely associated perspective with accumulated citations, implying its significance in representing information linked to a paper’s reputation.
+G.2. Disentanglement Correlation Analysis
+Table 9. Results of Disentanglement Correlation in Computer Science
+CS dif con contri dif (%) con (%) contri (%) citations final real
+dif 0.9579 0.9990 0.0389 -0.3237 0.4556 0.3209 0.9971 0.8276
+con 0.9579 0.9620 -0.1176 -0.1544 0.3495 0.3575 0.9765 0.8091
+contri 0.9990 0.9620 0.0276 -0.3236 0.4665 0.3239 0.9981 0.8287
+dif (%) 0.0389 -0.1176 0.0276 -0.7657 0.1967 -0.0364 -0.0022 -0.0003
+con (%) -0.3237 -0.1544 -0.3236 -0.7657 -0.7812 -0.0056 -0.2861 -0.2477
+contri (%) 0.4556 0.3495 0.4665 0.1967 -0.7812 0.0439 0.4382 0.3779
+citations 0.3209 0.3575 0.3239 -0.0364 -0.0056 0.0439 0.3331 0.2982
+final 0.9971 0.9765 0.9981 -0.0022 -0.2861 0.4382 0.3331 0.8298
+real 0.8276 0.8091 0.8287 -0.0003 -0.2477 0.3779 0.2982 0.8298
+CHM dif con contri dif (%) con (%) contri (%) citations final real
+dif 0.8441 0.9932 0.6134 -0.3982 -0.0478 0.1844 0.9828 0.7834
+con 0.8441 0.8723 0.2398 -0.0062 -0.3362 0.2581 0.9267 0.7317
+contri 0.9932 0.8723 0.5724 -0.3795 -0.0275 0.2035 0.9911 0.7912
+dif (%) 0.6134 0.2398 0.5724 -0.9104 0.4773 -0.0080 0.5075 0.4006
+con (%) -0.3982 -0.0062 -0.3795 -0.9104 -0.7981 0.0477 -0.2885 -0.2331
+contri (%) -0.0478 -0.3362 -0.0275 0.4773 -0.7981 -0.0897 -0.1263 -0.0883
+citations 0.1844 0.2581 0.2035 -0.0080 0.0477 -0.0897 0.2189 0.2022
+final 0.9828 0.9267 0.9911 0.5075 -0.2885 -0.1263 0.2189 0.7955
+real 0.7834 0.7317 0.7912 0.4006 -0.2331 -0.0883 0.2022 0.7955
+PSY dif con contri dif (%) con (%) contri (%) citations final real
+dif 0.8642 0.9972 0.5355 -0.5871 0.6302 0.3556 0.9903 0.8196
+con 0.8642 0.8740 0.2039 -0.2505 0.2997 0.4998 0.9244 0.7598
+contri 0.9972 0.8740 0.5014 -0.5609 0.6149 0.3572 0.9928 0.8199
+dif (%) 0.5355 0.2039 0.5014 -0.9869 0.9337 -0.0134 0.4522 0.3917
+con (%) -0.5871 -0.2505 -0.5609 -0.9869 -0.9792 0.0056 -0.5070 -0.4328
+contri (%) 0.6302 0.2997 0.6149 0.9337 -0.9792 0.0043 0.5570 0.4683
+citations 0.3556 0.4998 0.3572 -0.0134 0.0056 0.0043 0.4017 0.3416
+final 0.9903 0.9244 0.9928 0.4522 -0.5070 0.5570 0.4017 0.8256
+real 0.8196 0.7598 0.8199 0.3917 -0.4328 0.4683 0.3416 0.8256
+In Table 9, we illustrate the correlations among all disentangled perspective values/proportions and the predicted/real values, focusing on samples with all positively predicted values across all datasets.
+Taking the correlation analysis of Computer Science as an example, ”conformity” stands out among the values of all three perspectives.
+Its correlation with the predicted/real values shows a significant gap (-0.02) compared to other perspectives.
+Despite this, conformity emerges as the most closely associated perspective with accumulated citations, implying its significance in representing information about a paper’s reputation.
 The predicted increment appears as the most influential predictor of the real increment.
@@
 (2) The DPPDCC model adeptly captures the trend in a paper’s citation increment, validating its substantial potential and reliability.
-In our calculations of the proportions of all three perspectives within predicted values, a notable finding emerges: conformity diffusion stands out as the only perspective with numerous negative correlations.
-This observation suggests that in the field of Chemistry, papers solely characterized by accumulated citations tend to attract fewer new citations.
-Moreover, it shows an almost negative correlation with the proportion of contribution (-0.89), indicating their mutual exclusivity.
+In our calculations of the proportions of all three perspectives within predicted values, we find that conformity diffusion is the only perspective with numerous negative correlations.
+This suggests that in the field of Computer Science, papers characterized solely by accumulated citations tend to attract fewer new citations.
+Additionally, it shows an almost negative correlation with the proportion of contribution (-0.78), indicating their mutual exclusivity.
 Consequently, conformity and contribution have learned entirely distinct facets of citation increments.
-Overall, the diffusion perspective can be seen as the most pivotal factor for prediction, as it showcases the highest correlations in both values and proportions.
-This prominence aligns with the intuitive connection between citation increments and the information diffusion process.
-Conversely, the secondary perspective, contribution, likely delves into factors beyond the graph structure, aiming to uncover additional crucial elements.
-As a complement to the numerical citation increment, it offers a more objective evaluation, potentially aiding in the identification of valuable papers by encompassing diverse and complementary information.
-5.4.3. Disentangled Composition Visualization
-(a) time trend
-(b) group detail
-(c) value bin
+Overall, the contribution perspective is the most pivotal factor for prediction, showcasing the highest correlations in both values and proportions.
+However, the correlations remain a flexible threshold, where other perspectives also help determine the predicted increments.
+As a complement to the numerical citation increment, this approach offers a more objective evaluation, potentially aiding in identifying valuable papers by encompassing diverse and complementary information.
+In summary, contribution is the most significant perspective across all datasets.
+Its disentangled value has the largest correlation with both the predicted and real citation increments, as does its proportion in Computer Science and Psychology.
+Moreover, diffusion and conformity focus on different aspects of the popularity factor.
+Accumulated citations consistently explain the trend of conformity values, while diffusion complements the structural influence with more flexibility regarding graph features.
+G.3. Complete Disentanglement Composition
+(a) time trend (CS)
+(b) group detail (CS)
+(c) value bin (CS)
+(d) time trend (CHM)
+(e) group detail (CHM)
+(f) value bin (CHM)
+(g) time trend (PSY)
+(h) group detail (PSY)
+(i) value bin (PSY)
 Figure 9.
-Visualization of disentangled value proportions in Chemistry.
-(a) displays the trend that evolved with the publication time.
-(b) demonstrates the detailed composition of papers categorized into previous, fresh, and immediate ones.
-(c) is binning the samples based on the predicted values.
-We visualize the disentanglement composition with all positive values in the test set of Chemistry.
-For the stacked plot associated with time in Figure 6a, a clear trend emerges: the contribution proportion gradually increases, while the conformity proportion undergoes a significant drop.
-This trend is sensible since old papers tend to accumulate substantial citations, thus attracting more researchers to cite them due to inertia.
-Additionally, the diffusion proportion appears to be more stable across different times, as it represents the potential hidden properties within diffusion.
-This stability becomes more pronounced when papers are categorized into previous, fresh, and immediate ones in Figure 6b.
-Diffusion proportions remain steady in both mean and standard deviation.
-Moreover, contribution emerges as the most critical perspective, with its dominant proportion.
-It further holds particular significance for fresh and immediate papers.
-For immediate papers receiving almost no citations, both the diffusion and conformity perspectives struggle to offer predictive insights.
-However, the contribution enables the model to anticipate the potential of samples in advance by analyzing their citation context.
-This capability enhances the model’s robustness, particularly for challenging samples close to cold-start scenarios.
-Furthermore, in Figure 6c, when we bin the samples based on the predicted logarithmic values, we observe a descending trend for conformity and a growing trend for contribution.
-Notably, within the lowest bin group, the conformity proportion surpasses the contribution proportion for the first time.
-The alignment between time and impact groups can unveil an overall trend: in dynamic contexts, previous papers tend to attract fewer new citations and may primarily owe to their conformity, whereas recent and immediate papers are more likely to garner increased citations owing to their genuine contributions.
-5.4.4. Disentangled Conformity Representation Visualization
-(a) total
-(b) prev
-(c) fresh
-(d) immed
-Figure 7. Results of T-SNE visualization for representations from conformity perspective in Chemistry. Various colors correspond to the binning labels assigned to the samples.
-To validate that disentangled perspectives learn distinct aspects of the features, we visualize the representations extracted from the conformity encoder with T-SNE employed by scikit-learn (Pedregosa et al., 2011), as this perspective is the only one with available labels.
-Initially, the T-SNE algorithm is executed on the complete test sets of each category.
-Subsequently, 3000 points are randomly sampled to generate the figures.
-The points are color-coded based on the binning labels of the cumulative citations at the test time point.
-Figure 7 illustrates a consistent and structured distribution of points across all figures, indicating our model’s effective capture of group discrepancies across various binning labels in all categories.
-Across different categories, variations in the distribution of binning labels are evident due to differences in cumulative citations over time.
-Fresh Papers published within the last five years, particularly immediate ones, tend to fall under the lowest accumulated citation label (red), dominating a significant area.
-Nevertheless, discernible boundaries persist between distinct binning labels within these categories.
-This suggests the model’s capacity to generalize and transfer learned group differences to unseen samples, showcasing its potential applicability in real-world scenarios.
-5.5. Case Study (RQ4)
+Visualization of disentangled value proportions.
+(a)/(d)/(g) displays the trend that evolved with the publication time.
+(b)/(e)/(h) demonstrates the detailed composition of papers categorized into previous, fresh, and immediate ones.
+(c)/(f)/(i) is binning the samples based on the predicted values.
+Figure 9 illustrates the disentangled compositions across all datasets.
+We observe common trends, such as the similar pattern of conformity and the dominance of contribution.
+Additionally, details such as the respective slopes and nuanced composition for various categories vary according to the intrinsic characteristics of the corresponding fields.
+Appendix H Case Study
 Table 10. Detailed meta-data, disentangled values/proportions, and potential impact of the selected cases.
@@
 pub time refs citations dif con contri dif con contri pred real
-242481 2012 28 43 1.2272 0.8977 1.4735 34.10% 24.95% 40.95% 3.5984 3.0445
-15676 2001 19 103 1.2151 0.9443 1.4569 33.60% 26.11% 40.29% 3.6163 3.0445
-374215 2015 60 26 1.1949 0.8409 1.6318 32.58% 22.93% 44.49% 3.6675 3.2189
-(a) 242481-$c$
-(b) 242481-$e$
-(c) 242481-$\alpha$
-(d) 15676-$c$
-(e) 15676-$e$
-(f) 15676-$\alpha$
-(g) 374215-$c$
-(h) 374215-$e$
-(i) 374215-$\alpha$
+981400 2016 28 7 (3/2) 1.2175 0.7428 1.1862 38.69% 23.61% 37.70% 3.1466 3.0445
+47429 2009 12 51 (15/3) 1.1432 0.7895 1.1993 36.50% 25.21% 38.29% 3.1320 3.1355
+88781 2014 37 5 (1/0) 1.1984 0.7247 1.2307 38.00% 22.98% 39.02% 3.1538 4.7449
+(a) 981400-$c$
+(b) 981400-$e$
+(c) 981400-$\alpha$
+(d) 47429-$c$
+(e) 47429-$e$
+(f) 47429-$\alpha$
+(g) 88781-$c$
+(h) 88781-$e$
+(i) 88781-$\alpha$
 Figure 10.
-Weighted paper ”cites” subgraphs of selected cases organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ in Chemistry.
+Weighted paper ”cites” subgraphs of selected cases organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ in Computer Science.
 We filter the edges within 2-hop graphs by retaining only the top 30% accumulated weights of each node and preserve the direct citation or reference nodes of the target paper to attain these sparser graphs.
@@
 In this section, we visualize graph structures organized by co-cited strengths $c$, average original attention scores $e$, and sum scores $\alpha$ within ”cites” edge subgraphs of the selected papers.
-To maintain clarity, we focus on graphs extracted from the latest year of the snapshots and the final layer of DPPDCC.
+To maintain clarity, we focus on graphs extracted from the latest year of the snapshots and average scores across all layers of DPPDCC.
 We then filter the edges in 2-hop graphs based on the top 30% accumulated weights, resulting in sparser graphs.
@@
 Our labeling strategy encompasses the target paper and its directly connected highly-cited citations.
-As outlined in Table 7, 242481, 15676, and 374215 respectively embody the diffusion, conformity, and contribution perspectives.
-Despite their near-identical predicted and real increments, their disentangled value compositions differ from one another.
-Notably, 374215 stands as a recent paper not encountered during the training phase, while 242481 and 15673 are earlier publications, each with varying years since publication and distinct accumulated citations.
-These papers exhibit perspective proportions aligning with all trends discussed in Section 5.4.3.
+As outlined in Table 10, papers 981400, 47429, and 88781, respectively, embody the diffusion, conformity, and contribution perspectives.
+Despite their near-identical predicted increments, their disentangled value compositions differ from one another.
+Notably, 981400 and 88781 are recent papers not encountered during the training phase, while 47429 is an earlier publication, each with varying years since publication and distinct accumulated citations.
+These papers exhibit perspective proportions aligning with all trends discussed in Section E.
 As depicted in Figure 10, $c$ subgraphs manifest as the densest, while $e$ and $\alpha$ subgraphs appear considerably sparser, featuring fewer interconnected nodes.
 In the higher layers of DPPDCC, the model tends to concentrate on specific nodes, resulting in sparsity in these graphs.
-The introduced co-cited strengths $c$ serve as a complementary factor to the original attention scores $e$, aiding in identifying valuable citations from highly-cited papers.
-Notably, $c$ consistently incorporates highly-cited papers overlooked in $e$ subgraphs.
-For instance, Figures 8b and 8e consider only one highly-cited paper each (532725 and 91549).
-However, in Figures 8c and 8f, three additional highly-cited papers (559976, 532726, 180237 and 26005, 582958, 193949) - initially presented in Figures 8a and 8d - are taken into account.
-For 242481, within the test set, it stands as a recent publication with limited citations due to its relatively short existence.
-Despite its lower accumulated citations compared to 15676, 242481 has garnered significant attention from four highly-cited papers, equaling the number received by 15676.
-Consequently, its average influential citation surpasses that of 15676, leading to a higher proportion of the diffusion perspective in its incremental source.
-Furthermore, since subgraphs depicted in Figure 8 are based on the ”cites” edge, only edges from valuable citations are retained.
-In contrast to 15676, 242481 boasts numerous connections from references, indicating its significance as an important citation from the perspective of its references.
-Consequently, it emerges as a pivotal element in the information diffusion process, contributing to its high diffusion values.
-For 15767, being the oldest paper with a publication span of over 10 years by the test time point (2017), its conformity value stands out as the highest among all cases, contributing significantly to a larger proportion of its incremental impact.
-Despite its extensive citation count, considering the evolving context of the citation network, its significance has diminished.
-Over time, citations toward this paper tend to lean more towards its established reputation rather than other factors.
-For 374215, it is the most recent paper with limited citations.
-Because it has neither highly-cited citations nor enough accumulated citations, compared to either 242481 or 15767, both its diffusion and conformity perspectives are low in values and proportions.
-Conversely, its contribution perspective appears as the most pivotal, showcasing the largest values and proportion.
-Benefiting from its contribution, 374215 successfully surpasses the other two papers in predicted/real potential impact.
-Interestingly, beyond the factors highlighted in the graph structures or accumulated citation counts, something intrinsic to its contribution stands out, as delineated by DPPDCC.
-This factor distinguishes itself from mere popularity, emphasizing DPPDCC’s substantial potential for practical application and impact.
-Indeed, our proposed DPPDCC effectively disentangles citation increments derived from various factors for the target paper.
-This disentanglement, reflected in the values and proportions assigned to distinct perspectives, provides more comparable metrics for identifying valuable papers.
-In future work, we aim to extend this framework and assess its practical application in uncovering novel insights within real-world contexts.
-This ongoing exploration will enable us to delve deeper into new discoveries and applications facilitated by our disentanglement approach.
-6. Conclusion
-In this study, we propose a novel model named DPPDCC that disentangles the potential impacts of papers into diffusion, conformity, and contribution values.
-To better align with real-world conditions, we reformulate the problem by partitioning the datasets based on the observation time point, without excluding lowly-cited papers.
-Further, DPPDCC comprises two principal components: Citation-aware GNN Encoder (CGE) and Popularity-aware Disentanglement Module (PDM).
-Given the dynamic heterogeneous graph of a target paper, CGE employs CompGAT to encode the comparative and co-cited/citing information between papers and utilizes Type-specific Snapshot Readout to aggregate the information from each snapshot in an evolutionary manner.
-Subsequently, PDM identifies amplifying effects in information diffusion and addresses the Matthew effect stemming from accumulated citations from a collective conformity perspective.
-Finally, it constrains each perspective to address distinct aspects for preserving the genuine contribution.
-Experimental results on three datasets demonstrate that DPPDCC significantly outperforms alternative approaches for previous, fresh, and immediate papers.
-Disentanglement visualizations and a detailed case study provide additional evidence supporting DPPDCC’s reliable predictive capabilities.
+The introduced co-cited strengths $c$ act as a complementary factor to the original attention scores $e$, facilitating the identification of significant citations from highly-cited papers.
+Notably, $c$ consistently integrates significant highly-cited papers that were overlooked, replacing the original set with more valuable ones in $e$ subgraphs.
+For instance, Figure 10b considers only one highly-cited paper (541312).
+However, in Figure 10c, one additional highly-cited paper (437859) - initially presented in Figure 10a is taken into account.
+Similarly, although Figure 10c and 10b share the same number of highly-cited papers (3), the co-cited subgraph $c$ replaces one of them (37213 $\rightarrow$ 51553).
+For paper 981400, within the test set, it stands as a recent publication with limited citations due to its relatively short existence.
+Although it has fewer accumulated citations than paper 47429, 981400 has attracted significant attention from two highly-cited papers, which is noteworthy given its total of seven citations.
+Consequently, its average influential citation (28.6%) exceeds that of 47429 (5.9%), resulting in a higher proportion of the diffusion perspective in its incremental source.
+This positions it as a crucial element in the information diffusion process, contributing to its high diffusion values.
+On the other hand, paper 47429, the oldest paper in the test set with a publication span of nearly 8 years by 2017, stands out with the highest conformity value among all cases, significantly contributing to a larger proportion of its incremental impact.
+Despite its extensive citation count, its significance has diminished over time in the evolving context of the citation network, with citations tending to emphasize its established reputation rather than other factors.
+Paper 88781, also recent with limited citations, lacks both highly-cited and accumulated citations compared to papers 981400 or 47429.
+Consequently, it exhibits low values and proportions in both diffusion and conformity perspectives.
+However, its contribution perspective emerges as pivotal, showcasing the largest values and proportion.
+Leveraging its contribution, 88781 surpasses the other two papers in predicted/real potential impact.
+Interestingly, DPPDCC highlights intrinsic factors to its contribution beyond popularity-related factors highlighted in graph structures or accumulated citation counts, emphasizing DPPDCC’s substantial potential for practical application and impact.
+Indeed, our proposed DPPDCC effectively disentangles citation increments from different factors for the target paper.
+This disentanglement, as evidenced by the values and proportions assigned to distinct perspectives, offers more comparable metrics for identifying valuable papers.
+In future endeavors, we aspire to expand this framework and evaluate its practical utility in uncovering novel insights within real-world scenarios.
+This continued exploration will allow us to delve further into discoveries and new applications facilitated by our disentanglement approach.
+Appendix I Limitations
 In our future endeavors, we intend to develop extensive datasets spanning diverse fields of study to simulate practical scenarios.